### 09 Dec 2023 v1

# Chapter 1      Introduction to IoT

Abstract
The Internet of Things (IoT) has rapidly transformed the 21st century, enhancing decision-making processes and introducing innovative consumer services such as pay-as-you-use models. This integration of smart devices and automation technologies has revolutionized our lives. However, it is essential to recognize the significant concerns surrounding security, privacy, intellectual property rights, safety, and trust in this technological landscape, which require further exploration. This chapter serves as a comprehensive guide to newcomers interested in the IoT domain, providing a foundation for making future contributions. It begins by explaining the core concept of IoT, its historical evolution, network components, and key characteristics. The advantages of IoT deployment across various domains are discussed, along with an introduction to the foundational layered architectures supporting IoT. The chapter also delves into the taxonomy of IoT technologies, encompassing hardware, software, wireless communication technologies, hardware platforms, and cloud solutions. Existing applications in major domains like smart cities, healthcare, and agriculture are explored, offering a broad perspective on the IoT’s application potential. In addressing prevalent issues and challenges in designing and deploying IoT applications, the chapter examines security threats across architectural layers, ethical considerations, user privacy concerns, and trust-related issues. This discussion equips researchers with a solid understanding of diverse IoT aspects, providing a strong foundation for both practical IoT projects and the development of novel theoretical approaches within the IoT field.
Keywords: Taxonomy, Architectural Layers, Ethical Considerations, Security Threats, Hardware Platforms, Trust-related Issues.
1.1 Introduction
In today’s ubiquitous digital landscape, the Internet has profoundly impacted global existence, marking an ongoing journey towards even more pervasive connectivity—ushering in the era of the Internet of Things (IoT). A groundbreaking invention in recent decades, IoT revolutionizes the interaction between the physical and digital realms, as defined by Vermesan et al. [1]. In this interconnected landscape, the digital world engages with the physical world through an array of sensors and actuators. Pena-López et al. [2] offer an expansive interpretation, characterizing IoT as a paradigm where computing and networking seamlessly integrate into virtually any object, empowering remote querying and modification. Broadly, the term ”Internet of Things” describes a transformative realm where nearly every daily device is intricately linked to a network, enabling collaborative utilization for intelligent and automated tasks. The concept of IoT was first introduced by Peter T. Lewis in 1985, defining it as the fusion of individuals, processes, and technology with interconnected devices and sensors. This facilitates remote monitoring, status assessment, manipulation, and trend analysis of these devices [3]. The IoT journey, far from its conclusion, promises a future where diverse devices seamlessly connect to the web, reshaping human existence in unprecedented ways.
The Internet of Things (IoT) is a network of interconnected devices with sensors, actuators, processors, and various communication technologies. Sensors collect real-time data from both internal states and external surroundings, ranging from mobile phones to microwave ovens. Actuators, in turn, respond to data or commands, enabling automation and remote control of physical devices. Data collected by sensors undergoes processing either at the network edge or on central servers, with some preprocessing occurring directly in the sensors or end devices. Processed data is then transmitted to remote servers for further analysis, storage, and processing. These data form the basis for analysis, decision-making, and subsequent actions, which can be physical (e.g., adjusting a smart thermostat) or virtual (e.g., sending notifications) [4]. The applications of IoT are extensive and diverse, impacting various aspects of our lives. From personal convenience in smart homes to healthcare and fitness innovations, IoT has the potential to influence personal, financial, physical, educational, professional, and mental aspects of individuals. In smart homes, IoT enables remote control of electrical appliances, lighting, coffee brewing, thermostat adjustments, and even hands-free operations through voice commands [5]. In healthcare, wearable IoT devices offer remote monitoring, allowing caregivers and healthcare professionals to provide timely assistance in emergencies. Additionally, individuals can use wearable devices to track sleep patterns, physical activity, and overall fitness [6]. These examples only scratch the surface of IoT’s broad application landscape, indicating the exciting possibilities and challenges that researchers are exploring for the future.
The Internet of Things (IoT) has the potential to transform how people interact with technology, providing greater convenience, efficiency, and personalization in daily life. Despite its transformative impact, IoT faces challenges. The sheer number of devices and the substantial data generated pose significant challenges, with a projected 41.6 billion IoT devices producing 79.4 zettabytes of data by 2025 [7]. Addressing this requires scalable architectures and enhanced processing capabilities. Moreover, IoT heavily relies on wireless communication, leading to challenges such as distortion and unreliability in geographically dispersed locations. Ensuring dependable data transmission becomes a pivotal challenge, emphasizing the critical role of communication technologies in the IoT landscape. Beyond technical hurdles, various general and domain-specific challenges are crucial for the success of IoT. Identifying and addressing these multifaceted challenges collectively is essential to unlock the full potential of IoT and overcome obstacles for broader adoption and integration into our lives.
Motivation of this chapter: IoT is not a novel concept; researchers have been exploring this field for decades. Consequently, the question arises: what sets this chapter apart from others, and why should readers invest their time in exploring the basics presented here? While there are several book chapters introducing IoT, many share outdated concepts, lacking the latest insights from the ongoing exploration in this dynamic field. For instance, in [8], only IoT applications are discussed, and in [9], authors overlook major domains, including IoT applications, advantages, challenges, and technologies. Notably, Nagaraj et al. [10] present a well-structured discussion covering technologies, architectures, applications, and challenges. However, their discussion is limited to these aspects, with a less extensive exploration of the application section. Therefore, the need arises for a comprehensive chapter that covers all aspects and incorporates recent additions in the IoT field. This chapter goes beyond the basics, encompassing fundamental components, characteristics, and advantages. It delves into architectures, provides a taxonomy of technologies used in IoT, explores a significant number of applications across diverse domains, and addresses ethical considerations as well as legal and regulatory issues. In essence, our chapter serves as a holistic guide, covering the most significant facets of IoT, from foundational principles to emerging research challenges and future directions. Readers are encouraged to explore this chapter for a thorough and up-to-date understanding of the evolving landscape of IoT.
Authors’ Contributions: The authors’ contributions in this chapter can be summarized as follows:
  1. 1.
    This chapter provides an introductory overview of the IoT with the aim of assisting future contributors. It assesses the benefits, generic architecture, key technologies underpinning IoT, and its diverse applications across various domains.
  2. 2.
    Recent studies have been examined to facilitate this assessment. Additionally, this chapter encompasses a taxonomy of IoT technologies, including extensive coverage of field communication.
  3. 3.
    In conclusion, this chapter offers a comprehensive discussion of the open research challenges, ethical considerations, as well as the legal and regulatory aspects of IoT.
Chapter organization: It begins with an exploration of the fundamentals in Section 1.2, covering the history 1.2.1, components 1.2.2, and characteristics 1.2.3. In Section 1.3, the advantages of IoT are examined, with a focus on both the general perspective and specific domains such as healthcare and agriculture. By introducing a generic architecture in Section 1.4, the taxonomy of technologies that underpin IoT operations are discussed in detail in Section 1.5. Furthermore, Section 1.6 explores the extensive array of application domains of IoT, with a focus on multiple promising research works within each domain, aiming to provide a comprehensive understanding of its real-world applications. It concludes in Section 1.7 by addressing the key challenges that IoT faces today, considering insights from the reviewed literature, and sheds light on potential future scopes and developments in the ever-evolving realm of IoT technology.
1.2 Fundamentals of IoT
At its core, the IoT revolves around the concept of connecting everyday objects and devices to the internet, enabling them to communicate, collect data, and perform actions autonomously or in response to commands. This basic concept involves three key elements: sensors and actuators that gather and interact with data, a network infrastructure for data transmission, and cloud-based platforms for data storage, processing, and analysis. By interconnecting these elements, IoT empowers us to enhance efficiency, gain real-time insights, and create smart, responsive systems that impact various aspects of our lives, from smart homes and cities to industries and healthcare. The evolution of IoT, which has unfolded over several decades, began with the concept of smart objects. Since then, IoT has undergone numerous groundbreaking transformations that have left the world astounded, thanks to its unique and convenient characteristics. The advantages brought by this technology are unquantifiable, as it touches nearly every aspect of people’s lives, simplifying and improving them. This section comprehensively discusses these fundamental IoT concepts, covering the following topics: The History of IoT, Components of IoT, Characteristics of IoT.
1.2.1 Historical Development of IoT
Since the invention of the first landline, the telegraph, in the 1830s and 1840s, machines have been instrumental in facilitating direct communication. A significant step towards the IoT occurred on June 3, 1900, with the first radio voice transmission, often referred to as “wireless telegraphy”. This paved the way for IoT. The development of computers, which began in the 1950s, is another crucial aspect of IoT.
In 1962, the Internet, a fundamental component of IoT, started as a DARPA [^1] The Defense Advanced Research Projects Agency is a research and development agency of the United States Department of Defense responsible for the development of emerging technologies for use by the military. project. A group of renowned researchers initiated efforts to connect computers and systems. By 1969, DARPA had evolved into ARPANET [^2] The Advanced Research Projects Agency Network (ARPANET) was the first wide-area packet-switched network with distributed control and one of the first computer networks to implement the TCP/IP protocol suite. Both technologies became the technical foundation of the Internet., a precursor to today’s Internet [11].
While the term “Internet of Things” is relatively new, the concept of integrating computers and networks to monitor and manage devices has a rich history spanning decades. In the late 1970s, various stakeholders, including businesses, governments, and consumers, began exploring ways to connect personal computers (PCs) and other machinery. This led to the practical use of systems for remotely monitoring electrical grid meters via telephone lines during that era [12].
The 1980s saw a growing interest in enhancing physical objects with sensors and intelligence. Commercial service providers began supporting public access to ARPANET, an early precursor to the modern Internet, during this time. Satellites and landlines played pivotal roles in establishing the foundational communication infrastructure for the emerging IoT. The concept of a network of smart devices was initially explored as early as 1982 when a Coca-Cola vending machine [^3] The ”Only” Coke Machine on the Internet. Available at: Carnegie Mellon University at Carnegie Mellon University was modified to connect to ARPANET. This allowed local programmers to remotely monitor the vending machine’s contents, ensuring drinks were available and cold before making a purchase. However, the technology was challenging to manage, and progress in this field was limited during that period [13]. In parallel, during the 1980s, local area networks (LANs) gained popularity and proved effective for real-time communication and document sharing among groups of PCs.
In the 1990s, advancements in wireless technology set the stage for the widespread adoption of ”machine-to-machine” (M2M) solutions in enterprise and industrial contexts, particularly for equipment monitoring and operation. However, many of these early M2M solutions relied on closed, purpose-built networks and proprietary or industry-specific standards, rather than utilizing Internet Protocol (IP)-based networks and Internet standards [14]. By the mid-1990s, the Internet had expanded its global reach, offering new possibilities for researchers and technologists to explore ways to enhance connections between humans and machines. A significant milestone in this journey was the creation of the first Internet-connected ’device’ by John Romkey—an IP-enabled toaster that could be controlled over the Internet. This innovative toaster was showcased at an Internet conference in 1990, marking an early example of the IoT in action [15].
In 1991, Mark Weiser’s paper “The Computer of the 21st Century” [16] and academic events like UbiComp and PerCom shaped the contemporary IoT vision [14]. Global Positioning Satellites (GPS) became a reality in early 1993 with the Department of Defense establishing a stable system of 24 satellites. Privately owned commercial satellites soon followed, enhancing IIoT functionality [3]. In early 1994, Reza Raji introduced the IoT concept in IEEE Spectrum, describing it as “moving small data packets to integrate and automate from home appliances to entire factories” [17]. Steve Mann’s near-real-time WearCam, powered by a 64-processor setup, was invented later that year. Between 1993 and 1997, companies proposed IoT solutions, including Microsoft’s “at Work” and Novell’s NEST. Momentum grew as Bill Joy introduced device-to-device communication in his ”Six Web” framework at the 1999 World Economic Forum [3].
The term “Internet of Things” was coined by Peter T. Lewis in a 1985 speech during the Congressional Black Caucus Foundation’s 15th Annual Legislative Weekend in Washington, D.C. Lewis defined IoT as the integration of people, processes, and technology with connectable devices and sensors for remote monitoring, status assessment, manipulation, and trend evaluation related to these devices [3]. In 1997, Paul Saffo described sensors and their future role. British technologist Kevin Ashton, while serving as the executive director of the Auto-ID Centre at MIT, independently coined the term “Internet of Things”. During his time at Procter & Gamble, he explored radio-frequency identification (RFID), a technology framework enabling physical devices to connect via microchips and wireless signals. In the same year, they developed a global RFID-based item identification system [14].
In 1999, Kevin Ashton was the first to describe the IoT and proposed the name “Internet of Things” during a presentation for Procter & Gamble. He believed RFID technology, primarily designed for inventory tracking, was a significant prerequisite for the Internet of Things, allowing computers to efficiently manage and monitor individual objects. The concept of tagging objects has been realized through technologies like digital watermarking, barcodes, and QR codes, used for identification and tracking purposes [18]. Subsequent technological advancements, including the proliferation of smartphones, cloud computing, improved processing power, and enhanced software algorithms, along with the availability of sophisticated sensors capable of measuring various parameters, laid the foundation for robust data collection, storage, and processing for the IoT’s growth.
As a significant step forward in commercialising IoT, LG announced plans to launch a smart refrigerator capable of autonomously managing its contents in 2000. Walmart and the US Department of Defense pioneered inventory tracking using RFID and the IoT in $2002-2003$. RFID gained prominence in the US Army’s Savi program in 2003, and Walmart expanded its RFID usage worldwide that same year. In 2004, Cornelius “Pete” Peterson, CEO of NetSilicon, predicted that IoT devices would dominate the next era of information technology, particularly in fields like medical devices and industrial controls [3]. In 2005, numerous articles in mainstream newspapers such as The Guardian, Scientific American, and The Boston Globe discussed IoT’s future direction.
The IPSO Alliance was founded in 2008 to promote the use of IP in networks of “smart objects”, while the FCC allowed the use of the “white” label in 2008. Google initiated the development of autonomous cars in 2009, and in 2011, Google’s Nest smart thermostat entered the market, enabling remote heating management. In June 2012, major Internet service providers and web-based companies agreed to expand the global Internet’s address space by enabling IPv6 for their services and products, a significant step towards a viable IoT. This led to substantial growth and interest in the field. IT giants like Cisco, IBM, and Ericsson later took numerous educational and commercial initiatives related to IoT. Cisco Systems estimated that the IoT was ”born” between 2008 and 2009, with the things/people ratio growing from 0.08 in 2003 to 1.84 in 2010 [18].
1.2.2 Components of IoT
The IoT consists of several key components that serve as the essential building blocks for constructing an IoT system. This section provides an in-depth exploration of the principal components within IoT. IoT comprises three main components: (1) Sensors/Devices and Actuators, (2) Storage and Data Analytics, and (3) Interpretation and Visualization Tools. Each of these is further categorized into various subcomponents.
Sensors/Devices and Actuators
  1. 1.
    Sensors/Devices
    Sensors play a crucial and essential role within an IoT system. Given that IoT operates by gathering data from the surrounding environment, it is necessary for all IoT applications to incorporate one or more sensors to meet this need. A defining characteristic of IoT devices is their context awareness, which is made possible through the utilization of sensor technology. Sensors are not only compact and cost-effective but also energy-efficient. However, they are subject to limitations such as battery capacity and ease of deployment. [4]. An overview of various types of sensors has been provided below.
    1. (a)
      Mobile based Sensors
      Smartphones, which are widespread and commonly used, are equipped with various sensors, including accelerometers and environmental monitoring sensors. Given their extensive usage, researchers are exploring the potential of using smartphones as integral components in building smart IoT solutions. These applications can harness sensor data from smartphones to generate valuable insights and outcomes. Some of the general sensors found in smartphones include [19]:
      1. i.
        Accelerometer: This sensor monitors and detects the motion or acceleration of the device, typically measuring changes in the smartphone’s velocity in three dimensions [20]. The data patterns recorded by the accelerometer can be employed to identify the user’s physical activities, including running, walking, and cycling.
      2. ii.
        Gyroscope: The gyroscope introduces an extra dimension compared to the accelerometer by being able to detect rotation or tilting of the phone.
      3. iii.
        GPS: The Global Positioning System (GPS) employs satellites to determine the location of a mobile phone.
      4. iv.
        Magnetometer: The magnetometer detects the magnetic field and can function as a compass.
      5. v.
        Light Sensor: The light sensor measures the intensity of light in the environment around it. It can be applied in different ways, such as managing the room’s lighting or adjusting the screen’s brightness.
      6. vi.
        Proximity Sensor: The proximity sensor employs an infrared (IR) LED that emits infrared rays. These rays bounce off objects they encounter, and by measuring the time it takes for them to return, one can calculate the distance to those objects. This technology enables us to measure how far things are from the phone.
      Certain smartphones, like the Samsung Galaxy S4, come equipped with extra sensors, including a thermometer, barometer, and humidity sensor. These sensors are used to measure temperature, atmospheric pressure, and humidity levels, respectively [4].
    2. (b)
      Medical Sensors
      The healthcare industry is one of the most influenced fields where innovation and IoT have paved the way. Wearable devices and sensors have facilitated remote monitoring for physicians and enabled researchers to collect data continuously and in real-time. A variety of wearable devices are now on the market for measuring physiological data, including blood pressure, heart rate, respiration, and glucose levels, [21]. These devices come in different forms, such as wristbands, smartwatches, and monitoring patches. Smartwatches and fitness trackers, known for their versatility, have gained popularity among consumers. Likewise, monitoring patches have emerged as a valuable asset to the healthcare sector by enabling remote treatment for patients.
    3. (c)
      Neural Sensors
      Neural sensors play a crucial role in comprehending the workings of human neurons. They enable us to decode brain signals, assess the brain’s current state, and, when necessary, optimize it for improved focus and attention. This practice is commonly referred to as Neurofeedback [22].
    4. (d)
      Environmental and Chemical Sensors
      Environmental sensors serve the purpose of measuring various environmental factors, including temperature, humidity, pressure, and pollution, among others. Traditional tools like thermometers and barometers can effectively measure parameters such as temperature and pressure. When it comes to assessing air quality, specialized sensors come into play. These sensors detect the presence of gases and particulate matter in the air [23].
      Additionally, chemical sensors play a crucial role in detecting both chemical and biochemical substances. Among the innovative technologies available are the electronic nose (e-nose) and electronic tongue (e-tongue), which rely on pattern recognition to sense chemicals based on odor and taste. These sensors find valuable applications in smart cities for monitoring pollution levels [24].
    5. (e)
      RFID
      RFID technology finds widespread use in various IoT applications, serving as sensors. For instance, it is employed for tracking products within extensive inventories or monitoring items within large retail stores.
  2. 2.
    Actuators
    Actuators hold a crucial role and operate in direct contrast to sensors. They transform energy into physical motion and are typically positioned on the outer periphery of a system. Take, for instance, a scenario involving a smart home system that incorporates numerous sensors and actuators. In this setup, the actuators receive signals from the sensors and, depending on the context, carry out actions such as locking or unlocking doors, toggling lights or electrical devices on or off, regulating the house’s temperature, or setting alarms for emergencies. Essentially, actuators respond to and execute commands based on the signals they receive from sensors or other devices.
Storage and Data Analytics
Another crucial aspect of IoT is the management of the substantial volume of data generated and exchanged by IoT devices continuously. Storing this data presents a significant challenge within IoT networks. Furthermore, the data collected from these devices must undergo filtering, processing, and analysis to enable the effective functioning of the IoT system. In this process, gateways, cloud services, and analytics collaborate to handle data storage and processing tasks.
  1. 1.
    Gateway
    The IoT gateway are designed to simplify the IoT system. The gateway functions as the intermediate medium of communication between the devices and the central cloud system. The major functionality of IoT gateways are listed below:
    1. (a)
      Data Preprocessing
      The IoT gateway acts as an intermediary between sensor devices and the central cloud. Prior to forwarding data directly to the cloud, basic data analytics are performed in this stage. This layer performs the tasks of local data filtering, cleaning, preprocessing, and protocol translation. During this process, it may also aggregate, remove duplicates, or summarize the data to enhance response times and lower transmission costs [25].
    2. (b)
      Data Acquisition
      In this layer, data is collected from multiple sources and then transferred to the processing layers. Data is collected and converted into desired format in this layer by utilizing Data acquisition systems. The role of gateway in this stage is to provide secure connectivity between IoT devices and processing structure [26].
    3. (c)
      Data forwarding and Temporary Storage
      The primary role of the gateway is to ensure secure data transfer between the sensor layer and the central cloud [27]. Additionally, this layer serves as the temporary storage repository for the collected data.
    4. (d)
      Device Management
      This layer cacilitates real-time device configuration, allowing adjustments to device statuses, operational modes, error acknowledgments, and more [27].
    5. (e)
      Diagnostics
      The IoT gateway identifies errors and faults within the entire technology layer, including self-diagnostics for the IoT gateway itself [27].
  2. 2.
    Cloud
    The cloud serves as the central hub of an IoT network, taking on pivotal roles in data processing, storage, and management. Among its various responsibilities, the following key characteristics stand out:
    1. (a)
      Data Storage and Processing
      The cloud infrastructure serves as the storage solution for the extensive data generated by IoT devices, addressing the impracticality of storing such massive volumes within individual systems. Additionally, data processing takes place within this stage.
    2. (b)
      Scalability
      Cloud infrastructures, including compute, storage, and networking resources, can effortlessly handle thousands of devices.
    3. (c)
      Flexibility
      Cloud-based IoT offers considerable flexibility, allowing devices to be added or removed as needed without requiring a complete system reconfiguration.
    4. (d)
      Maintenance
      In cloud-based IoT, network devices are supervised and managed by the cloud service provider.
    5. (e)
      Cost Effective
      Cloud usage is cost-effective as users only incur charges for the resources they actively utilize.
    While cloud services are not mandatory for IoT, the recent shift toward edge and fog computing empowers local data processing. Nevertheless, the cloud is incorporated into the system for its scalability. storage and cost-effective service provision [28]. Furthermore, cloud-based services offer security functionalities such as encryption and authentication while enabling remote access and control of IoT devices.
  3. 3.
    Analytics
    This represents one of the most intricate and vital layers within IoT. It involves the analysis of data, generating valuable insights through the application of diverse machine learning algorithms and statistical analysis techniques. Numerous applications of analytics in IoT encompass Anomaly Detection, Environmental Monitoring, Energy Management, Smart Cities, and Agriculture [29].
Interpretation and Visualization Tools
This segment essentially serves as the user interface (UI). The UI offers a platform for users to interact directly with the application or system, facilitating communication. A user interface is not always reliant on a screen. For instance, a TV remote utilizes a user interface comprising multiple buttons, while devices like the Amazon Echo respond to voice commands for control. Receiving Automatic Notification, Monitoring Information Proactively, and Controlling the System Remotely are some common examples for user interfaces in IoT systems [30].
1.2.3 Characteristics of IoT
Characteristics pertain to the distinct qualities or capabilities of a system that set it apart from others. When it comes to IoT, providing a fixed set of characteristics can be challenging due to the evolving nature of IoT devices and their dependency on the specific application domain. Nevertheless, there are several common traits that help differentiate IoT technology from other systems. This section describes the characteristics of IoT system [31, 32]. Fig. 1.1 provides a visual representation of the characteristics of IoT.
Figure 1.1: IoT characteristics
  1. 1.
    Connectivity
    Connectivity is the most vital requirement of IoT. The main aspect of Internet of Things is a network with millions of devices connected to each other. The connectivity remains constant, allowing anyone from anywhere to connect to the IoT network at any given moment.
  2. 2.
    Intelligent decision making capability
    The extraction of knowledge from the data generated is highly significant. Consider a sensor that produces data; however, the true value of that data lies in its proper interpretation. This represents a crucial aspect of IoT, wherein IoT devices possess the capability to transform raw data collected by sensors into meaningful information and make decisions based on it.
  3. 3.
    Dynamic and Self Adapting
    IoT devices should have the ability to adapt to changes in context, their surrounding environment, and the existing situation. For instance, within a surveillance system, cameras are capable of switching between day and night modes or adjusting their resolution in response to motion detection, demonstrating their adaptability.
  4. 4.
    Unique Identity
    Every IoT device should have a unique identity and unique identifier. IoT device interfaces enable users to inquire about device information, monitor their status, and remotely manage them. Having a distinct identity is essential to empower users to safeguard their devices, whether through password protection or alternative security measures [33].
  5. 5.
    Self Configuring
    IoT devices possess the capability to autonomously update their systems in response to the situation, eliminating the need for user intervention. Moreover, they exhibit flexibility in network management, allowing new devices to seamlessly join the network and permitting any device to depart from the network at any time.
  6. 6.
    Scalability
    The IoT network is experiencing a continual growth in the number of connected devices, resulting in a substantial and continuous generation of data. Consequently, scalability emerges as the foremost feature for any IoT system.
    Figure 1.2: Advantages of Internet of Things
  7. 7.
    Interoperability
    IoT devices rely on standardized protocols and technologies to guarantee seamless communication among themselves and with other systems. Interoperability represents a fundamental block of the IoT, signifying the capacity for various IoT devices and systems to interact and share data, irrespective of the underlying technology or the manufacturer. Hence, IoT devices employ standardized protocols, data formats and technologies to uphold interoperability.
  8. 8.
    Heterogeneity
    The devices within an IoT network exhibit heterogeneity, showcasing the network’s capability to accommodate diverse elements.
  9. 9.
    Energy Efficiency
    Energy efficiency stands out as a significant characteristic of IoT. Numerous IoT devices are purposefully crafted to minimize energy consumption and create low-power devices. Furthermore, various IoT approaches are tailored to optimize power usage, such as opting for fog/edge computing over cloud computing to decrease bandwidth requirements and reduce power consumption when constructing an IoT system [34].
  10. 10.
    Security
    The presence of millions of internet-connected devices and the vast data generation underscore the vulnerability of the IoT network to security threats. Hence, safety and security emerge as pivotal characteristics of IoT. Ensuring safety is of paramount importance to preserve the effectiveness of IoT’s advantages, including efficiency and novel experiences.
1.3 Advantages of IoT
The amount of benefit IoT offers to people’s life is the main reason IoT is becoming more popular by the day. IoT solutions are developed and invented to make people’s life easier and more convenient. Healthcare, education, and business—there is hardly any field that is not affected by IoT technology. This section elaborates on the most visible benefits of IoT in people’s daily lives [35, 36] and Fig. 1.2 provides a visual overview.
  1. 1.
    Efficient Data Collection
    IoT based data collection has been particularly beneficial in sectors such as healthcare and finance [37]. A straightforward illustration of efficient data collection is the integration of IoT in the retail sector. Internet-connected tags can provide data on purchase decisions and sales trends, whether weekly or monthly. This enhanced data gathering can improve inventory management and reveal valuable insights into customer behavior, ultimately contributing to the prosperity of the business.
  2. 2.
    Control and Automation
    IoT has provided its customers a more convenient lifestyle and allows them to control their daily activities with the touch of a button. One simple yet remarkable example is the smart bulb, which can be controlled without even touching the switches; the user can simply turn off or on the light remotely. Not only can the light bulb, coffee maker, or any other electrical device in the home be controlled with the tip of the device, but they can also be controlled using a voice command when they are linked to Google’s or Amazon’s voice assistants.
  3. 3.
    Real-time Access to Information
    IoT devices offer immediate access to information, proving invaluable across healthcare, business, and everyday applications. A prime illustration of the advantages of real-time data access is within the healthcare sector. Physicians can access patient data in real-time, enabling continuous health monitoring. This capability becomes particularly crucial in delivering emergency medical assistance swiftly when unexpected health issues arise.
  4. 4.
    Improved Efficiency
    IoT systems operate autonomously, a valuable asset in various domains. Reduced human intervention leads to increased efficiency and decreased labor reliance. For instance, a company with a fleet of delivery vehicles can effortlessly track their real-time locations, eliminating the need for manual employee involvement in this task.
  5. 5.
    Improved Quality of Life
    The advent of IoT has significantly improved the lives of its users in numerous ways. Real-time health monitoring, including devices like blood pressure monitors and fitness trackers, empowers users to maintain their well-being effectively. Smart homes offer a stress-free and effortless lifestyle. These advantages extend beyond individuals and can benefit entire industries or communities. Smart devices, linked not only to intelligent traffic lights but also road safety monitors and toll gates, can provide drivers with real-time information about road conditions on their route.
  6. 6.
    Cost and Time Effective
    IoT minimizes human effort and relies heavily on real-time data transmission, leading to time savings. For example, real-time patient monitoring benefits both patients and doctors by eliminating the need for physical meetings, thus saving time for both parties. IoT aids businesses in streamlining their workflows by offering valuable insights and real-time information, resulting in cost reductions. In addition to businesses, individuals can reduce their everyday expenses through the use of IoT.
  7. 7.
    Asset Tracking
    This process involves tracking products within a business or logistics management system. Manual asset tracking is labor-intensive and time-consuming, but it can be streamlined through the application of IoT technologies like barcodes and RFID tags. These technologies allow for remote monitoring of goods and provide stakeholders with information about any faults or problems in real time [38].
  8. 8.
    Data Analysis for Research Purpose
    The enormous amount of data collected from the IoT devices has become a blessing for the researchers in various fields like healthcare, education, business etc. The healthcare researchers can use the data collected via biosensors and utilize them in inventing cures and vaccines for disease, the finance industry can utilises the data to understand the trend and improve customer experience, the super shops can analyze the customer behaviour and improve their business and so on.
  9. 9.
    Big Data and Predictive Analysis
    “Big data” has been a widely recognized term in the world long before the emergence of IoT. It involves the collection and analysis of massive volumes of data. One of the primary objectives of IoT is to amass data from diverse sources, sending this information back to systems for analysis. Effective analysis of big data can yield valuable insights, spanning from stock market predictions to understanding customer behavior, and thereby enhancing the business landscape.
  10. 10.
    Increased Productivity
    IoT utilization in both industry and homes has the potential to boost productivity significantly. For instance, in a smart home, users can streamline various household tasks using voice commands, enabling efficient multitasking. Similarly, in a business setting, analyzing customer behavior can enhance customer satisfaction, ultimately contributing to the prosperity of the enterprise. As an example, 46% of businesses that embraced IoT strategies saw improvements in efficiency, even though only 29% initially anticipated such improvements [39]. In the healthcare sector, doctors can offer more extensive services to their patients if they do not need to make physical visits to each patient individually.
  11. 11.
    Safety and Security
    The incorporation of IoT offers users a means of security not only in their homes but also in their businesses, schools, offices, and virtually anywhere. Individuals can remotely monitor their valuable assets such as vehicles and other important items. Parents can even keep track of their children’s whereabouts from their workplaces, providing peace of mind. IoT allows for vehicle tracking and the setup of alert systems in case of unusual incidents. Financial companies and banks can enhance the security of their confidential rooms or vehicles by utilizing IoT [34].
  12. 12.
    Enhance Customer Engagement
    The IoT offers several avenues for enhancing customer engagement. It achieves this by leveraging valuable customer data, personalizing experiences, improving convenience, and enabling real-time interactions.
  13. 13.
    Efficient Resource Utilization
    IoT facilitates efficient resource utilization through various mechanisms and capabilities. Its ability to collect and monitor real-time information allows organizations to track the status of their resources, such as equipment and machinery, enabling the identification of inefficiencies. Furthermore, in the industrial sector, predictive maintenance powered by IoT can anticipate machinery failures, reducing downtime and optimizing resource allocation for maintenance.
  14. 14.
    Improved Technology
    The innovation brought about by IoT leads to the creation of newer and more advanced technologies in the market. For instance, consider the scenario of an air conditioner that was initially controlled manually with a remote. With the advent of IoT, users can now operate it using voice commands or control it remotely. When such innovations hit the market, they spark competition, driving the development of improved solutions based on user feedback. This cycle of innovation in response to IoT advancements contributes to the continuous progression of technology [37].
Along with the generic benefits that apply to practically all fields that are integrating IoT, various domain-specific benefits are also apparent in many sectors. Some of the domains are described below.
Healthcare
The healthcare industry is one of the most benefited by IoT among all the sectors. The innovation of IoT has revolutionized the healthcare industry by various means. Here are some common examples [6]:
  1. 1.
    Remote Medical Assistance
    The idea of remote monitoring has brought significant advantages to the healthcare sector. Doctors can now observe their patients’ conditions from their offices without the need for physical visits, just by using the sensors attached to the patients’ bodies.
  2. 2.
    Improved Treatment
    Through the utilization of IoT technology, physicians can make well-informed decisions based on evidence and ensure full transparency.
  3. 3.
    Faster Disease Diagnosis
    Remote and continuous monitoring allows doctors to identify diseases in their early stages, which can save many lives.
  4. 4.
    Emergency monitoring and reporting
    During remote monitoring, if any unusual or emergency situations arise, such as heart failure or asthma attacks, doctors can quickly provide emergency treatment to the patient.
Agriculture
The use of IoT in agriculture brings about numerous advantages. This section highlights some of the most noticeable benefits [40].
  1. 1.
    Community Farming
    The implementation of IoT can facilitate the development and advancement of community farming. Through IoT, it becomes feasible to support services that enable shared data storage, the exchange of data and information, and the enhancement of communication among farmers and agriculture experts [41]. Additionally, IoT facilities or mobile phones can be used to share various agricultural equipment within the community, either for free or through paid services.
  2. 2.
    Safety Control and Fraud Prevention
    In agriculture, the challenge extends beyond just increasing production; it also involves ensuring the production of safe and nutritious food. There have been instances of food fraud, including adulteration, counterfeiting, and artificial enhancements, which pose significant health risks and negative economic impact [42]. Researchers in [43] have explored various aspects of food fraud, including product integrity, process integrity, people integrity, and data integrity, which can be effectively addressed through IoT technology. Additionally, IoT can be integrated into both logistics traceability and the qualitative tracking of food [44].
  3. 3.
    Wealth Creation and Distributions
    The adoption of IoT will uncover novel business models, allowing individual farmers to directly connect with customers, eliminating the need for intermediaries. Consequently, this will lead to increased profits for the farmers.
  4. 4.
    Cost Reduction and Wastage
    A significant benefit of IoT is its capability for remote monitoring of devices and equipment. Farmers can oversee extensive fields from a distance, eliminating the need for physical visits and saving time. Additionally, precise knowledge of when and where to apply pesticides and fertilizers will reduce costs and minimize material wastage [45].
  5. 5.
    Operational Efficiency
    IoT technology brings operational efficiency not only to farmers but also to decision-makers in the agriculture sector, including government and non-governmental agencies. By collecting data through IoT-based agriculture surveillance, this information can guide various interventions such as disease prevention, wildfire management, compensation programs, and resource allocation.
    Farmers can leverage IoT and data analytics for making informed decisions about farm management and processes. Automated health monitoring of livestock or crops enables efficient diagnosis and prescription of medicine by veterinary or agriculture officers, reducing losses. Additionally, IoT optimizes the agri-food supply chain, ensuring real-time balance between demand and supply.
  6. 6.
    Awareness
    IoT is anticipated to create cost-effective mobile apps and wireless network services for the agriculture industry. This development can make information about product markets, prices, government services, and regulatory standards for various farm produce easily accessible. As a result, people can receive notifications when fresh products hit the market and connect with farmers interested in organic foods.
  7. 7.
    Asset Management
    IoT simplifies asset management by enabling remote monitoring of farm equipment, preventing theft, and issuing alerts for necessary maintenance.
Education
The educational sector can derive similar benefits from the utilization of IoT, much like the advantages observed in other industries such as healthcare or retail, as previously described [34]. Some common examples are presented below.
  1. 1.
    Improvement in Learning Environment
    IoT integration has the potential to revolutionize the learning experience for all students. One innovative approach involves attaching RFID tags to various objects, allowing students to interact with these objects and instantly access relevant information. This not only promotes greater student engagement but also makes learning a more enjoyable and interactive process.
  2. 2.
    Distance Learning
    IoT has the potential to enhance distance learning, enabling students facing physical classroom attendance challenges, whether due to health concerns or other reasons, to actively engage in virtual classrooms.
  3. 3.
    Enhanced Assistant for Special Students
    IoT makes learning more accessible for special students who require additional support and assistance. As referenced in Selinger et al. [46] and Lutz et al. [47], IoT is confident to offer minority and disabled learners the opportunity for high-quality education and peer-to-peer interaction. It will also facilitate student-to-teacher engagement, ultimately enhancing their prospects for greater academic success.
Warehouse
The implementation of IoT in warehouse management and processing has paved the way for smart warehousing. As a result, big industries like Alibaba, Amazon, and DHL have already integrated this technology into their inventory, logistics, and warehouse management systems. Here are descriptions of several common benefits of smart warehousing:
  1. 1.
    Efficiency and Transparency
    The primary and most essential requirement in inventory and warehouse management is to ensure efficiency and transparency across the entire process. Barreto et al. [48] have presented the key aspects of Logistics 4.0 where they propose the use of an intelligent system and a database to facilitate data sharing through an IoT platform. This interconnected approach integrates all warehouse operations and workforces, leading to increased transparency in the process and enhanced system efficiency.
  2. 2.
    Efficient Order Monitoring
    Implementation of IoT can help in monitoring the order without any human intervention. The authors of [49] has created a system that employs the following components: (1) Bluetooth tags for transmitting relevant product data; (2) a receiver for monitoring the movement of objects and the position of pallets; and (3) an integrated software package and database server for warehouse monitoring and data exchange. This developed system significantly enhances the efficiency of stock management, delivery timing, tracking, and product management.
  3. 3.
    Ensure Quick Tally
    In the traditional tally process, which relies entirely on manual labor, the task of updating product numbers and information is labor-intensive. However, through IoT integration, the tally process can be significantly expedited by simply scanning the barcodes on the products [50].
  4. 4.
    Enhance Warehouse Automation Management
    IoT technology can enable warehouse workers to inspect storage boxes without the need for physical contact, a departure from the traditional method of physically touching each box to check its contents.
Supply Chain Management
The supply chain represents a intricate logistics system comprising services that transform raw materials into finished products and subsequently deliver them to customers. Furthermore, Ensuring a continuous and on-demand supply of products is a crucial market requirement. IoT introduces groundbreaking advantages in supply chain management, some of which are detailed below [38].
  1. 1.
    Decision Making
    In the field of logistics, the frequency of information flow and the need for instant decision-making surpasses that of most other industries due to its unique nature. Sensors integrated with products facilitate the real-time and continuous generation of data, simultaneously allowing the analysis of this data to predict market trends. This predictive capability aids in making efficient decisions across the entire supply chain.
  2. 2.
    Smart Transportation
    The integration of IoT technology facilitates intelligent transportation solutions, encompassing the monitoring of delivery vehicles, tracking vehicle locations, and promoting safe driving practices. By employing sensors connected to all delivery vehicles, real-time tracking of vehicle locations becomes possible, and notifications can be dispatched in the event of accidents. Additionally, these sensors can monitor factors like excessive vehicle speed, fuel tank levels, and tire pressure to ensure safe and efficient transportation. Moreover, by analyzing traffic conditions and road conditions, the system can send notifications to the driver regarding the status of the road leading to their destination. Additionally, the system can suggest alternative routes that promote safe driving.
  3. 3.
    Smart Loading and Unloading
    Efficiently handling the loading, unloading, and stacking of products is crucial for item distribution. It is essential to achieve this with cost-efficiency and timeliness. Automation lifts equipped with RFID tags can streamline this process by precisely locating and delivering goods without requiring human intervention. Furthermore, robots can be employed as warehouse workers, leading to a reduction in human labor [51].
  4. 4.
    Smart Carriage
    Integrating Automatic Guided Vehicles (AGV) into both internal and external transportation systems can lead to cost savings and time efficiency gains [52]. Various communication technologies facilitate internal communication among these vehicles, contributing to more efficient management processes as a result [53].
  5. 5.
    Smart Packaging
    IoT has seamlessly replaced traditional packaging processes with intelligent and smart packaging solutions. These smart packages are capable of making real-time decisions, and product information is easily accessible through smart labels affixed to the items [54]. In [55], authors have employed a similar process to categorize products based on their temperature sensitivity. Furthermore, to ensure the safe delivery of products, an automatic reminder system have been proposed that activates if the products deviate from the preset conditions.
  6. 6.
    Smart Distribution
    In the logistics and warehouse management of product distribution, several stages are involved both before and during the process. Before the final goods are dispatched, there is a crucial step of labeling products according to their types and ensuring proper label attachment. This task is efficiently handled by an autonomous wheeled robot [56]. In specific industries like the food sector, smart labeling tags equipped with microcontrollers are employed to effectively track food items [57]. IoT plays a significant role in various aspects of the distribution process, including route planning, item location, sorting, and labeling.
1.4 IoT Architecture
Architecture refers to a structured framework depicting the tangible components of a network, their operational arrangement and setup, underlying principles and structures, as well as the manner in which data is organized and utilized within its functioning. IoT architecture comprises of a collection of devices, sensors, actuators, end users, cloud services and most importantly various communication layers and IoT protocols. Owing to the dynamic characteristics of the IoT domain and the crucial importance of sensors, it is impossible to identify one fundamental and universally agreed architecture for diverse IoT domains. Nevertheless, certain foundational elements will exhibit similarity across various projects.
The architecture of the IoT has been categorized based on a range of factors. Certain researchers have undertaken a division grounded in distinct layers, while others have pursued a division predicated on the specific objectives of the IoT project. However, there are some fundamental blocks present in more or less in every IoT architecture which provides scalability and flexibility in IoT architecture [58].
  1. 1.
    Sensing Layer
    The primary function of this layer involves real-time detection of changes in the connected devices. This layer comprises two main actors: sensors and actuators. Sensors are responsible for acquiring diverse data types from the environment, including but not limited to temperature, humidity, caloric information, heart rate, and object identification. Meanwhile, actuators are mechanized components tasked with device manipulation, encompassing actions such as activating and deactivating door lights and similar functionalities.
  2. 2.
    Communication Layer
    Functioning as an intermediary system bridging the sensing layer with the upper layers, this layer manages connectivity, facilitates data transmission between the sensing and cloud layers, or upper layers, and conducts message routing among remote devices.
  3. 3.
    Cloud Layer
    Situated at the core of the IoT architecture, this layer functions as the central processing unit. It receives data collected from the sensing layer to conduct storage, processing, and analysis procedures. Within this framework, the cloud employs a data centre as a pivotal server for tasks like data refinement by eliminating redundant and noisy information, data format conversion, and more. Nonetheless, the advancement of cloud computing research is progressively shifting towards the decentralisation of multiple processing tasks from the cloud to edge nodes that are currently concentrated in data centres [59].
  4. 4.
    Management Layer
    Tasked with monitoring and administering the entire IoT system, this layer handles responsibility for device configuration and firmware updates, as well as providing analytics and generating reports. These capabilities empower system administrators to comprehensively grasp the system’s dynamics and find opportunities for enhancement.
  5. 5.
    Services and Application Layer
    This layer provides value to the end user of an IoT system. This layer utilizes processed data and provide various functionalities or services.
  6. 6.
    Security Layer
    With the responsibility of ensuring security and safeguarding against unauthorized access, this layer bears the duties of authentication and authorization, intrusion detection, threat mitigation, and related responsibilities [60].
Figure 1.3: Generic architecture of IoT
The IoT architecture can be divided in two ways: (1) Layered Architecture and (2) Domain-Specific Architecture. This section presents several existing architectures for IoT. However, detailed explanations are beyond the scope of this chapter; instead, it provides an overview of a generalized architecture.
1.4.1 Generic Architecture of IoT
IoT follows a layered architecture, which refers to the structured framework used to design and organize the various components and functions of an IoT system. This architecture is composed of multiple layers, each with its specific role and responsibilities, allowing for efficient communication, data processing, and management within the IoT ecosystem. The typical layers in IoT architecture include the perception layer, network layer and application layer. These layers collaborate to ensure the smooth operation of IoT devices, data transmission, and the provision of IoT services to end users.
According to this generalized architecture, also known as the three layer architecture, the IoT system is divided into three layers [61] namely (i) application layer, (ii) network layer, and (iii) perception layer. Every one of these layers possesses inherent security challenges [62]. A visual representation of the Three Layer Architecture is featured in Fig. 1.3. The details are described below.
  1. 1.
    Perception Layer
    The perception layer, also known as the sensor layer is the foundational layer of IoT architecture [61]. This layer engages with smart devices, including but not limited to smartwatches and smart rings, employing an array of sensors and actuators. The principal objective of this layer pertains to the collection of data from these intelligent devices via sensors, subsequently transmitting the acquired data to the upper layer known as the network layer.
  2. 2.
    Network Layer
    The network layer, also known as the transmission layer is the middle layer of IoT architecture [63]. This layer is responsible for receiving the information passed from the perception layer and determine the routes to transmit the processed data to various connected IoT devices and applications using integrated network such as wired or wireless secure connection. Network layer is core layer of IoT three layer architecture as this uses various devices such as routing devices, gateways, switches and hub and operate them by using various communication technologies such as WiFi, Bluetooth, 3G, LTE, Zigbee etc. In summary, the network layer is responsible for transmitting data to and from several applications through interfaces and gateways using multiple communication technologies and protocols.
  3. 3.
    Application Layer
    Serving as the uppermost tier within the IoT architecture, denoted as the application layer or business layer, as referenced in [58], this layer is tasked with the aggregation of data from the network layer, thereby striving to attain the objective of establishing an smart environment, the ultimate aim of the IoT paradigm. This layer accommodates a diverse array of applications, each characterised by their own requirements. Examples of such applications are smart grids, smart cities, and smart transportation, as elaborated in [62]. Moreover, this layer assumes the responsibility of upholding the data’s authenticity, integrity, and confidentiality, as elaborated in [64].
Figure 1.4: Visual representation of Fog vs Cloud
The three layer architecture is the generalized and most common architecture and several systems have integrated this architecture [64]. Although this multi layer architecture seems to be simple at first glance, but the functionalities of the network and application layer might get complex at times. For example, the network layer is not only responsible for data transmission but also provides data services such as data aggregation and processing etc. On the other hand, the application layer is not solely responsible for providing service to customers and users but also provides data analytic and conducts data mining etc. Therefore, in response to specific requirements, additional layers have been incorporated, building upon the fundamental layers. For instance, the four-layered or the five-layered architectures enhance the system’s flexibility. Nonetheless, the three-layer architecture serves as the foundation for all these variations. Furthermore, new-generation applications require shorter response times and low energy consumption as IoT devices have limited capacity [65, 66]. Therefore, researchers have utilized fog and cloud layers which are visualized in Fig. 1.4.
1.5 Key Technologies of IoT
To function properly, the Internet of Things requires a variety of technologies. Different technologies are deployed in various layers of an IoT architecture, and there are many different sorts of technologies, including hardware technology, software technology, and, most crucially, communication technology. This section discusses the many IoT technologies that are used to ensure the successful operation of an IoT system. Fig. 1.5 depicts the taxonomoy of IoT technologies.
Figure 1.5: Technology Taxonomy of IoT
1.5.1 Hardware Platforms
The major components of an IoT system are devices attached with sensors or wearable devices which are used to data collection. Therefore, various types of hardware platforms are used in order to build these sensor devices. Several key points need to be considered before selecting the hardware platforms such as what is the purpose of the IoT device, what kind of connectivity it requires and so on.
Two most accessible and popular hardware platforms are Raspberry Pi and Arduino. Raspberry Pi is a Linux based single board computer whereas Arduino makes different microcontrollers, open-source development tools, kits, and software. Both of them have strong data acquisition, processing ans storage capabilities and and provide both wireless and wired connectivity. However, in terms of power management Arduino is superior than Raspberry Pi because Raspberry Pi does not contain sleep or suspend modes for power utilization but Arduino does [67]. Intel has developed several development boards named Intel Galileo Gen and Intel Edison which uses Arduino IDE as development environment. Beagle Bone Black is a small sized Linux based board designed to build powerful IoT devices. Other than these platforms there are several other hardware platforms present in the markets such as Electric Imp 003, ARM mbed NXP LPC1768, WiPy and LoPy boards by Pycom company and many more [32].
1.5.2 Wireless Communication Technology
Given the vast number of devices in the current IoT network and the expectation of diverse connected devices in future IoT applications, there’s a urge to develop various technologies to facilitate their connectivity. This subsection explores the existing wireless technologies designed for IoT connectivity and categorize them into three groups.
Table 1.1: Comparison between popular short-range technologies based on various parameters. (Abbreviations: CSMA/CA (Carrier Sense Multiple Access with Collision Avoidance), CSMA/CD (Carrier Sense Multiple Access with Collision Detection), FHSS (Frequency Hopping Spread Spectrum), O-QPSK (Offset Quadrature Phase Shift Keying), OOK (On-Off Keying), PPM (Pulse Position Modulation), LBT (Listen Before Talk)
OWC
Parameters Bluetooth ZigBee LR-WPAN Wi-Fi VLC BS-ILC
Standard IEEE 802.15.1 IEEE 802.15.4 IEEE 802.15.4
IEEE 802.11a/b/c/
d/g/n/ac/ah
IEEE 802.15.7m
IEEE 802.15.13
LoRaWAN
Frequency Band 1MHz - 2.48GHz
Mainly at 2.4 GHz
Optionally 868MHz
or 915MHz
868/915 MHz,
2.4 GHz
a: 5GHz, b: 2.4GHz,
g: 2.4GHz, n: 2.4GHz
802.11ah: 1/2/16MHz
400-800 THz
Varies by region,
Europe: 868 MHz
USA: 915 MHz
Data Rate 1Mbps - 3 Mbps 20kbps to 250kbps 40-250 kbps
a: 54Mbps, b: 11Mbps,
g: 54Mbps, ah: 300Mbps
n: 600Mbps, ac: 7Gbps
15.13: multi gigabit
Recent: 100 Gbps
100Gbps
Transmission range
Classic: 100m
BLE: 240m
10-100 meters 10-100 meters 100m to 1km
Typically, within a room
7m: 200 meters
15.13: Several meters
200 meters
Energy consumtion
Classic: High
BLE: Low
Low Low Moderate to high
Transmitters: Moderate
Receivers:Minimal
Very low
Cost Cost-effective cost-effective Cost-effective moderate to high moderate to high Cost-effective
RA protocol
FHSS, FDMA
TDMA based polling
CSMA/CA CSMA/CA CSMA/CD, CSMA/CA
CSMA/CA,
TDMA/CDMA
LBT
Modulation type
GFSK, DQPSK,
$\pi/4$-DQPSK
BPSK/O-QPSK BPSK/O-QPSK BPSK/QPSK/QAM OOK/PPM/OFDM CSS
Short-Range Technologies
Short-range technologies are commonly used in IoT applications to enable communication between devices within a limited proximity. These technologies are well-suited for scenarios where devices need to exchange data within a small coverage area, typically spanning from a few meters to a few kilometers. There are several short-range technologies, each with unique characteristics and advantages suited for specific purposes. Table 1.1 provides a comprehensive overview of various short-range technologies used in the IoT environment, categorized based on various parameters such as frequency band, data rate, transmission range, and more. The table summarizes the technical specifications of Bluetooth, ZigBee, Wi-Fi, LR-WPAN, VLC, and BS-ILC [68, 32].
Bluetooth, ZigBee, LR-WPAN, and BS-ILC offer cost-effective communication solutions, whereas the overall cost of Wi-Fi and VLC may vary depending on usage. In terms of frequency bands, most technologies operate around 2.4 GHz, except for VLC, which utilizes a much larger frequency band. The frequency for BS-ILC varies by region. VLC and BS-ILC prioritize achieving high transmission rates of approximately 100 Gbps, while Wi-Fi transmission rates can vary across different standards. Notably, VLC and BS-ILC exhibit very low energy consumption, as indicated in the table.
Long Range Technologies
Long range technology in contrast with short range technologies refers to technologies that enable communication over extended distances. These technologies are designed to transmit data or signals over considerable ranges, typically much farther than short-range technologies like Bluetooth or Wi-Fi. Long-range technologies are commonly used in various applications such as long-distance wireless communication, remote monitoring, tracking, and more. They are essential for scenarios where data needs to be transmitted reliably over extensive geographical areas.
Long-range wireless technologies can be classified into two main categories: Mobile Communication Technology and LPWAN Technologies. Table 1.2 presents a detailed overview of various long-range technologies utilized in the IoT environment, organized based on several parameters such as frequency band, data rate, transmission range, energy consumption, cost, and more. it provides a summary of the technical specifications for various versions of Mobile Communication Technologies, which include 2G, 3G, 4G, 5G, and WiMax, as well as different versions of LPWAN Technologies, which encompass UNLICENSED LPWAN and LICENSED LPWAN.
It is evident from the table that when it comes to mobile communication technologies, 2G, 3G, and 4G operate within a frequency band ranging from approximately 850MHz to 2100MHz, while 5G utilizes a much broader frequency band of up to 80GHz. Additionally, all five technological variations, 5G offers the highest data rate, exceeding 20Gbps. In terms of energy efficiency, both 2G and 5G have low consumption rates, whereas WiMax exhibits high energy consumption. However, it’s worth noting that 3G, 4G, and 5G all come with higher implementation costs. In the UNLICENSED LPWAN category, the top contenders are LoRa and Sigfox, while the most promising options in the LICENSED LPWAN technology are LTE-M and NB-IoT (Narrowband Internet of Things). The key distinction highlighted in the table is that UNLICENSED LPWAN technology operates within unlicensed spectrum resources, leading to lower deployment costs, whereas LICENSED LPWAN technology relies on licensed spectrum resources, resulting in relatively higher expenses for both devices and deployment [68].
Emerging For Massive Connectivity
While current wireless IoT technologies have achieved some success in supporting various IoT applications, they still face challenges in meeting the future demands of IoT, where there will be a need to connect hundreds of billions of objects or things. For example, handling the connectivity among vast number of IoT devices with limited transmission payload and several random access protocols used by existing technologies often result in significant problems such as frequent access collisions, increased delays, and a high amount of signaling overhead for IoT devices. Furthermore, the constrained availability of wireless resources for connecting IoT devices creates shortages and inefficient utilization of these resources. Hence, there have been numerous ongoing initiatives to tackle the limitations of current technologies. Among these, the solutions CS, NOMA, mMIMO, and ML stand out as the most promising [68].
These technologies all have the potential to be incorporated into future IoT connectivity standards. However, it is essential to acknowledge that there are existing issues and constraints that must be resolved for their effective implementation. These emerging technologies offer the capability not only to support massive connectivity but also to deliver high reliability and low-latency transmissions. In the future, it is anticipated that the development of even more advanced technologies aimed at addressing critical IoT challenges. Simultaneously, efforts should focus on intelligently integrating existing and emerging technologies to unlock their full potential and optimize system performance [68].
Table 1.2: Comparison between popular long-range technologies based on various parameters. (Abbreviations: GSM (Global System for Mobile Communications), UMTS (Universal Mobile Telecommunications System), LTE-A (Long-Term Evolution Advanced), LTE-M (Long-Term Evolution for Machines), 3GPP (3rd Generation Partnership Project), TDMA (Time Division Multiple Access), CDMA (Code Division Multiple Access), FDMA ( Frequency Division Multiple Access), WCDMA (Wideband Code Division Multiple Access), OFDMA (Orthogonal Frequency Division Multiple Access), SC-FDMA (Single Carrier Frequency Division Multiple Access), CP-OFDM (Cyclic Prefix Orthogonal Frequency Division Multiplexing), TDD (Time Division Duplex), FDD (Frequency Division Duplex), ALHOA (Adaptive Link Hopping On Air), QPSK (Quadrature Phase Shift Keying), BPSK (Binary Phase Shift Keying), QAM (Quadrature Amplitude Modulation), GFSK (Gaussian Frequency Shift Keying), DBPSK (Differential Binary Phase Shift Keying), CSS (Chirp Spread Spectrum), OFDM (Orthogonal Frequency Division Multiplexing), GMSK (Gaussian Minimum Shift Keying))
Mobile Communication Technology LPWAN TECHNOLOGIES
UNLICENSED LPWAN LICENSED LPWAN
Parameters Technologies 2G 3G 4G 5G WiMax LoRa Sigfox LTE-M NB-IoT
Standard GSM, CDMA
UMTS,
CDMA2000
LTE, LTE-A,
IEEE 802.16
5G NR,
Wi-Fi 6
IEEE 802.16 LoRaWAN R1.0
Proprietary
technology
3GPP LTE 3GPP LTE
Frequency band
850MHz,
900MHz,
1800MHz,
1900MHz
850MHz,
1900MHz,
2100MHz
700MHz,
1700/2100MHz,
2500MHz
60MHz -
80GHz
2.3GHz,
2.5GHz,
3.5GHz
Unlicensed ISM
bands, 125kHz,
250kHz
Unlicensed ISM
bands, 100Hz
Licensed LTE
bands, 1.4MHz
Licensed LTE
bands, 200kHz
Data rate
9Kbps -
384Kbps
384Kbps -
several Mbps
100Mbps -
1Gbps
1Gbps -
20Gbps or
higher
16d: 75Mbps
16e: 1Gbps
0.3Kbps -
50Kbps
100bps -
600bps
300bps -
1Mbps
200bps -
250kbps
Transmission range
several
kilometers
several
kilometers
several
kilometers
several
kilometers
several
kilometers
Urban: 5km
Rural: 20km
Urban: 10km
Rural: 50km
Urban: 1km
Rural: 10km
several
kilometers
Energy consumtion
low Moderate
Moderate
to high
Low (Energy-
efficient)
high Extremely low Low Moderate Low
RA protocol
TDMA/
CDMA/
FDMA
CDMA/
WCDMA
OFDMA/
SC-FDMA
Massive MIMO/
beamforming/
CP-OFDM
OFDMA/
TDD/FDD
ALOHA/
Slotted-ALOHA
ALOHA Slotted-ALOHA Slotted-ALOHA
Modulation type GMSK/QPSK QPSK/16-QAM
OFDM/QPSK/
16-QAM/64-QAM/
256-QAM
256-QAM
OFDM, QPSK,
16-QAM,
64-QAM
CSS GFSK/DBPSK
QPSK/QAM/
BPSK
QPSK/BPSK
Cost Cost-effective Moderate high high high Cost-effective
Depends on
subscription
model
Moderae to high Moderae to high
1.5.3 Cloud Solutions
IoT cloud solutions offer various services like real-time data collection, data transmission, monitoring, data analytics, improved decision-making, and device management. These services are available on a pay-as-you-use basis, allowing users to pay only for the services they actually use. Cloud services are becoming increasingly popular, and cloud-based architectures are discussed in detail in Section 1.4. Cloud platforms can be integrated into numerous domains, including healthcare, smart cities, agriculture, education, and supply chain management. For simplicity purpose, Table 1.3 presents several popular platforms in the agriculture domain [32].
Table 1.3: Popular IoT cloud solutions in agriculture domain
IoT Cloud
platform
Provider
Real-time
Data Capture
Data
Visualization
Cloud
Service Type
Data
Analytics
Developer
Cost
ThingSpeak MathWorks Yes Yes (Matlab) Public Yes Free
Plotly
Plotly
Technologies Inc.
Yes Yes Public Yes Free
Carriots Altair Yes Yes Private No
Limited up to:
10 devices
Exosite Exosite Yes Yes IoTSaaS Yes 2 devices
GroveStreams
GroveStreams
LLC
Yes Yes Private Yes
Limited up to:
20 stream,
10K transaction,
5 SMS, 500 Email
ThingWorx PTC Yes Yes Private Yes Pay per use
Nimbits Open-source Yes Yes Hybrid No Free
1.5.4 Hardware and Software Technology
This category encompasses a range of hardware and software technologies responsible for different aspects of IoT communication channels. Several technologies, such as NFC and M2M, are considered as both hardware and software technologies, whereas RFID is classified as a hardware-only technology. The subsequent sections will provide brief insights into these technologies.
Hardware Technology
RFID
RFID, or Radio-Frequency Identification which has also been introduced in sub section 1.2.2, is a technology that comprises of one or more readers and several RFID tags. These tags are small microchips, have unique codes and can be on items like products in a store or even access cards. When a device called an RFID reader sends out electromagnetic radio waves, the tags respond with their codes. The reader captures these codes and sends them to a computer, which can then figure out what the tags belong to. This technology is used in various ways, like keeping track of inventory in stores, allowing access to secure areas, or monitoring packages as they move through the delivery process. Hitachi has developed a tag with dimensions that comes in three configurations Passive Reader Active Tag (PRAT), Active Reader Passive Tags (ARPT) and Active Reader Active Tag (ARAT). An Electronic Product Code (EPC) is a typical piece of information stored in an RFID tag. These codes make it possible to uniquely track objects because they are encoded on RFID tags [69].
Near Field Communication (NFC)
Near Field Communication (NFC) is a short range wireless connectivity technology that uses magnetic field induction to establish communicating between devices when they are brought within near proximity of each other without the need of prior connection establishment. It is an extension of RFID technology with additional and more dynamic features by integrating smartphones. Most modern phones are equipped with NFC chips and support applications like Apple Pay and Google Pay. NFC operates in the unlicensed Radio Frequency band at 13.56MHz. Its typical range is about 20 meters, with the actual distance often determined by the size of the antenna in the device. NFC technology is expected to play a vital role in the future of IoT by providing a wireless connection tool to link with other smart objects. For instance, using NFC on a mobile device, a user can turn their phone into various objects, like using it as a credit card for transactions [69].
Machine to Machine (M2M)
Machine-to-Machine (M2M) communication is gaining popularity and involves direct communication between computers, embedded processors, smart sensors, actuators, and mobile devices. It comprises four main components: sensing, heterogeneous access, information processing, and applications & processing. In practical terms, M2M functions within a five-part framework:
  1. 1.
    M2M Device: This is a device capable of responding to requests for data stored within it.
  2. 2.
    M2M Area Network (Device Domain): It provides the connectivity that links M2M Devices with M2M Gateways.
  3. 3.
    M2M Gateway: These gateways utilize M2M capabilities to ensure that M2M Devices can interact and connect to the communication network.
  4. 4.
    M2M Communication Networks (Network Domain): These networks facilitate communication between the M2M Gateway(s) and M2M applications.
  5. 5.
    M2M Applications: This layer serves as middleware, where data passes through various application services and is utilized by specific business processing engines.
Machine-to-Machine (M2M) technology finds applications across various sectors, including healthcare, smart robotics, cyber transportation systems (CTS), manufacturing systems, smart home technologies, and smart grids [69].
Vehicle-to-Vehicle Communications (V2V)
Vehicle-to-Vehicle Communications (V2V) treats each vehicle as a node and enables wireless data exchange between vehicles regarding their speed, location, and more. This network allows vehicles to broadcast and receive messages in all directions. Vehicles equipped with appropriate software, often referred to as safety applications, can use messages from nearby vehicles to detect potential collision risks in real-time. There are two types of communication within this network: one is between vehicles, and the other involves road infrastructure. However, the structure or arrangement of this communication is flexible as vehicles move from one location to another. This network can be divided into four main categories: safety and collision avoidance, traffic infrastructure management, vehicle telematics, and entertainment services with Internet connectivity [69].
1.6 Applications of IoT
IoT has been effectively integrated into various domains, leading to the development of intelligent applications. While some of these applications are already available, others are still in the research phase. Nevertheless, the essence of these applications indicates that IoT is set to enhance people’s lives by providing convenience and flexibility. The following section has categorized and provided brief descriptions of the areas where IoT integration has been implemented or proposed.
1.6.1 Smart Cities
The concept of smart cities can be seen as a complex IoT paradigm, where the management of public affairs incorporates the introduction of Information and Communication Technology (ICT) solutions [70]. A smart city utilizes public resources to digitize the city and enhance the quality of life. A real-world example of a smart city implementation is “Padova Smart City”, which has been put into practice in the city of Padova, Italy [71]. Barcelona and Stockholm are two noteworthy examples of smart cities. Barcelona has embarked on the CityOS project, with the primary goal of creating a centralized operating system to manage all the smart devices and services available within the city. Their focus has predominantly centered on enhancing smart transportation and water systems. Similarly, Stockholm has also placed significant emphasis on these two domains and holds the distinction of being one of the pioneering cities to implement the concept of congestion charges, wherein users are imposed fees for entering congested areas [4].
Smart cities are characterized by technology-based infrastructure and highly efficient smart public services. These smart cities are divided into multiple subsystems, and some of them are described below. In addition, Table 1.4 summarises related research works in this domain.
Table 1.4: Relevant applications in different Smart Cities domain
Applications Related Works
Structural health of buildings [72, 70]
Waste management [70]
Air quality and noise monitoring [73, 74]
Smart Transport & Traffic congestion
[4, 75, 76, 77, 78]
Smart grid [79, 80]
Smart water systems [81, 4]
Structural Health of Buildings
This service involves continuous monitoring of areas prone to various external factors, and measuring the health of any building. IoT sensors connected to the buildings can store information about the building’s strength, which will help analyze how sturdy the building is or if it requires any refinement [72]. Additionally, various kinds of sensors can be employed depending on the usage, such as sensors for tracking vibrations to measure the stress on the building, temperature and humidity sensors, or other types of atmospheric sensors to assess the level of pollution in an area [70]. Employing IoT in this area can reduce manual labor, where humans have to manually assess the building’s health and environmental conditions, and it can also reduce the overall cost to a great extent.
Waste Management
Waste management is a crucial operation in any city, whether it is considered a smart city or not, as it directly impacts the livability of the area. Thus, an efficient waste management system is essential for any society. Integrating IoT into waste management offers multiple benefits. It enables the detection of waste levels and real-time tracking of garbage truck routes, leading to more efficient route planning. Moreover, it can streamline manual labor involved in waste separation and monitor the disposal process. These tasks are achieved using sensors on garbage vehicles connected to a central software system that analyzes and controls the system based on the data collected by the sensors [70]. This approach reduces the cost of manual processes and enhances recycling management.
Air Quality and Noise Monitoring
Creating a healthy and safe environment for all living beings is crucial, and monitoring air quality and noise levels in any area is a key part of achieving this goal. Various environmental sensors, including soil sensors, temperature and humidity sensors, and gas sensors, can detect the presence of toxic and pollutant substances in the air and assess pollution levels. This information enables local authorities to control pollution, implement effective measures to reduce pollutant levels, identify highly polluted or toxic areas, and designate suitable locations for outdoor activities with good air quality [73, 74]. Likewise, it is essential to maintain a balanced noise level for all living beings in society, including humans and animals. Noise sensors can be used to measure the decibel levels, and the data collected can help the central authority identify noisy areas and regulate noise levels to keep them within acceptable limits.
Smart Transport & Traffic Congestion
In today’s world, traffic congestion is a widespread issue affecting nearly every country and city, particularly as more than half of the world’s population now resides in urban areas [4]. To tackle this issue, many cities have adopted IoT solutions to establish smart transportation systems aimed at managing traffic congestion.These initiatives include smart traffic lights and smart parking systems, which collectively enhance transportation capacity and improve safety and speed for travelers.
The primary advantages of smart transportation systems are reducing traffic congestion, ensuring hassle-free travel, and facilitating easy parking. Moreover, these systems enable quicker responses in case of accidents and contribute to accident reduction by effectively managing traffic flow [4]. A variety of sensor technologies are employed to achieve these objectives, such as accelerometers to measure speed, RFIDs for vehicle identification, GPS sensors for location tracking, gyroscopes for direction detection, and cameras for recording traffic patterns and vehicle movements. Practical implementations of these sensors are elaborated upon below.
  1. 1.
    Traffic Management and Monitoring Applications
    In a smart transportation system, vehicles are seamlessly interconnected via wireless networks. Each vehicle is equipped with essential IoT sensors, such as RFID and GPS. Traffic congestion is identified through the use of GPS and accelerometers, which are capable of tracking vehicle movements across different areas. The collected data is then analyzed by mapping systems to pinpoint congested areas. This information is subsequently transferred to the driver, providing guidance on selecting the optimal route and avoiding potential traffic jams [4, 75].
  2. 2.
    Application to Ensure Safety
    Numerous IoT applications have been developed with a focus on monitoring driver behavior continuously to enhance road safety. This is achieved by detecting signs of drowsiness or fatigue in drivers and providing appropriate recommendations, such as suggesting rest breaks. Various IoT applications utilize techniques like eye movement detection, facial expression analysis, and pressure sensing. These applications are typically integrated into smartphones, equipped with sensors like gyroscopes, accelerometers, and GPS, which analyze the collected data to make informed decisions. Additionally, through remote vehicle monitoring, it becomes feasible to track factors such as speed and tire pressure. This information allows centralized systems to make informed decisions aimed at preventing potential accidents [76].
  3. 3.
    Parking Guidance and Information System
    To address vehicle parking challenges, a range of IoT-based applications have been introduced. The primary goal of these applications is to assist drivers in locating available parking spaces remotely, eliminating the need to spend valuable time searching for vacant spots. This, in turn, contributes to reducing traffic congestion [77]. Parking guidance and information (PGI) applications employ barcodes to track the number of vehicles entering or leaving a parking area. Furthermore, parking zones are equipped with cameras and sensors that detect and monitor vehicles, facilitating the identification of vacant and suitable parking spots for incoming vehicles [78].
  4. 4.
    Smart Traffic Lights
    Smart traffic lights utilize data collected from sensors to make decisions based on the degree of traffic congestion. These intelligent systems can serve various purposes, such as prioritizing public transportation vehicles or implementing special controls for emergency service vehicles when needed [4].
  5. 5.
    Applications for Detecting Accidents
    These applications have the capability to identify unusual situations, such as road accidents, through the use of accelerometer sensors or the assistance of acoustic information. Additionally, they can transmit images of the surrounding area to the central authority. Based on this information, the authority can make informed decisions regarding the extent of medical assistance required.
Smart grid
A traditional grid system is an electrical grid that includes transmission lines, transformers, and various communication utilities, responsible for delivering electricity from power plants to homes or businesses. One significant limitation of the traditional grid is its one-way communication, which prevents power plants from efficiently responding to increasing power demands from consumers. To address this challenge, the smart grid establishes a two-way communication system between utilities and consumers which enables more effective management of economic, sustainable, and secure power resources [79].
By integrating IoT into grid systems, houses and businesses can be equipped with smart meters that monitor energy generation, storage, and consumption, transmitting this data to the smart grid. The smart grid, which allows for two-way communication between utilities and consumers, helps consumers in better understanding the power consumption, reports real-time increases in energy demand and calculates electricity pricing for customers, departing from the traditional monthly billing cycle [80]. Additionally, with the aid of fog/edge computing, utility providers can efficiently dispatch energy from the grid, enabling consumers to reduce their energy consumption, utilize resources efficiently, and ultimately reduce their electricity costs.
Smart Water Systems
Water is one of the most critical natural resources, and its scarcity is a prevalent issue in many parts of the world. Therefore, implementing smart water systems is not a luxury but a necessity. The primary role of smart water systems is to monitor, measure, and efficiently distribute water usage. One notable model in this field is the smart water meter designed by Hauber-Davidson and Idris [81]. These meters can detect water inflow and outflow and identify any potential leaks. Smart water meters can also be used in conjunction with data from smart river sensors and weather information to aid in flood prediction [4].
1.6.2 Medical and Healthcare
The healthcare sector has experienced remarkable advancements through the integration of IoT, offering solutions to real-life healthcare challenges and enhancing people’s lifestyles. Researchers have proposed various applications in healthcare that utilize wearable sensor devices to monitor patients’ health, diagnose diseases, issue emergency alerts, and notify users when necessary. Remote monitoring saves time for both patients and doctors while reducing overall healthcare costs. Furthermore, sharing data collected from these wearable devices with healthcare researchers contributes to the development of safer and more timely healthcare solutions and aids in the discovery of cures and vaccines for emerging diseases. This section explores several IoT applications in healthcare and a summary of the associated research papers is presented in Table 1.5.
Table 1.5: Relevant works in medical and healthcare applications
Applications Related Works
ECG Monitoring [82, 83]
Glucose level monitoring [84, 85]
Temperature monitoring [86, 87]
Blood pressure monitoring [88, 89]
Oxygen saturation monitoring [90, 91]
Mood monitoring [92, 93]
Medication management [94, 95]
Wheelchair management [96, 97]
Rehabilitation system [98, 99]
Fitness [100, 101]
Other notable applications
[102, 103, 104, 105],
[106, 107, 108, 109]
ECG Monitoring
An Electrocardiogram (ECG) measures the heart’s electrical signals and serves as an indicator of heart health, aiding in the detection of conditions like arrhythmia, prolonged QT interval, and myocardial ischemia. IoT integration in ECG monitoring has enabled early detection of heart abnormalities.
Wu et al. [82] have introduced an ECG data monitoring system utilizing a bipotential chip attached to the user’s t-shirt, transmitting data to a smartphone via Bluetooth. Real-time ECG data monitoring becomes feasible when IoT systems are combined with big data analytics. Bansals and Gandhi [83] have presented a scheme that amalgamated nanoelectronics, IoT, and big data analytics for long-term and continuous ECG data monitoring.
Glucose Level Monitoring
Diabetes is a medical condition characterized by higher blood glucose levels than those found in individuals without diabetes. Various types of diabetes exist, and doctors use three primary methods to identify them: random plasma glucose tests, fasting plasma glucose tests, and oral glucose tolerance tests. Nevertheless, the fingerstick method, involving a small pinprick to the fingertip followed by blood glucose level measurement, remains the most commonly used diagnostic approach.
The advent of IoT technology has brought about improvements in this process, making it quicker and more convenient for patients. Istepanian et al. [84] have introduced an IoT-integrated noninvasive blood glucose monitoring device to continuously monitor glucose levels, eliminating the need for fingerstick testing. It is worth noting that optical sensors, such as infrared LED and near-infrared photodiode setups, have also been employed for glucose level measurements. In these approaches, the reflected light signal from the human body is used to calculate the glucose level [85].
Temperature Monitoring
Monitoring body temperature is crucial for maintaining homeostasis and plays a pivotal role in various diagnostic processes. Traditionally, temperature measurement methods involve using thermometers placed in the mouth, ear, or rectum, but these methods often cause discomfort for the patient and pose an elevated risk of infection. However, recent advancements in IoT-based temperature monitoring applications have effectively addressed these issues.
For instance, one innovative approach by [86] where authors have introduced a 3D printed wearable device designed to be inserted into the ear. This device utilizes an infrared sensor to measure temperature from the tympanic membrane, ensuring accuracy while remaining unaffected by external factors like environmental conditions or physical activities. In another significant development, Gunawan [87] has created an IoT-based temperature monitoring system employing both Arduino and Raspberry Pi. This system records temperature data into a database, which users can conveniently access through a web page via their mobile phones or computers.
Blood Pressure Monitoring
In many diagnostic processes, measuring blood pressure is a compulsory step. However, the major issue with the traditional method is that it requires one person to record the blood pressure. Therefore, the integration of IoT into blood pressure monitoring has been a blessing for both doctors and patients. In [88], a wearable cuffless gadget has been proposed, capable of measuring both systolic and diastolic pressure, with the results stored in the cloud. A cloud and fog-based architecture have been proposed in [89] for long-term, real-time blood pressure monitoring, with the data also being stored for future references.
Oxygen Saturation Monitoring
Pulse oximetry, a highly beneficial noninvasive device for measuring oxygen saturation, addresses the limitations of traditional methods and allows for real-time monitoring. The integration of IoT-based technology has led to significant advancements in pulse oximetry, particularly in the healthcare industry. In the he study [90], an advanced noninvasive pulse oximetry system has been proposed, capable of measuring oxygen levels, heart rate, and pulse parameters, while transmitting this data to a central server. Another study [91], introduced an integrated alert system that sends notifications when oxygen levels reach a critical threshold. This system incorporates a pulse oximeter and WLAN router connected through the Blynk server.
Mood Monitoring
Monitoring a person’s mood is crucial for understanding their mental state and promoting a mentally healthy life. It also aids healthcare specialists and psychologists in treating various mental health conditions such as depression, stress, and bipolar disorder. Self-monitoring emotional stability empowers individuals to proactively address their mental well-being.
The integration of IoT into this domain offers numerous advantages. Continuous remote monitoring becomes possible, greatly benefiting individuals with severe mental illnesses by enabling real-time responses to emergencies. Furthermore, IoT can detect a person’s mental state by analyzing heartbeats through wearable devices. Kaur et al. [92] have proposed a wearable device capable of tracking a driver’s emotions, including anger, stress, terror, and sadness. The intelligent system, by analyzing emotion variations, determines whether the driver has entered a subconscious state and stops the vehicle’s DC motor accordingly. In another study, machine learning and IoT were combined to monitor a user’s mood by analyzing their social networking activities and heart rate through a wearable device [93].
Medication Management
Adhering to a prescribed medication schedule is crucial for maintaining one’s health. However, this can be particularly challenging for elderly individuals who may experience memory issues, dementia, or cognitive decline. Managing daily medication routines becomes increasingly difficult for them. Fortunately, the integration of IoT offers a solution to this problem, and numerous research efforts have explored using IoT to track patients’ medication compliance.
In one study [94], a medical box was developed with the primary function of reminding individuals to take their medications. Notably, this system features three separate trays for different times of the day (morning, afternoon, and night). Additionally, it can measure vital health parameters such as blood pressure, temperature, blood oxygen levels, and ECG readings. A mobile application was designed to facilitate two-way communication between patients and doctors, and the recorded information is accessible to both parties. Furthermore, authors in [95] have proposed storage conditions, including temperature and humidity solutions, to help users maintain an optimal environment for their medications.
Wheelchair Management
Wheelchairs are vital tools in the lives of individuals who are physically unable to move independently, providing both physiological assistance and psychological support. However, for individuals with brain damage who lack the capability to operate a wheelchair, researchers have been exploring the addition of navigation and tracking systems, with IoT playing a crucial role in these wheelchair advancements.
In one study [96], an IoT-based steering and real-time obstacle detection system for wheelchairs has been proposed. This system can identify obstacles through image processing methods, enhancing the wheelchair’s safety and usability. An advanced and automated smart wheelchair was reported in another study [97]. This wheelchair not only monitored movement but also offered features such as an umbrella, foot mat, head mat, and obstacle detection. These innovations significantly improved interaction with the living environment and enhanced the user’s overall experience.
Rehabilitation System
Rehabilitation is a critical aspect of restoring an individual’s functional abilities, aiming to help those with diseases or disabilities regain their normal lifestyle. The application of IoT in this field is versatile and has proven effective in various areas, including cancer treatment, sports injury recovery, stroke rehabilitation, and addressing physical disabilities.
For instance, an innovative smart walker was introduced in a particular study [98]. This walker utilizes a multi-modal sensor to monitor the patient’s walking pattern, and the collected data can be accessed by doctors and caregivers through a mobile application, facilitating better monitoring and support. A sports rehabilitation system was reported which monitors various parameters such as temperature, motion posture, electromyography (EMG), electrocardiography (ECG), among others, and provides real-time feedback to athletes [99]. These examples highlight the potential of IoT integration in modernizing rehabilitation systems, making them more efficient and effective.
Fitness
The quality of an individual’s life is significantly influenced by their level of fitness and daily physical activity. To facilitate fitness monitoring and promote healthier lifestyles, various applications leveraging the IoT have been developed. One approach involves utilizing smartphone accelerometer data to assess users’ activity levels. This data can reveal metrics such as the number of steps taken, the duration of physical activity, and periods of inactivity. Wearable fitness trackers are readily available in the market and have gained popularity as convenient devices for monitoring fitness levels. Additionally, IoT sensors can be integrated with gym equipment to provide real-time monitoring of individuals’ fitness progress. For instance, in a notable study [100], a “smart mat” have been introduced. This mat is designed to measure the number of exercise steps performed on its surface, offering valuable insights into users’ workout routines.
In the field of sports science, fitness assessment and training load monitoring have proven to be invaluable tools for both coaches and athletes. IoT technology enables the continuous and non-invasive monitoring of physiological biomarkers, such as saliva or sweat. This data can be leveraged to optimize hydration strategies, ultimately enhancing athletes’ overall performance [101]. These IoT-driven fitness monitoring applications and devices contribute to a more proactive approach to health and fitness, promoting active lifestyles and well-being.
Other Notable Applications
The application of IoT in the healthcare industry is incredibly diverse, extending far beyond the previously mentioned areas. There are numerous domains where IoT has already been implemented and where its potential benefits are being realized, leading to a significant increase in the adoption of Healthcare IoT (HIoT) technology.
For instance, in cancer treatment, IoT-based methods have emerged as powerful tools. An innovative IoT-based cancer treatment approach is introduced in a recent study, that encompasses various stages, including chemotherapy and radiotherapy [102]. Additionally, this system securely stores lab test results on a cloud server, allowing physicians to monitor medication dosages and enabling remote consultations through a dedicated mobile application. Furthermore, HIoT has found applications in detecting skin lesions [103], with notable advancements in lung cancer detection through state-of-the-art machine learning algorithms and IoT-based systems [104, 105, 106].
In the realm of surgical training and medical procedures, IoT has been harnessed to create next-generation solutions. One such development involves a surgical training framework that employs virtual reality to simulate realistic training environments. This framework also enables interaction with surgeons from around the world, fostering collaborative learning and expertise sharing [107]. Additionally, a human-robot collaborative system has been devised for performing minimally invasive surgeries [108]. Monitoring hemoglobin levels in the blood has become more accessible through portable devices equipped with photoplethysmography (PPG) sensors, light-emitting diodes (LEDs), and photodiodes. These devices enable the non-invasive measurement of hemoglobin levels, enhancing healthcare monitoring and diagnosis [109]. Asthma Monitoring is another notable sector where sensor-based asthma monitoring system is used to track the user’s respiratory rate [110].
Numerous other HIoT applications are currently in use or under research, underscoring the ongoing revolutionary impact of IoT in the field of healthcare. It is expected that IoT will continue to drive advancements and improvements in healthcare delivery and patient outcomes.
1.6.3 Smart Agriculture and Environment
Agriculture holds a crucial position in a country’s economic progress. However, crop yields can be significantly impacted by various factors, including soil moisture and environmental variables like carbon dioxide levels, temperature, and humidity. To enhance agricultural outcomes, it becomes essential to implement robust surveillance systems in the fields. Achieving this efficiently is made possible through the integration of the IoT. The following section has explored several applications in this agricultural domain as well as a summary of the related research work in Table 1.6.
Table 1.6: Relevant studies in smart agriculture and environment domain
Applications Related Works
Water-saving irrigation [111, 112]
Crop growth environment monitoring [113, 114]
Animal and plant life information monitoring [115, 116]
Intelligent agricultural machinery [117, 118]
Agricultural product quality safety and traceability [119, 120]
Diseases Monitoring [121, 122]
Water Saving Irrigation
Water is a crucial factor in crop growth, yet its scarcity is an escalating concern. Different crops have varying water requirements, necessitating a dynamic irrigation approach. The integration of IoT technology in agriculture addresses this challenge by revolutionizing traditional flood irrigation methods, ultimately resolving water shortage problems.
To address these issues, Yang et al. [111] have proposed an intelligent system leveraging neural networks for water-efficient irrigation. This method enhances irrigation efficiency by minimizing human intervention and reducing wastage due to excessive drainage. Li et al. [112] have developed a series of GPRS DTU-based intelligent agricultural irrigation control systems capable of tailoring irrigation to crop growth needs.
Crop Growth Environment Monitoring
Various environmental factors, including temperature, humidity, air pressure, carbon dioxide levels, soil temperature, and soil pH, play a crucial role in crop growth. IoT devices integrated into agricultural systems can sense and analyze these environmental factors, enabling remote field monitoring and the creation of an optimal farming environment tailored to these variables.
Chen et al. [113] have proposed an environmental monitoring solution that utilizes WSNs for millet farmland. In contrast, Lin et al. [114] have designed a wireless environmental monitoring system that harnesses soil energy to enable cost-effective remote monitoring of farmland environments.
Animal and Plant Life Information Monitoring
Efficient agricultural production necessitates the monitoring of both animal and plant information, a pivotal phase in the process. This monitoring plays a substantial role in advancing improved production, increasing profitability, and ensuring high-quality product development.
  1. 1.
    Animal Life Information Monitoring
    Monitoring diverse aspects of animal behavior, including their food consumption, body temperature, activity levels, and health status, enables the tracking of their physiological and nutritional well-being, ensuring their healthy development. Similar to other monitoring systems, animal information monitoring includes various types of sensors, actuators, wireless connections, and terminal equipment.
    In [115], the authors have proposed an infrared-based body temperature measurement system for pigs, allowing the early detection of diseases .
  2. 2.
    Plant Life Information Monitoring
    Plant life information can be assessed through two distinct methods: external information and internal information. External data involves analyzing visual factors such as diseases, pests, and leaf color, while internal information includes parameters like chlorophyll content and photosynthetic rate. Various wireless sensor devices can be connected to plants to facilitate remote and continuous monitoring of their information, ultimately allowing for the early detection of diseases and promoting healthy plant growth.
    Porto et al. [116] have introduced a system for citrus traceability that examines the environmental conditions suitable for the optimal growth of citrus. This system is capable of identifying and preventing plant diseases, ensuring the robust health of citrus crops.
Intelligent Agricultural Machinery
Intelligent machinery demonstrates the capacity to autonomously manage a wide range of agricultural operations, encompassing tasks such as cultivation, sowing, transplanting, fertilization, pesticide application, feeding, irrigation, picking, and harvesting, all executed with precision and efficiency. Moreover, it has the capability to gather a variety of data about the farm, including soil moisture and water quality, as well as ambient information like temperature and humidity. These data resources can be effectively harnessed to implement precision agriculture and enhance breeding practices [123]. IoT technology plays a pivotal role in reducing manual labor in agricultural machinery operations. It enables remote monitoring and standardizes machinery functions through various technologies such as sensors and wireless communication methods.
Sowjanya et al. have introduced a versatile autonomous robot vehicle in [117] where the vehicle is equipped with Bluetooth technology for remote control and is capable of independently executing a range of tasks including farming, seeding, and irrigation. In a similar vein, researchers in [118] have presented a robot system designed for fruit detection. This system employs a single-lens multi-box detection method and a stereo camera to identify the position of fruits in three dimensions, enabling efficient fruit picking.
Agricultural Product Quality Safety and Traceability
IoT plays a pivotal role in enhancing agricultural product quality, safety, and traceability, particularly in the domains of product warehousing, logistics, and distribution. It enables automatic product identification, tracking, and the accurate tallying of incoming and outgoing items within warehouses. By leveraging WSNs, real-time remote monitoring of vehicles or containers used for product delivery can be achieved, ensuring the traceability of agricultural product origins and destinations. Numerous countries have recognized the imperative for real-time traceability and have developed intelligent systems to address this need. These systems include the American agricultural product traceability system, the European beef traceability system, the Swedish agricultural traceability system, the Japanese food traceability system, and the Australian livestock traceability system.
In the pursuit of agricultural product safety and traceability, Jiang et al. [119] have established a comprehensive agricultural product safety traceability platform. This platform enables the real-time automatic collection, processing, and display of agricultural product data, thereby enhancing product traceability and reducing the costs associated with tracking and monitoring agricultural items. Gu et al. have introduced a system in [120] designed to trace the complete lifecycle of pigs, covering stages such as slaughter, processing, and sales, thereby ensuring transparency and traceability throughout the pork production process.
Diseases Monitoring
The integration of IoT for real-time and continuous monitoring offers farmers and relevant authorities the capability to identify diseases at an early stage and implement preventive measures before they escalate. IoT, when combined with machine learning, artificial intelligence, or image processing, has yielded various systems for early disease detection and prevention in agriculture.
The farm’s environment plays a pivotal role in disease occurrence, and in [121], authors have presented a framework that combines various sensor devices through wireless sensor networks for monitoring various environmental factors. This system collects real-time data and employs multiple machine learning models like KNN, Random Forest, Logistic Regression, and Linear Regression to analyze insights from this data. Additionally, a framework for environmental observation-based crop disease monitoring, which continuously monitors plant growth and diseases and notifies farmers when necessary, has been proposed in [122].
1.6.4 Smart Home (SH)
The adoption of smart home technology has seen a significant and continuous increase in recent years. As technologies such as sensors, actuators, and wireless networks have matured, people’s trust in these advancements has grown, leading to a greater reliance on technology to enhance their quality of life.
In a smart home, various types of sensors are strategically deployed, each with its specific function. Smart homes simplify daily tasks for users, proving especially beneficial for those prone to forgetting routine actions like locking doors or turning off appliances. From smart door locks to the maintenance of household items like coffee machines, heaters, and smart bulbs, and even the use of surveillance cameras for enhanced security, smart homes offer a wide range of possibilities. Furthermore, users can control these devices through voice commands and remotely monitor their home equipment. Smart homes contribute to improved energy efficiency by automatically turning off devices not in use and notifying users of any unusual incidents.
MavHome, short for “Managing An Intelligent Versatile Home” [5], represents a smart home system that functions as an intelligent agent. It employs prediction algorithms to perform various tasks in response to user-initiated events. As for energy conservation, an intelligent home achieves energy conservation through the utilization of sensors and the context-aware capabilities of IoT. Data gathered by these sensors are transmitted to a context aggregator, which then forwards the data to a context-aware service engine. This engine analyzes the data and determines appropriate actions. For instance, it may decide to turn off the air conditioning if the temperature is too cold, shut off the gas supply in case of a detected leak, or switch off the lights when there are no occupants at home [124].
1.6.5 Smart Manufacturing System (SMS)
With the development and evolution of IoT (Internet of Things), Industrial IoT, Artificial Intelligence, and Cyber-Physical Systems, many countries have opted to transform their manufacturing systems into smart manufacturing systems. Through the integration of smart technologies, these systems facilitate a rapid and extensive flow of data within and among manufacturing processes. Equipped with these data and employing advanced information and communication technology, smart manufacturing systems possess the capability to swiftly respond to global demands, efficiently utilise materials, energy, and labour resources, and deliver customised products on time [125].
What sets the smart manufacturing model apart from other manufacturing paradigms is its vision of the next generation of manufacturing with enhanced capabilities [126]. Smart manufacturing systems adapt to new circumstances by leveraging real-time information for intelligent decision-making and by proactively predicting and preventing potential failures.
1.6.6 Internet of Robotics Things (IoRT)
The Internet of Robotic Things (IoRT) is a concept that combines the principles of IoT (Internet of Things) and Robotics. IoRT represents an emerging technology that incorporates robots within an IoT ecosystem as objects, enabling communication, collaboration, and automation. These robots are seamlessly integrated into smart environments, where they perform a wide range of tasks. These tasks span from personal activities within smart homes to applications in the healthcare industry. Furthermore, they extend to professional activities such as monitoring, delivery, and object control within manufacturing industries or warehouses.
1.6.7 Oil and Gas Industry
The integration of IoT into the oil and gas industry can enable automated and optimized execution of industry operations, leading to more efficient decision-making. IoT-based applications in the oil and gas sector offer numerous advantages, including enhanced operational efficiency through automation and optimization, resulting in overall cost reduction. Environmental monitoring sensors deployed in this context can detect and alert to the presence of poisonous or harmful substances in the environment. Moreover, IoT technology provides real-time data, significantly improving decision-making processes. Importantly, IoT integration enables the early detection of potential hazards, empowering authorities to take prompt and effective preventive measures.
The oil and gas industry entails substantial expenditures and significant health and safety risks, highlighting the importance of continuous and real-time monitoring within this sector. To achieve this objective, the oil and gas industry is increasingly adopting IoT-based remote monitoring solutions. These solutions allow for the remote monitoring of field equipment, the analysis of field data, collaborative data-driven decision-making, and the implementation of control commands to optimize asset performance while mitigating health, safety, and environmental (HSE) hazards [127]. Furthermore, IoT integration in this industry aims to reduce human labor, minimize time wastage, and enhance accuracy through automation. An illustrative example is Equinor’s deployment of a well optimization system in the Bakken oil field. In this case, Equinor have utilized IoT devices and machine learning algorithms in approximately 50 wells. This implementation led to optimized well operation and maintenance, resulting in a remarkable 33% increase in oil production from these.
1.6.8 Smart Retail
The adoption of IoT in the retail industry has created a flexible environment that benefits both customers and sellers. This shift allows the entire retail sector to migrate from offline to online, enabling customers to independently conduct their shopping through self-service while facilitating smooth interactions between retailers and their customers. Furthermore, IoT technologies like RFID can be employed to monitor products, and sensors can be deployed to collect customer data, which can then be utilized for analyzing customer buying behavior and enhancing business profitability. Additionally, customers have the option to make payments through online transactions and monitor their orders using online services [128].
In the study presented in [129], the authors have introduced a smart unstaffed retail store that relies on AI and IoT for commodity detection. Furthermore, in [130], two models, namely the Retail 4.0 IoT Consumer Model and the Retail 4.0 IoT Retailer Model, have been put forward. These models are capable of analyzing consumer behavior and discerning consumer buying patterns with ease.
1.6.9 Industrial Internet of Things (IIoT)
The IIoT, often recognized as The Industrial Internet of Things (IIoT), holds significant potential according to numerous market researchers. It serves as an extension of IoT specifically customized for the industrial sector and its applications. IIoT empowers industries and enterprises to enhance and optimize their operations by leveraging machine-to-machine communication, big data analytics, and machine learning. The scope of IIoT is extensive, encompassing a wide array of connected industrial devices and systems. These include, but are not limited to, connected electric meters, wastewater systems, flow gauges, pipeline monitors, manufacturing robots, and various other types of industrial equipment and devices [131].
One notable application of IIoT is in the mining industry, where companies like CISCO have implemented IIoT solutions to improve safety and efficiency in underground mines. These solutions involve connecting people, tracking the locations of miners and vehicles, monitoring vehicle statuses, and automating building controls. Additionally, the mine’s blasting system is integrated with the personnel location tracking system to enhance miner safety [132]. Moreover, IoT has the potential to revolutionize the manufacturing field by enhancing productivity, maintenance practices, machine monitoring, and ensuring the safety of workers [133].
1.6.10 Social Life and Entertainment
In addition to work or professional activities, social life and entertainment are integral parts of a person’s life, and several applications have been developed to monitor and enhance human social activities. Portable devices such as mobile phones and tablets possess sensing capabilities and communication technologies that facilitate interactions between individuals. Integrating IoT into an individual’s social life can contribute to emotion detection, community building, and emotional support.
For instance, CircleSense [134] is an application designed to identify a person’s social circle by analyzing their social activities using various sensors. It also tracks the person’s location via location sensors and employs Bluetooth technology to identify people in proximity. Affective computing is a technology that analyzes various aspects of a person’s behavior, including facial expressions, speech, body gestures, hand movements, and sleep patterns, to identify and respond to their emotions. An application of affective computing is Camy [135]. Camy is an artificial pet dog designed to interact with human beings, displaying affection and empathy. It assesses the owner’s mood and provides suggestions, emotional support, and encourages playful behavior accordingly.
A Table 1.7 has been provided for presenting information about relevant works in the aforementioned domains.
Table 1.7: Relevant works in other applications
Applications Related Works
Smart home (SH) [5, 136]
Smart manufacturing system (SMS) [125, 126]
Internet of robotics things (IoRT) [137, 138]
Oil and gas [127]
Smart Retail [129, 130]
Industrial IoT [132, 133]
Social Life and Entertainment [134, 135]
1.7 Challenges and Future Directions
The remarkable position that IoT now occupies in today’s world was once only a dream a few years ago. Today, IoT has captivated the entire globe and continues to extend its reach into various domains. Nonetheless, IoT also grapples with numerous issues and challenges that pose hurdles to its seamless implementation and expansion. This section offers an in-depth exploration of the key challenges confronting IoT.
1.7.1 Broad and Open Research Challenges
Broad and open research challenges refers to complex and critical challenges in the IoT system that do no have any straightforward or predefined solutions. These challenges require extensive research, exploration in order to propose efficient solutions. They are “broad” in the sense that they encompass a wide range of related issues and considerations, and they are “open” because they may not have clear-cut or definitive answers, leaving room for ongoing research, experimentation, and discovery. This subsection provides a comprehensive listing of challenges in the IoT domain.
Building Intelligent Environments based on IoT Paradigm
Creating a smart environment requires a vast number of devices, sensors, and complementary technologies to facilitate their inter-connectivity. Managing this large volume of objects constitutes the initial challenge in the realm of IoT and intelligent environments. Moreover, the substantial task of collecting, storing, and conducting efficient analyses on enormous amounts of data remains a significant concern and creates collision issues within the IoT framework [139]. Handling massive amounts of devices and data in the IoT requires. One approach is to employ decentralized systems instead of centralized ones, reducing the volume of data sent to the cloud for processing. Techniques like data filtering, compression, and load balancing can further minimize data size. Utilizing IoT technologies with robust device management and maintenance capabilities is also beneficial. Additionally, leveraging big data technologies like Hadoop and Spark can efficiently handle the substantial IoT data volumes. This holistic approach ensures readiness for the expanding IoT landscape.
Privacy and Security Challenges of IoT Applications
The heterogeneity of IoT devices and the diversity of various IoT applications result in various security and privacy issues. People are primarily concerned about potential privacy invasion and security threats when using these technological devices. Further details are provided below.
  1. 1.
    Security
    Different layers of IoT are vulnerable to various kinds of attacks based on the technologies and protocols used in these respective layers. According to [140], IoT layers including perception layer, network layer, and application layer face various security attacks. As the major role of perception layer is to collect data, so the security challenges in this layer focuses in Falsifying the data and destroying perception devices. Attacks including Node Capture Attacks, Malicious code Injection Attacks, False Data Injection Attacks, Replay Attacks, Cryptanalysis Attacks & Side Channel Attacks, Eavesdropping and Interference and Sleep Deprivation Attacks are faced by this layer [140].
    Since the basic function of network layer is to transmit collected data, particularly using wireless technologies, security challenges in this layer revolve around the availability of network resources and the wireless network. Challenges at the Network Layer encompass Denial-of-Service (DoS) Attacks, Spoofing Attacks, Sinkhole Attacks, Wormhole Attacks, Man-in-the-Middle Attacks, Routing Information Attacks, Sybil Attacks, and Unauthorized Access [140]. Application layer, which focuses on providing user-requested services, challenges primarily revolve around software attacks, including: Phishing Attack, Malicious Virus/Worm, Malicious Scripts [140].
    It is imperative to propose secure, robust, and reliable authentication schemes to detect and defend unauthorized access [141]. To defend against worms, viruses, and malicious scripts bridging firewalls, virus detection techniques and script detection techniques such as honeypot techniques, static code analysis, and dynamic action detection must be implemented. Along with that, secured routing protocols must be introduced to ensure secure routing.
  2. 2.
    Privacy
    As previously mentioned, IoT devices continuously generate vast amounts of real-time data, which undergo three main stages: (1) Data Collection, (2) Data Aggregation , and (3) Data Mining and Analytics. While these processes enhance our lives by providing various services, they also raise concerns about data privacy in IoT. Privacy breaches in IoT can have serious repercussions for both the IoT network and its users, including financial losses, property damage, and even risks to human safety and security [140, 142].
    For instance, consider the smart grid, Adversaries can readily seize control of the smart meters, allowing them to access or manipulate the collected data. This could potentially compromise the confidentiality and privacy of energy consumption data. By using these altered data, utility providers may make inaccurate assessments of energy supply and demand within the grid, resulting in erroneous energy dispatch decisions. This, in turn, could lead to imbalances in energy supply and demand, potentially causing widespread power outages. In the healthcare industry, if an adversary manages to acquire a patient’s health data, they could manipulate medication prescriptions or medical records, leading to significant health risks and potential insurance fraud. Therefore, it is imperative to deploy privacy preservation schemes to prevent data leakage and ensure that private data remains inaccessible to adversaries [140, 143].
    Privacy-preserving mechanisms in the context of IoT data processing can be categorized into three main groups: (i) privacy preservation during data collection, (ii) privacy preservation during data aggregation, and (iii) privacy preservation during data mining and analytics. While various techniques such as encryption and key management can be applied to protect privacy in data collection, mining, and analytics, the majority of efforts in IoT privacy preservation have focused on data aggregation. Data aggregation involves processing relevant data in multiple locations, making it challenging to ensure privacy using traditional encryption methods. Therefore, several privacy-preserving mechanisms have been developed specifically for data aggregation, falling into the following categories:
    1. (a)
      Anonymity-based privacy preservation, which employs techniques like K-anonymity, L-diversity, and T-closeness to protect the privacy of identification information during data aggregation.
    2. (b)
      Encryption-based privacy preservation, which utilizes encryption techniques like homomorphic encryption, commitment mechanisms, secret sharing, and zero-knowledge proofs to prevent data from being eavesdropped by adversaries during aggregation [144, 145].
    3. (c)
      Perturbation-based privacy preservation, where techniques such as data customization, data sharing, and random noise injection are used to perturb raw data, ensuring privacy during aggregation.
    Among these, perturbation-based privacy-preserving schemes are popular in IoT due to their direct operation on raw data. However, many of these perturbation-based privacy preserving schemes sacrifice data utility to achieve privacy. This reduction in data utility can hinder the support of services requested by IoT applications. Therefore, a significant challenge, in the field of data privacy preservation in IoT, is designing schemes that strike a balance between privacy and data utility, making it a crucial area for future research [21, 140]. In summary, safeguarding data privacy in IoT is crucial to prevent these adverse outcomes and maintain the security and integrity of both individuals and the IoT ecosystem.
Compatibility
In an IoT network, devices from various vendors are interconnected, which can lead to monitoring and management challenges. Different industries currently rely on a multitude of standards to support their applications. Given the vast amounts of data, diverse device types, and the presence of various entities, utilising standard interfaces becomes crucial. This importance is amplified, especially for applications that need to accommodate both cross-organisational collaborations and a wide array of system limitations. Addressing these issues requires all industries to adhere to specific standards, but achieving such universal compliance can be a daunting and impractical task.
Scalability
IoT is an ever-expanding network, and heterogeneous devices are expected to continuously join the IoT network in the future. As a result, as the number of devices increases, ensuring smooth connectivity, effective data management, and overall system performance at a small scale becomes increasingly challenging. Therefore, the scalability of IoT poses an ongoing challenge for the future of this technology. To effectively address scalability challenges, it is essential to construct a scalable architecture, utilizing technologies like modular components, load balancers, and distributed systems.
Energy Efficiency
Small smart devices that make up IoT systems often have limited battery power, which cannot be easily replaced. This limitation can lead to a global energy crisis and high power consumption, as well as constraints on memory and processing capabilities. Consequently, routing processes and compute-intensive applications may not run efficiently on these devices. While some routing protocols do support low-power communication, they are still in the early stages of development, and the constrained energy of smart devices may not be sufficient to fully utilize these WSN routing protocols. To tackle these challenges effectively, it’s essential to emphasize the creation of low-power hardware and the adoption of energy-efficient protocols like MQTT-SN or CoAP, rather than relying on more power-intensive alternatives such as HTTP. Additionally, harnessing over-the-air (OTA) firmware updates can ensure devices remain optimized and bug-free, thus diminishing the necessity for physical maintenance visits. Duty cycling, as another viable approach, aids in curbing power consumption significantly.
Mobility Management
Mobility management in IoT refers to the ability to handle devices that move within the network seamlessly. It is a crucial aspect because many IoT devices are not stationary and need to communicate as they change locations. The presence of mobile devices in IoT setups can lead to challenges in how routing protocols and IoT networks work efficiently. The current methods used for devices that move, like in sensor networks, mobile adhoc networks, and vehicular networks, can’t effectively handle the various problems related to routing because of these sensors having limited processing power and energy resources. To address these challenges, IoT systems employ various mobility management techniques and protocols, aiming to provide reliable and seamless communication for mobile devices in the IoT ecosystem.
Cost of Maintenance and Services
The IoT network consists of a vast number of devices, utilizing various costly communication technologies. This inevitably leads to increased maintenance and service costs for these numerous devices and connections. Hence, a significant challenge lies in addressing this issue by designing devices and sensors that demand minimal maintenance.
Internet Disconnection Problem
The disruption of internet connectivity, which is central to IoT operations, results in inferior performance from IoT applications and a decline in service quality. Furthermore, restrictions on the number of devices that can concurrently interact with the base station limit user access to these services. This issue is especially problematic in remote or unreliable network settings, where sustaining a consistent internet connection proves challenging. Consequently, addressing the problem of internet disconnections in IoT is imperative to maintain the reliability and efficiency of IoT systems.
Processing, Analysis and Management of Data
The procedure for Processing, Analysis and Management of Data is tremendously challenging because of the heterogeneous nature of IoT devices and large sale of data generation. Currently, most systems utilize the centralized cloud based system for performing computationally intensive tasks and delivery of data. However, an ongoing concern revolves around the limitations of traditional cloud architectures when it comes to efficiently handling the vast amounts of data generated and utilized by IoT-enabled devices. Additionally, these architectures struggle to support the associated computational demands while also meeting precise timing constraints. To address this challenge, most systems are currently relying on existing solutions like mobile cloud computing and fog computing, both of which utilize edge processing [146].
Figure 1.6: Ethical Issues in IoT
Other Challenges
In addition to the previously mentioned challenges, IoT technology faces several other issues. The widespread adoption of IoT devices and technology, coupled with our increasingly reliant lifestyles, has led to users becoming highly dependent on IoT applications. This reliance is particularly critical in healthcare, where patients heavily depend on healthcare applications. Moreover, IoT devices can sometimes unexpectedly interfere with human activities, resulting in unanticipated and autonomous behaviors. The IoT network introduces ambiguity, making it challenging to distinguish between physical and virtual devices and even humans due to the ease of transformation between these categories. Quality and traffic control have become more complex due to the miniaturization and the huge number of IoT devices. Managing unique identifications for each IoT device is also a growing concern. Furthermore, IoT goes beyond geographical boundaries, with applications like healthcare offering services internationally. This global reach poses challenges for nations, as data generated within their borders can be collected and transmitted to service providers located anywhere in the world, raising issues related to data privacy and jurisdiction. Addressing these multifaceted challenges will require careful consideration and international collaboration to ensure the effective and secure implementation of IoT technology.
1.7.2 Ethical Considerations
The term “ethical issue” pertains to a situation or quandary characterized by a clash of moral principles, values, or ethical norms [147]. The realm of IoT confronts these ethical quandaries, necessitating individuals or organizations to make challenging choices amidst conflicting interests. These decisions frequently revolve around determining what is ethically correct or incorrect. Figure 1.6 illustrates prevalent such ethical dilemmas within the IoT domain.
The ethical concerns related to IoT can be divided into five main categories, as adapted from [148], and are outlined below. These issues are put forth with the aim of safeguarding privacy rights through the regulation of how organizations manage information generated by IoT devices. While these ethical standards are designed to establish guidelines for organizations, they should also prompt individual concerns regarding their privacy, as these principles serve as legal safeguards to protect individuals.
  • Information privacy considerations
    Organizations must handle the produced data with both openness and transparency. Except in specific cases, they should offer choices to individuals allowing them not to disclose their identity or to use a pseudonym.
  • Information collection
    Organizations can gather requested data by implementing more stringent criteria for acquiring ‘sensitive’ information. Conversely, they should specify their approach to handling unsolicited information. In both situations, organizations are required to delineate the circumstances surrounding the collection of this information and provide prior notification to relevant parties.
  • Managing data
    Organizations should specify the situations in which they might utilize or share the information they have gathered. Under specific conditions, an organization can employ personal data for direct marketing purposes. Nonetheless, they have the option of disclosing it internationally; but, before doing so, they must set the safeguards that will be used to protect this information.
  • Integrity of information
    The information gathered and shared by entities should be precise, current, and comprehensive. Reasonable precautions must be in place to prevent misuse, interference, loss, and unauthorized access, changes, or disclosure.
  • Information rectification and availability
    When requesting access to their information, entities should clearly state their responsibilities for allowing access and making corrections to the information they possess. This involves the obligation to grant access and make necessary modifications, except in cases where a particular exception is applicable.
1.7.3 Legal and Regulatory Issues
After identifying the challenges and ethical concerns, there is a need to consider the legal aspects concerning the effectiveness of current laws in safeguarding users within this context. This concern is significant because the boundary between the physical and virtual realms in IoT is increasingly blurred. The ensuing questions serve as examples of the issues that require discussion, as referenced in [147].
  1. 1.
    What are the consequences in the event of an Internet connection failure?
  2. 2.
    Who bears the responsibility or liability for maintaining the security of IoT devices, routers, and connections to the cloud?
  3. 3.
    Can one be certain that unauthorized access to a home’s internal network will not occur due to hacking within IoT services on the cloud side?
  4. 4.
    What are the consequences when an IoT service provider encounters a period of unavailability for essential life-support devices?
  5. 5.
    What are the consequences of an IoT device operating without the explicit consent of its owner or functioning in unintended manners, such as erroneously placing orders or engaging in excessive vacuuming during unconventional hours?
  6. 6.
    What occurs if the vendor of an IoT product ceases operations and discontinues product support?
  7. 7.
    To whom does the data gathered by IoT devices belong?
  8. 8.
    Are there situations in which it is inappropriate for IoT devices to gather data?
  9. 9.
    What occurs when the user decides to withdraw their participation?
  10. 10.
    How can one address the issue of individuals who lack access to smart devices or the skills to operate them? (Digital divide)
Considering the challenges mentioned earlier, it is evident that IoT faces several obstacles, including financial constraints, security vulnerabilities, and data privacy concerns, which can have life-threatening implications, especially in healthcare data breaches. To address these challenges effectively, a proactive approach involving extensive research is crucial. This research should focus on identifying IoT’s specific issues, followed by the implementation of robust technical solutions. Effective execution of this process will not only ensure a secure IoT system but also encourage user trust in enrolling themselves within the IoT network.
To mitigate ethical concerns, raising user awareness is essential, along with the integration of self-adaptive security policies and dynamically modifiable policies during IoT application development. The introduction of new laws and standards is also necessary, integrating existing regulations like HIPPA, FIPPS, the Electronic Communication Privacy Act, and others to comprehensively address security, privacy, and legal issues. Additionally, addressing technical challenges involves the introduction of adaptable and new standards, as well as the implementation of standard address identification. Some examples of these technical solutions include advanced encryption techniques, electronic signatures, the integration of standard protocols, and regulations to limit third-party data usage. This holistic approach aims to overcome the multifaceted challenges that IoT faces.
1.8 Conclusions
The IoT has rapidly become an integral part of the 21st century, enhancing daily decision-making and ushering in innovative consumer services like pay-as-you-use. The seamless integration of smart devices and automation technologies has revolutionized our lives. However, amidst this technological marvel, we must acknowledge significant concerns related to security, privacy, intellectual property rights, safety, and trust. These concerns continue to demand further investigation. This chapter has provided a comprehensive overview of IoT for newcomers seeking to explore this domain and gain a thorough understanding to make future contributions. Specifically, the chapter begins by discussing the fundamental concept of IoT, its historical development, network components, and key characteristics. Subsequently, it explores the advantages of employing IoT in various domains and introduces readers to the fundamental layered architectures that underpin IoT. Furthermore, the chapter delves into the taxonomy of IoT technologies, which not only encompasses hardware, software, and wireless communication technologies but also various hardware platforms and cloud solutions. It also reviews existing applications across major domains such as smart cities, healthcare, and agriculture, offering readers insights into the breadth of IoT applications. Finally, the chapter addresses the prevalent issues and challenges associated with the deployment and design of IoT applications. It explores security threats across architectural layers, ethical dilemmas, privacy concerns faced by users, and issues related to trust. This discussion serves as a solid foundation for researchers interested in developing practical IoT projects or pioneering new theoretical approaches within the IoT field, equipping them with a deep understanding of various aspects of IoT. This, in turn, provides a good ground for the researchers who are whether interested in designing realistic IoT projects or developing novel theoretical approaches in the IoT field by acquiring deep knowledge in different IoT aspects.

### 12 Jan 2024 v2

@@
 Abstract
-The Internet of Things (IoT) has rapidly transformed the 21st century, enhancing decision-making processes and introducing innovative consumer services such as pay-as-you-use models.
-This integration of smart devices and automation technologies has revolutionized our lives.
-However, it is essential to recognize the significant concerns surrounding security, privacy, intellectual property rights, safety, and trust in this technological landscape, which require further exploration.
-This chapter serves as a comprehensive guide to newcomers interested in the IoT domain, providing a foundation for making future contributions.
-It begins by explaining the core concept of IoT, its historical evolution, network components, and key characteristics.
-The advantages of IoT deployment across various domains are discussed, along with an introduction to the foundational layered architectures supporting IoT.
-The chapter also delves into the taxonomy of IoT technologies, encompassing hardware, software, wireless communication technologies, hardware platforms, and cloud solutions.
-Existing applications in major domains like smart cities, healthcare, and agriculture are explored, offering a broad perspective on the IoT’s application potential.
+The Internet of Things has rapidly transformed the 21st century, enhancing decision-making processes and introducing innovative consumer services such as pay-as-you-use models.
+The integration of smart devices and automation technologies has revolutionized every aspect of our lives, from health services to the manufacturing industry, and from the agriculture sector to mining.
+Alongside the positive aspects, it is also essential to recognize the significant safety, security, and trust concerns in this technological landscape.
+This chapter serves as a comprehensive guide for newcomers interested in the IoT domain, providing a foundation for making future contributions.
+Specifically, it discusses the overview, historical evolution, key characteristics, advantages, architectures, taxonomy of technologies, and existing applications in major IoT domains.
 In addressing prevalent issues and challenges in designing and deploying IoT applications, the chapter examines security threats across architectural layers, ethical considerations, user privacy concerns, and trust-related issues.
-This discussion equips researchers with a solid understanding of diverse IoT aspects, providing a strong foundation for both practical IoT projects and the development of novel theoretical approaches within the IoT field.
-Keywords: Taxonomy, Architectural Layers, Ethical Considerations, Security Threats, Hardware Platforms, Trust-related Issues.
+This discussion equips researchers with a solid understanding of diverse IoT aspects, providing a comprehensive understanding of IoT technology along with insights into the extensive potential and impact of this transformative field.
+Keywords: Architectural Layers, Ethical Considerations, Security Threats, Hardware Platforms, Trust-related Issues, Taxonomy of technologies.
 1.1 Introduction
@@
 The IoT journey, far from its conclusion, promises a future where diverse devices seamlessly connect to the web, reshaping human existence in unprecedented ways.
-The Internet of Things (IoT) is a network of interconnected devices with sensors, actuators, processors, and various communication technologies.
+The IoT is a network of interconnected devices with sensors, actuators, processors, and various communication technologies.
 Sensors collect real-time data from both internal states and external surroundings, ranging from mobile phones to microwave ovens.
@@
 These examples only scratch the surface of IoT’s broad application landscape, indicating the exciting possibilities and challenges that researchers are exploring for the future.
-The Internet of Things (IoT) has the potential to transform how people interact with technology, providing greater convenience, efficiency, and personalization in daily life.
+The IoT has the potential to transform how people interact with technology, providing greater convenience, efficiency, and personalization in daily life.
 Despite its transformative impact, IoT faces challenges.
@@
 Beyond technical hurdles, various general and domain-specific challenges are crucial for the success of IoT.
-Identifying and addressing these multifaceted challenges collectively is essential to unlock the full potential of IoT and overcome obstacles for broader adoption and integration into our lives.
+Identifying and addressing these multifaceted challenges collectively is essential to unlocking the full potential of IoT and overcoming obstacles for broader adoption and integration into our lives.
 Motivation of this chapter: IoT is not a novel concept; researchers have been exploring this field for decades.
 Consequently, the question arises: what sets this chapter apart from others, and why should readers invest their time in exploring the basics presented here?
-While there are several book chapters introducing IoT, many share outdated concepts, lacking the latest insights from the ongoing exploration in this dynamic field.
+While there are several book chapters introducing IoT, many share outdated concepts, lacking the latest insights from the ongoing exploration of this dynamic field.
 For instance, in [8], only IoT applications are discussed, and in [9], authors overlook major domains, including IoT applications, advantages, challenges, and technologies.
@@
 Readers are encouraged to explore this chapter for a thorough and up-to-date understanding of the evolving landscape of IoT.
-Authors’ Contributions: The authors’ contributions in this chapter can be summarized as follows:
+Contributions: The contributions of this chapter can be summarized as follows:
   1. 1.
@@
     In conclusion, this chapter offers a comprehensive discussion of the open research challenges, ethical considerations, as well as the legal and regulatory aspects of IoT.
+Table 1.1: List of abbreviations
+Name Abbreviation Name Abbreviation Name Abbreviation
+FDMA Frequency Division Multiple Access LBT Listen Before Talk CDMA Code Division Multiple Access
+FHSS Frequency Hopping Spread Spectrum CSS Chirp Spread Spectrum OOK On-Off Keying
+TDMA Time Division Multiple Access GSM
+Global System for Mobile
+Communications
+PPM Pulse Position Modulation
+GFSK Gaussian Frequency Shift Keying UMTS
+Universal Mobile Telecommunications
+System
+OFDM
+Orthogonal Frequency Division
+Multiplexing
+DWPSK
+Differential Quadrature Phase Shift
+Keying
+GMSK Gaussian Minimum Shift Keying TDD Time Division Duplex
+CSMA/CA
+Carrier Sense Multiple Access with
+Collision Avoidance
+LTE-A Long-Term Evolution Advanced FDD Frequency Division Duplex
+BPSK Binary Phase Shift Keying 3GPP 3rd Generation Partnership Project ALHOA Adaptive Link Hopping On Air
+CSMA/CD
+Carrier Sense Multiple Access with
+Collision Detection
+WCDMA
+Wideband Code Division Multiple
+Access
+DBPSK
+Differential Binary Phase Shift
+Keying
+O-QPSK Offset Quadrature Phase Shift Keying OFDMA
+Orthogonal Frequency Division
+Multiple Access
+QAM Quadrature Amplitude Modulation
+QPSK Quadrature Phase Shift Keying SC-FDMA
+Single Carrier Frequency Division
+Multiple Access
+CP-OFDM
+Cyclic Prefix Orthogonal Frequency
+Division Multiplexing
 Chapter organization: It begins with an exploration of the fundamentals in Section 1.2, covering the history 1.2.1, components 1.2.2, and characteristics 1.2.3.
-In Section 1.3, the advantages of IoT are examined, with a focus on both the general perspective and specific domains such as healthcare and agriculture.
+In Section 1.3, the advantages of IoT are examined.
 By introducing a generic architecture in Section 1.4, the taxonomy of technologies that underpin IoT operations are discussed in detail in Section 1.5.
@@
 It concludes in Section 1.7 by addressing the key challenges that IoT faces today, considering insights from the reviewed literature, and sheds light on potential future scopes and developments in the ever-evolving realm of IoT technology.
+Table 1.1 represents some abbreviations mentioned in this chapter.
 1.2 Fundamentals of IoT
@@
 The advantages brought by this technology are unquantifiable, as it touches nearly every aspect of people’s lives, simplifying and improving them.
-This section comprehensively discusses these fundamental IoT concepts, covering the following topics: The History of IoT, Components of IoT, Characteristics of IoT.
-1.2.1 Historical Development of IoT
+This section comprehensively discusses these fundamental IoT concepts, covering the following topics: history, components, characteristics.
+1.2.1 Historical Development
 Since the invention of the first landline, the telegraph, in the 1830s and 1840s, machines have been instrumental in facilitating direct communication.
@@
 This led to the practical use of systems for remotely monitoring electrical grid meters via telephone lines during that era [12].
-The 1980s saw a growing interest in enhancing physical objects with sensors and intelligence.
+In the 1980s, there was a growing interest in enhancing physical objects with sensors and intelligence.
 Commercial service providers began supporting public access to ARPANET, an early precursor to the modern Internet, during this time.
@@
 In the 1990s, advancements in wireless technology set the stage for the widespread adoption of ”machine-to-machine” (M2M) solutions in enterprise and industrial contexts, particularly for equipment monitoring and operation.
-However, many of these early M2M solutions relied on closed, purpose-built networks and proprietary or industry-specific standards, rather than utilizing Internet Protocol (IP)-based networks and Internet standards [14].
+However, many of these early M2M solutions relied on closed, purpose-built networks and proprietary or industry-specific standards rather than utilizing Internet Protocol (IP)-based networks and Internet standards [14].
 By the mid-1990s, the Internet had expanded its global reach, offering new possibilities for researchers and technologists to explore ways to enhance connections between humans and machines.
@@
 In 1991, Mark Weiser’s paper “The Computer of the 21st Century” [16] and academic events like UbiComp and PerCom shaped the contemporary IoT vision [14].
-Global Positioning Satellites (GPS) became a reality in early 1993 with the Department of Defense establishing a stable system of 24 satellites.
+Global Positioning System (GPS) became a reality in early 1993 with the Department of Defense establishing a stable system of 24 satellites.
 Privately owned commercial satellites soon followed, enhancing IIoT functionality [3].
 In early 1994, Reza Raji introduced the IoT concept in IEEE Spectrum, describing it as “moving small data packets to integrate and automate from home appliances to entire factories” [17].
-Steve Mann’s near-real-time WearCam, powered by a 64-processor setup, was invented later that year.
+Later that year, Steve Mann invented the near-real-time WearCam, powered by a 64-processor setup.
 Between 1993 and 1997, companies proposed IoT solutions, including Microsoft’s “at Work” and Novell’s NEST.
@@
 Lewis defined IoT as the integration of people, processes, and technology with connectable devices and sensors for remote monitoring, status assessment, manipulation, and trend evaluation related to these devices [3].
-In 1997, Paul Saffo described sensors and their future role.
-British technologist Kevin Ashton, while serving as the executive director of the Auto-ID Centre at MIT, independently coined the term “Internet of Things”.
-During his time at Procter & Gamble, he explored radio-frequency identification (RFID), a technology framework enabling physical devices to connect via microchips and wireless signals.
+In 1997, Paul Saffo described sensors and their future roles.
+British technologist Kevin Ashton, while serving as the executive director of the Auto-ID Center at MIT, independently coined the term “Internet of Things”.
+During his time at Procter and Gamble, he explored radio-frequency identification (RFID), a technology framework enabling physical devices to connect via microchips and wireless signals.
 In the same year, they developed a global RFID-based item identification system [14].
-In 1999, Kevin Ashton was the first to describe the IoT and proposed the name “Internet of Things” during a presentation for Procter & Gamble.
-He believed RFID technology, primarily designed for inventory tracking, was a significant prerequisite for the Internet of Things, allowing computers to efficiently manage and monitor individual objects.
+In 1999, Kevin Ashton was the first to describe the IoT and proposed the name “Internet of Things” during a presentation for Procter and Gamble.
+He believed RFID technology, primarily designed for inventory tracking, was a significant prerequisite for the IoT, allowing computers to efficiently manage and monitor individual objects.
 The concept of tagging objects has been realized through technologies like digital watermarking, barcodes, and QR codes, used for identification and tracking purposes [18].
 Subsequent technological advancements, including the proliferation of smartphones, cloud computing, improved processing power, and enhanced software algorithms, along with the availability of sophisticated sensors capable of measuring various parameters, laid the foundation for robust data collection, storage, and processing for the IoT’s growth.
-As a significant step forward in commercialising IoT, LG announced plans to launch a smart refrigerator capable of autonomously managing its contents in 2000.
+As a significant step forward in commercializing IoT, LG announced plans to launch a smart refrigerator capable of autonomously managing its contents in 2000.
 Walmart and the US Department of Defense pioneered inventory tracking using RFID and the IoT in $2002-2003$.
@@
 Cisco Systems estimated that the IoT was ”born” between 2008 and 2009, with the things/people ratio growing from 0.08 in 2003 to 1.84 in 2010 [18].
-1.2.2 Components of IoT
+1.2.2 Components
 The IoT consists of several key components that serve as the essential building blocks for constructing an IoT system.
-This section provides an in-depth exploration of the principal components within IoT.
-IoT comprises three main components: (1) Sensors/Devices and Actuators, (2) Storage and Data Analytics, and (3) Interpretation and Visualization Tools.
+This section provides an in-depth exploration of the principal components of the IoT.
+IoT comprises three main components: (1) sensors/devices, and actuators; (2) storage and data analytics; and (3) interpretation and visualization tools.
 Each of these is further categorized into various subcomponents.
@@
   1. 1.
-    Sensors/Devices
-    Sensors play a crucial and essential role within an IoT system.
+    Sensors/Devices: Sensors play a crucial and essential role within an IoT system.
 Given that IoT operates by gathering data from the surrounding environment, it is necessary for all IoT applications to incorporate one or more sensors to meet this need.
@@
     1. (a)
-      Mobile based Sensors
-      Smartphones, which are widespread and commonly used, are equipped with various sensors, including accelerometers and environmental monitoring sensors.
+      Mobile-based Sensors
+      Smartphones, which are widespread and commonly used, are equipped with various sensors.
 Given their extensive usage, researchers are exploring the potential of using smartphones as integral components in building smart IoT solutions.
 These applications can harness sensor data from smartphones to generate valuable insights and outcomes.
-Some of the general sensors found in smartphones include [19]:
-      1. i.
-        Accelerometer: This sensor monitors and detects the motion or acceleration of the device, typically measuring changes in the smartphone’s velocity in three dimensions [20].
-The data patterns recorded by the accelerometer can be employed to identify the user’s physical activities, including running, walking, and cycling.
-      2. ii.
-        Gyroscope: The gyroscope introduces an extra dimension compared to the accelerometer by being able to detect rotation or tilting of the phone.
-      3. iii.
-        GPS: The Global Positioning System (GPS) employs satellites to determine the location of a mobile phone.
-      4. iv.
-        Magnetometer: The magnetometer detects the magnetic field and can function as a compass.
-      5. v.
-        Light Sensor: The light sensor measures the intensity of light in the environment around it.
-It can be applied in different ways, such as managing the room’s lighting or adjusting the screen’s brightness.
-      6. vi.
-        Proximity Sensor: The proximity sensor employs an infrared (IR) LED that emits infrared rays.
-These rays bounce off objects they encounter, and by measuring the time it takes for them to return, one can calculate the distance to those objects.
-This technology enables us to measure how far things are from the phone.
-      Certain smartphones, like the Samsung Galaxy S4, come equipped with extra sensors, including a thermometer, barometer, and humidity sensor.
-These sensors are used to measure temperature, atmospheric pressure, and humidity levels, respectively [4].
+Some of the general sensors found in smartphones include an accelerometer, gyroscope, GPS, magnetometer, light sensor, and proximity sensor [19].
+Certain smartphones, like the Samsung Galaxy S4, come equipped with extra sensors, including a thermometer, barometer, and humidity sensor [4].
     2. (b)
       Medical Sensors
-      The healthcare industry is one of the most influenced fields where innovation and IoT have paved the way.
+      The healthcare industry is one of the most influential fields where innovation and IoT have paved the way.
 Wearable devices and sensors have facilitated remote monitoring for physicians and enabled researchers to collect data continuously and in real-time.
-A variety of wearable devices are now on the market for measuring physiological data, including blood pressure, heart rate, respiration, and glucose levels, [21].
 These devices come in different forms, such as wristbands, smartwatches, and monitoring patches.
@@
       Neural Sensors
-      Neural sensors play a crucial role in comprehending the workings of human neurons.
-They enable us to decode brain signals, assess the brain’s current state, and, when necessary, optimize it for improved focus and attention.
+      Neural sensors play a crucial role in comprehending the workings of human neurons by enabling us to decode brain signals, assess the brain’s current state, and, when necessary, optimize it for improved focus and attention.
 This practice is commonly referred to as Neurofeedback [22].
@@
       Environmental and Chemical Sensors
-      Environmental sensors serve the purpose of measuring various environmental factors, including temperature, humidity, pressure, and pollution, among others.
-Traditional tools like thermometers and barometers can effectively measure parameters such as temperature and pressure.
-When it comes to assessing air quality, specialized sensors come into play.
-These sensors detect the presence of gases and particulate matter in the air [23].
-      Additionally, chemical sensors play a crucial role in detecting both chemical and biochemical substances.
+      While conventional tools manage parameters such as temperature and pressure, specialized environmental sensors play a crucial role in evaluating air quality.
+These sensors detect gases and particulate matter [23], while also measuring factors like temperature, humidity, pressure, and pollution.
+Besides, chemical sensors play a crucial role in detecting both chemical and biochemical substances.
 Among the innovative technologies available are the electronic nose (e-nose) and electronic tongue (e-tongue), which rely on pattern recognition to sense chemicals based on odor and taste.
@@
     5. (e)
-      RFID
-      RFID technology finds widespread use in various IoT applications, serving as sensors.
+      Radio-Frequency Identification (RFID)
+      RFID technology, which serves as sensors, finds widespread use in various IoT applications.
 For instance, it is employed for tracking products within extensive inventories or monitoring items within large retail stores.
   2. 2.
-    Actuators
-    Actuators hold a crucial role and operate in direct contrast to sensors.
+    Actuators: Actuators hold a crucial role and operate in direct contrast to sensors.
 They transform energy into physical motion and are typically positioned on the outer periphery of a system.
@@
   1. 1.
-    Gateway
-    The IoT gateway are designed to simplify the IoT system.
-The gateway functions as the intermediate medium of communication between the devices and the central cloud system.
-The major functionality of IoT gateways are listed below:
+    Gateway: Gateways, designed to simplify the IoT system, function as the intermediate medium of communication between the devices and the central cloud system.
+The major functionalities of IoT gateways are listed below:
     1. (a)
       Data Preprocessing
-      The IoT gateway acts as an intermediary between sensor devices and the central cloud.
-Prior to forwarding data directly to the cloud, basic data analytics are performed in this stage.
+      The IoT gateway serves as an intermediary between sensor devices and the central cloud, conducting basic data analytics before forwarding information directly to the cloud.
 This layer performs the tasks of local data filtering, cleaning, preprocessing, and protocol translation.
@@
       Data Acquisition
-      In this layer, data is collected from multiple sources and then transferred to the processing layers.
-Data is collected and converted into desired format in this layer by utilizing Data acquisition systems.
-The role of gateway in this stage is to provide secure connectivity between IoT devices and processing structure [26].
+      In this layer, data is collected from multiple sources, converted into the desired format, and then transferred to the processing layers.
+The role of the gateway in this stage is to provide secure connectivity between IoT devices and processing structures [26].
     3. (c)
@@
       Device Management
-      This layer cacilitates real-time device configuration, allowing adjustments to device statuses, operational modes, error acknowledgments, and more [27].
+      This layer facilitates real-time device configuration, allowing adjustments to device statuses, operational modes, error acknowledgments, and more [27].
     5. (e)
@@
   2. 2.
-    Cloud
-    The cloud serves as the central hub of an IoT network, taking on pivotal roles in data processing, storage, and management.
-Among its various responsibilities, the following key characteristics stand out:
-    1. (a)
-      Data Storage and Processing
-      The cloud infrastructure serves as the storage solution for the extensive data generated by IoT devices, addressing the impracticality of storing such massive volumes within individual systems.
-Additionally, data processing takes place within this stage.
-    2. (b)
-      Scalability
-      Cloud infrastructures, including compute, storage, and networking resources, can effortlessly handle thousands of devices.
-    3. (c)
-      Flexibility
-      Cloud-based IoT offers considerable flexibility, allowing devices to be added or removed as needed without requiring a complete system reconfiguration.
-    4. (d)
-      Maintenance
-      In cloud-based IoT, network devices are supervised and managed by the cloud service provider.
-    5. (e)
-      Cost Effective
-      Cloud usage is cost-effective as users only incur charges for the resources they actively utilize.
+    Cloud: The cloud serves as the central hub of an IoT network, taking on pivotal roles in data processing, storage, and management.
+The key characteristics of the cloud include the ability to store and process extensive data generated by devices, scalability to effortlessly handle thousands of devices, flexibility by allowing devices to be added or removed as needed without requiring a complete system reconfiguration, supervision and management by the cloud service provider, and cost-effectiveness.
     While cloud services are not mandatory for IoT, the recent shift toward edge and fog computing empowers local data processing.
@@
   3. 3.
-    Analytics
-    This represents one of the most intricate and vital layers within IoT.
-It involves the analysis of data, generating valuable insights through the application of diverse machine learning algorithms and statistical analysis techniques.
+    Analytics: This represents one of the most intricate and vital layers within IoT.
+It involves the analysis of data, generating valuable insights through the application of diverse machine learning (ML) algorithms and statistical analysis techniques.
 Numerous applications of analytics in IoT encompass anomaly detection, environmental monitoring, energy management, smart cities, and agriculture [29].
+Figure 1.1: IoT characteristics
 Interpretation and Visualization Tools
@@
 Receiving Automatic Notification, Monitoring Information Proactively, and Controlling the System Remotely are some common examples for user interfaces in IoT systems [30].
-1.2.3 Characteristics of IoT
-Characteristics pertain to the distinct qualities or capabilities of a system that set it apart from others.
-When it comes to IoT, providing a fixed set of characteristics can be challenging due to the evolving nature of IoT devices and their dependency on the specific application domain.
-Nevertheless, there are several common traits that help differentiate IoT technology from other systems.
-This section describes the characteristics of IoT system [31, 32].
-Fig.
-1.1 provides a visual representation of the characteristics of IoT.
-Figure 1.1: IoT characteristics
+1.2.3 Characteristics
+This segment essentially serves as the user interface (UI).
+The UI offers a platform for users to interact directly with the application or system, facilitating communication.
+A user interface is not always reliant on a screen.
+For instance, a TV remote utilizes a user interface comprising multiple buttons, while devices like the Amazon Echo respond to voice commands for control.
+Receiving automatic notifications, monitoring information proactively, and controlling the system remotely are some common examples of user interfaces in IoT systems [30].
+The IoT characteristics discussed in this chapter are illustrated in Figure 1.1.
   1. 1.
-    Connectivity
-    Connectivity is the most vital requirement of IoT.
-The main aspect of Internet of Things is a network with millions of devices connected to each other.
+    Connectivity: Connectivity is the most vital requirement of IoT.
+The main aspect of IoT is a network with millions of devices connected to each other.
 The connectivity remains constant, allowing anyone from anywhere to connect to the IoT network at any given moment.
   2. 2.
-    Intelligent decision making capability
-    The extraction of knowledge from the data generated is highly significant.
+    Intelligent decision making capability: The extraction of knowledge from the data generated is highly significant.
 Consider a sensor that produces data; however, the true value of that data lies in its proper interpretation.
@@
   3. 3.
-    Dynamic and Self Adapting
-    IoT devices should have the ability to adapt to changes in context, their surrounding environment, and the existing situation.
+    Dynamic and Self Adapting: IoT devices should have the ability to adapt to changes in context, their surrounding environment, and the existing situation.
 For instance, within a surveillance system, cameras are capable of switching between day and night modes or adjusting their resolution in response to motion detection, demonstrating their adaptability.
   4. 4.
-    Unique Identity
-    Every IoT device should have a unique identity and unique identifier.
+    Unique Identity: Every IoT device should have a unique identity and unique identifier.
 IoT device interfaces enable users to inquire about device information, monitor their status, and remotely manage them.
@@
   5. 5.
-    Self Configuring
-    IoT devices possess the capability to autonomously update their systems in response to the situation, eliminating the need for user intervention.
+    Self Configuring: IoT devices possess the capability to autonomously update their systems in response to the situation, eliminating the need for user intervention.
 Moreover, they exhibit flexibility in network management, allowing new devices to seamlessly join the network and permitting any device to depart from the network at any time.
   6. 6.
-    Scalability
-    The IoT network is experiencing a continual growth in the number of connected devices, resulting in a substantial and continuous generation of data.
-Consequently, scalability emerges as the foremost feature for any IoT system.
-    Figure 1.2: Advantages of Internet of Things
+    Scalability: The IoT network is experiencing a continual growth in the number of connected devices, resulting in a substantial and continuous generation of data.
+Consequently, scalability emerges as the foremost feature of any IoT system.
+    Figure 1.2: Advantages of IoT
   7. 7.
-    Interoperability
-    IoT devices rely on standardized protocols and technologies to guarantee seamless communication among themselves and with other systems.
+    Interoperability: IoT devices rely on standardized protocols and technologies to guarantee seamless communication among themselves and with other systems.
 Interoperability represents a fundamental block of the IoT, signifying the capacity for various IoT devices and systems to interact and share data, irrespective of the underlying technology or the manufacturer.
-Hence, IoT devices employ standardized protocols, data formats and technologies to uphold interoperability.
+Hence, IoT devices employ standardized protocols, data formats, and technologies to uphold interoperability.
   8. 8.
-    Heterogeneity
-    The devices within an IoT network exhibit heterogeneity, showcasing the network’s capability to accommodate diverse elements.
+    Heterogeneity: The devices within an IoT network exhibit heterogeneity, showcasing the network’s capability to accommodate diverse elements.
   9. 9.
-    Energy Efficiency
-    Energy efficiency stands out as a significant characteristic of IoT.
+    Energy Efficiency: Energy efficiency stands out as a significant characteristic of IoT.
 Numerous IoT devices are purposefully crafted to minimize energy consumption and create low-power devices.
@@
   10. 10.
-    Security
-    The presence of millions of internet-connected devices and the vast data generation underscore the vulnerability of the IoT network to security threats.
-Hence, safety and security emerge as pivotal characteristics of IoT.
+    Security: The presence of millions of internet-connected devices and the vast amount of data generated underscore the vulnerability of the IoT network to security threats.
+Hence, safety and security emerge as pivotal characteristics of the IoT.
 Ensuring safety is of paramount importance to preserve the effectiveness of IoT’s advantages, including efficiency and novel experiences.
 1.3 Advantages of IoT
-The amount of benefit IoT offers to people’s life is the main reason IoT is becoming more popular by the day.
-IoT solutions are developed and invented to make people’s life easier and more convenient.
-Healthcare, education, and business—there is hardly any field that is not affected by IoT technology.
-This section elaborates on the most visible benefits of IoT in people’s daily lives [35, 36] and Fig.
+The amount of benefits IoT offers to people’s lives is the main reason IoT is becoming more popular by the day.
+IoT solutions are developed and invented to make people’s lives easier and more convenient.
+IoT technology affects almost every field, including healthcare, education, and business.
+This section elaborates on the most visible benefits of IoT in people’s daily lives [35, 36], and Fig.
 1.2 provides a visual overview.
   1. 1.
-    Efficient Data Collection
-    IoT based data collection has been particularly beneficial in sectors such as healthcare and finance [37].
+    Efficient Data Collection: IoT-based data collection has been particularly beneficial in sectors such as healthcare and finance [37].
 A straightforward illustration of efficient data collection is the integration of IoT in the retail sector.
@@
   2. 2.
-    Control and Automation
-    IoT has provided its customers a more convenient lifestyle and allows them to control their daily activities with the touch of a button.
+    Control and Automation: IoT has provided its customers with a more convenient lifestyle and allows them to control their daily activities with the touch of a button.
 One simple yet remarkable example is the smart bulb, which can be controlled without even touching the switches; the user can simply turn off or on the light remotely.
@@
   3. 3.
-    Real-time Access to Information
-    IoT devices offer immediate access to information, proving invaluable across healthcare, business, and everyday applications.
+    Real-time Access to Information: IoT devices offer immediate access to information, proving invaluable across healthcare, business, and everyday applications.
 A prime illustration of the advantages of real-time data access is within the healthcare sector.
@@
   4. 4.
-    Improved Efficiency
-    IoT systems operate autonomously, a valuable asset in various domains.
+    Improved Efficiency: IoT systems operate autonomously, which is a valuable asset in various domains.
 Reduced human intervention leads to increased efficiency and decreased labor reliance.
@@
   5. 5.
-    Improved Quality of Life
-    The advent of IoT has significantly improved the lives of its users in numerous ways.
+    Improved Quality of Life: The advent of IoT has significantly improved the lives of its users in numerous ways.
 Real-time health monitoring, including devices like blood pressure monitors and fitness trackers, empowers users to maintain their well-being effectively.
@@
 These advantages extend beyond individuals and can benefit entire industries or communities.
-Smart devices, linked not only to intelligent traffic lights but also road safety monitors and toll gates, can provide drivers with real-time information about road conditions on their route.
+Smart devices, linked not only to intelligent traffic lights but also to road safety monitors and toll gates, can provide drivers with real-time information about road conditions on their route.
   6. 6.
-    Cost and Time Effective
-    IoT minimizes human effort and relies heavily on real-time data transmission, leading to time savings.
+    Cost and Time Effective: IoT minimizes human effort and relies heavily on real-time data transmission, leading to time savings.
 For example, real-time patient monitoring benefits both patients and doctors by eliminating the need for physical meetings, thus saving time for both parties.
@@
   7. 7.
-    Asset Tracking
-    This process involves tracking products within a business or logistics management system.
+    Asset Tracking: This process involves tracking products within a business or logistics management system.
 Manual asset tracking is labor-intensive and time-consuming, but it can be streamlined through the application of IoT technologies like barcodes and RFID tags.
@@
   8. 8.
-    Data Analysis for Research Purpose
-    The enormous amount of data collected from the IoT devices has become a blessing for the researchers in various fields like healthcare, education, business etc.
-The healthcare researchers can use the data collected via biosensors and utilize them in inventing cures and vaccines for disease, the finance industry can utilises the data to understand the trend and improve customer experience, the super shops can analyze the customer behaviour and improve their business and so on.
+    Data Analysis for Research Purpose: The enormous amount of data collected from IoT devices has become a blessing for researchers in various fields like healthcare, education, business, etc.
+The healthcare researchers can use the data collected via biosensors to invent cures and vaccines for disease; the finance industry can use the data to understand trends and improve customer experience; the super shops can analyze customer behavior and improve their businesses; and so on.
   9. 9.
-    Big Data and Predictive Analysis
-    “Big data” has been a widely recognized term in the world long before the emergence of IoT.
+    Big Data and Predictive Analysis: “Big data” has been a widely recognized term in the world long before the emergence of IoT.
 It involves the collection and analysis of massive volumes of data.
 One of the primary objectives of IoT is to amass data from diverse sources, sending this information back to systems for analysis.
-Effective analysis of big data can yield valuable insights, spanning from stock market predictions to understanding customer behavior, and thereby enhancing the business landscape.
+Effective analysis of big data can yield valuable insights, spanning from stock market predictions to understanding customer behavior, thereby enhancing the business landscape.
   10. 10.
-    Increased Productivity
-    IoT utilization in both industry and homes has the potential to boost productivity significantly.
+    Increased Productivity: IoT utilization in both industry and homes has the potential to boost productivity significantly.
 For instance, in a smart home, users can streamline various household tasks using voice commands, enabling efficient multitasking.
@@
   11. 11.
-    Safety and Security
-    The incorporation of IoT offers users a means of security not only in their homes but also in their businesses, schools, offices, and virtually anywhere.
-Individuals can remotely monitor their valuable assets such as vehicles and other important items.
+    Safety and Security: The incorporation of IoT offers users a means of security not only in their homes but also in their businesses, schools, offices, and virtually anywhere.
+Individuals can remotely monitor their valuable assets, such as vehicles and other important items.
 Parents can even keep track of their children’s whereabouts from their workplaces, providing peace of mind.
@@
   12. 12.
-    Enhance Customer Engagement
-    The IoT offers several avenues for enhancing customer engagement.
+    Enhance Customer Engagement: The IoT offers several avenues for enhancing customer engagement.
 It achieves this by leveraging valuable customer data, personalizing experiences, improving convenience, and enabling real-time interactions.
   13. 13.
-    Efficient Resource Utilization
-    IoT facilitates efficient resource utilization through various mechanisms and capabilities.
+    Efficient Resource Utilization:IoT facilitates efficient resource utilization through various mechanisms and capabilities.
 Its ability to collect and monitor real-time information allows organizations to track the status of their resources, such as equipment and machinery, enabling the identification of inefficiencies.
-Furthermore, in the industrial sector, predictive maintenance powered by IoT can anticipate machinery failures, reducing downtime and optimizing resource allocation for maintenance.
+Predictive maintenance powered by IoT in the industrial sector can anticipate machinery failures, reducing downtime and optimizing resource allocation for maintenance.
   14. 14.
-    Improved Technology
-    The innovation brought about by IoT leads to the creation of newer and more advanced technologies in the market.
+    Improved Technology: The innovation brought about by IoT leads to the creation of newer and more advanced technologies on the market.
 For instance, consider the scenario of an air conditioner that was initially controlled manually with a remote.
@@
 This cycle of innovation in response to IoT advancements contributes to the continuous progression of technology [37].
-Along with the generic benefits that apply to practically all fields that are integrating IoT, various domain-specific benefits are also apparent in many sectors.
-Some of the domains are described below.
-Healthcare
-The healthcare industry is one of the most benefited by IoT among all the sectors. The innovation of IoT has revolutionized the healthcare industry by various means. Here are some common examples [6]:
-  1. 1.
-    Remote Medical Assistance
-    The idea of remote monitoring has brought significant advantages to the healthcare sector.
-Doctors can now observe their patients’ conditions from their offices without the need for physical visits, just by using the sensors attached to the patients’ bodies.
-  2. 2.
-    Improved Treatment
-    Through the utilization of IoT technology, physicians can make well-informed decisions based on evidence and ensure full transparency.
-  3. 3.
-    Faster Disease Diagnosis
-    Remote and continuous monitoring allows doctors to identify diseases in their early stages, which can save many lives.
-  4. 4.
-    Emergency monitoring and reporting
-    During remote monitoring, if any unusual or emergency situations arise, such as heart failure or asthma attacks, doctors can quickly provide emergency treatment to the patient.
-Agriculture
-The use of IoT in agriculture brings about numerous advantages. This section highlights some of the most noticeable benefits [40].
-  1. 1.
-    Community Farming
-    The implementation of IoT can facilitate the development and advancement of community farming.
-Through IoT, it becomes feasible to support services that enable shared data storage, the exchange of data and information, and the enhancement of communication among farmers and agriculture experts [41].
-Additionally, IoT facilities or mobile phones can be used to share various agricultural equipment within the community, either for free or through paid services.
-  2. 2.
-    Safety Control and Fraud Prevention
-    In agriculture, the challenge extends beyond just increasing production; it also involves ensuring the production of safe and nutritious food.
-There have been instances of food fraud, including adulteration, counterfeiting, and artificial enhancements, which pose significant health risks and negative economic impact [42].
-Researchers in [43] have explored various aspects of food fraud, including product integrity, process integrity, people integrity, and data integrity, which can be effectively addressed through IoT technology.
-Additionally, IoT can be integrated into both logistics traceability and the qualitative tracking of food [44].
-  3. 3.
-    Wealth Creation and Distributions
-    The adoption of IoT will uncover novel business models, allowing individual farmers to directly connect with customers, eliminating the need for intermediaries.
-Consequently, this will lead to increased profits for the farmers.
-  4. 4.
-    Cost Reduction and Wastage
-    A significant benefit of IoT is its capability for remote monitoring of devices and equipment.
-Farmers can oversee extensive fields from a distance, eliminating the need for physical visits and saving time.
-Additionally, precise knowledge of when and where to apply pesticides and fertilizers will reduce costs and minimize material wastage [45].
-  5. 5.
-    Operational Efficiency
-    IoT technology brings operational efficiency not only to farmers but also to decision-makers in the agriculture sector, including government and non-governmental agencies.
-By collecting data through IoT-based agriculture surveillance, this information can guide various interventions such as disease prevention, wildfire management, compensation programs, and resource allocation.
-    Farmers can leverage IoT and data analytics for making informed decisions about farm management and processes.
-Automated health monitoring of livestock or crops enables efficient diagnosis and prescription of medicine by veterinary or agriculture officers, reducing losses.
-Additionally, IoT optimizes the agri-food supply chain, ensuring real-time balance between demand and supply.
-  6. 6.
-    Awareness
-    IoT is anticipated to create cost-effective mobile apps and wireless network services for the agriculture industry.
-This development can make information about product markets, prices, government services, and regulatory standards for various farm produce easily accessible.
-As a result, people can receive notifications when fresh products hit the market and connect with farmers interested in organic foods.
-  7. 7.
-    Asset Management
-    IoT simplifies asset management by enabling remote monitoring of farm equipment, preventing theft, and issuing alerts for necessary maintenance.
-Education
-The educational sector can derive similar benefits from the utilization of IoT, much like the advantages observed in other industries such as healthcare or retail, as previously described [34].
-Some common examples are presented below.
-  1. 1.
-    Improvement in Learning Environment
-    IoT integration has the potential to revolutionize the learning experience for all students.
-One innovative approach involves attaching RFID tags to various objects, allowing students to interact with these objects and instantly access relevant information.
-This not only promotes greater student engagement but also makes learning a more enjoyable and interactive process.
-  2. 2.
-    Distance Learning
-    IoT has the potential to enhance distance learning, enabling students facing physical classroom attendance challenges, whether due to health concerns or other reasons, to actively engage in virtual classrooms.
-  3. 3.
-    Enhanced Assistant for Special Students
-    IoT makes learning more accessible for special students who require additional support and assistance.
-As referenced in Selinger et al.
-[46] and Lutz et al.
-[47], IoT is confident to offer minority and disabled learners the opportunity for high-quality education and peer-to-peer interaction.
-It will also facilitate student-to-teacher engagement, ultimately enhancing their prospects for greater academic success.
-Warehouse
-The implementation of IoT in warehouse management and processing has paved the way for smart warehousing.
-As a result, big industries like Alibaba, Amazon, and DHL have already integrated this technology into their inventory, logistics, and warehouse management systems.
-Here are descriptions of several common benefits of smart warehousing:
-  1. 1.
-    Efficiency and Transparency
-    The primary and most essential requirement in inventory and warehouse management is to ensure efficiency and transparency across the entire process.
-Barreto et al.
-[48] have presented the key aspects of Logistics 4.0 where they propose the use of an intelligent system and a database to facilitate data sharing through an IoT platform.
-This interconnected approach integrates all warehouse operations and workforces, leading to increased transparency in the process and enhanced system efficiency.
-  2. 2.
-    Efficient Order Monitoring
-    Implementation of IoT can help in monitoring the order without any human intervention.
-The authors of [49] has created a system that employs the following components: (1) Bluetooth tags for transmitting relevant product data; (2) a receiver for monitoring the movement of objects and the position of pallets; and (3) an integrated software package and database server for warehouse monitoring and data exchange.
-This developed system significantly enhances the efficiency of stock management, delivery timing, tracking, and product management.
-  3. 3.
-    Ensure Quick Tally
-    In the traditional tally process, which relies entirely on manual labor, the task of updating product numbers and information is labor-intensive.
-However, through IoT integration, the tally process can be significantly expedited by simply scanning the barcodes on the products [50].
-  4. 4.
-    Enhance Warehouse Automation Management
-    IoT technology can enable warehouse workers to inspect storage boxes without the need for physical contact, a departure from the traditional method of physically touching each box to check its contents.
-Supply Chain Management
-The supply chain represents a intricate logistics system comprising services that transform raw materials into finished products and subsequently deliver them to customers.
-Furthermore, Ensuring a continuous and on-demand supply of products is a crucial market requirement.
-IoT introduces groundbreaking advantages in supply chain management, some of which are detailed below [38].
-  1. 1.
-    Decision Making
-    In the field of logistics, the frequency of information flow and the need for instant decision-making surpasses that of most other industries due to its unique nature.
-Sensors integrated with products facilitate the real-time and continuous generation of data, simultaneously allowing the analysis of this data to predict market trends.
-This predictive capability aids in making efficient decisions across the entire supply chain.
-  2. 2.
-    Smart Transportation
-    The integration of IoT technology facilitates intelligent transportation solutions, encompassing the monitoring of delivery vehicles, tracking vehicle locations, and promoting safe driving practices.
-By employing sensors connected to all delivery vehicles, real-time tracking of vehicle locations becomes possible, and notifications can be dispatched in the event of accidents.
-Additionally, these sensors can monitor factors like excessive vehicle speed, fuel tank levels, and tire pressure to ensure safe and efficient transportation.
-Moreover, by analyzing traffic conditions and road conditions, the system can send notifications to the driver regarding the status of the road leading to their destination.
-Additionally, the system can suggest alternative routes that promote safe driving.
-  3. 3.
-    Smart Loading and Unloading
-    Efficiently handling the loading, unloading, and stacking of products is crucial for item distribution.
-It is essential to achieve this with cost-efficiency and timeliness.
-Automation lifts equipped with RFID tags can streamline this process by precisely locating and delivering goods without requiring human intervention.
-Furthermore, robots can be employed as warehouse workers, leading to a reduction in human labor [51].
-  4. 4.
-    Smart Carriage
-    Integrating Automatic Guided Vehicles (AGV) into both internal and external transportation systems can lead to cost savings and time efficiency gains [52].
-Various communication technologies facilitate internal communication among these vehicles, contributing to more efficient management processes as a result [53].
-  5. 5.
-    Smart Packaging
-    IoT has seamlessly replaced traditional packaging processes with intelligent and smart packaging solutions.
-These smart packages are capable of making real-time decisions, and product information is easily accessible through smart labels affixed to the items [54].
-In [55], authors have employed a similar process to categorize products based on their temperature sensitivity.
-Furthermore, to ensure the safe delivery of products, an automatic reminder system have been proposed that activates if the products deviate from the preset conditions.
-  6. 6.
-    Smart Distribution
-    In the logistics and warehouse management of product distribution, several stages are involved both before and during the process.
-Before the final goods are dispatched, there is a crucial step of labeling products according to their types and ensuring proper label attachment.
-This task is efficiently handled by an autonomous wheeled robot [56].
-In specific industries like the food sector, smart labeling tags equipped with microcontrollers are employed to effectively track food items [57].
-IoT plays a significant role in various aspects of the distribution process, including route planning, item location, sorting, and labeling.
 1.4 IoT Architecture
 Architecture refers to a structured framework depicting the tangible components of a network, their operational arrangement and setup, underlying principles and structures, as well as the manner in which data is organized and utilized within its functioning.
-IoT architecture comprises of a collection of devices, sensors, actuators, end users, cloud services and most importantly various communication layers and IoT protocols.
-Owing to the dynamic characteristics of the IoT domain and the crucial importance of sensors, it is impossible to identify one fundamental and universally agreed architecture for diverse IoT domains.
-Nevertheless, certain foundational elements will exhibit similarity across various projects.
-The architecture of the IoT has been categorized based on a range of factors.
-Certain researchers have undertaken a division grounded in distinct layers, while others have pursued a division predicated on the specific objectives of the IoT project.
-However, there are some fundamental blocks present in more or less in every IoT architecture which provides scalability and flexibility in IoT architecture [58].
-  1. 1.
-    Sensing Layer
-    The primary function of this layer involves real-time detection of changes in the connected devices.
-This layer comprises two main actors: sensors and actuators.
-Sensors are responsible for acquiring diverse data types from the environment, including but not limited to temperature, humidity, caloric information, heart rate, and object identification.
-Meanwhile, actuators are mechanized components tasked with device manipulation, encompassing actions such as activating and deactivating door lights and similar functionalities.
-  2. 2.
-    Communication Layer
-    Functioning as an intermediary system bridging the sensing layer with the upper layers, this layer manages connectivity, facilitates data transmission between the sensing and cloud layers, or upper layers, and conducts message routing among remote devices.
-  3. 3.
-    Cloud Layer
-    Situated at the core of the IoT architecture, this layer functions as the central processing unit.
-It receives data collected from the sensing layer to conduct storage, processing, and analysis procedures.
-Within this framework, the cloud employs a data centre as a pivotal server for tasks like data refinement by eliminating redundant and noisy information, data format conversion, and more.
-Nonetheless, the advancement of cloud computing research is progressively shifting towards the decentralisation of multiple processing tasks from the cloud to edge nodes that are currently concentrated in data centres [59].
-  4. 4.
-    Management Layer
-    Tasked with monitoring and administering the entire IoT system, this layer handles responsibility for device configuration and firmware updates, as well as providing analytics and generating reports.
-These capabilities empower system administrators to comprehensively grasp the system’s dynamics and find opportunities for enhancement.
-  5. 5.
-    Services and Application Layer
-    This layer provides value to the end user of an IoT system. This layer utilizes processed data and provide various functionalities or services.
-  6. 6.
-    Security Layer
-    With the responsibility of ensuring security and safeguarding against unauthorized access, this layer bears the duties of authentication and authorization, intrusion detection, threat mitigation, and related responsibilities [60].
+IoT architecture comprises a collection of devices, sensors, actuators, end users, cloud services, and most importantly, various communication layers and IoT protocols.
+All IoT systems inherently follow the generic three-layer architecture.
+However, based on necessities or specific application requirements, the generic model can be modified by adding extra layers, thus forming four- or five-layer architectures [4].
 Figure 1.3: Generic architecture of IoT
 The IoT architecture can be divided in two ways: (1) layered architecture and (2) domain-specific architecture.
-This section presents several existing architectures for IoT.
-However, detailed explanations are beyond the scope of this chapter; instead, it provides an overview of a generalized architecture.
-1.4.1 Generic Architecture of IoT
+In this section, we have presented a generic architecture of IoT.
 IoT follows a layered architecture, which refers to the structured framework used to design and organize the various components and functions of an IoT system.
-This architecture is composed of multiple layers, each with its specific role and responsibilities, allowing for efficient communication, data processing, and management within the IoT ecosystem.
-The typical layers in IoT architecture include the perception layer, network layer and application layer.
+This architecture is composed of multiple layers, each with its own specific role and responsibilities, allowing for efficient communication, data processing, and management within the IoT ecosystem.
+The typical layers in the IoT architecture include the perception layer, network layer, and application layer.
 These layers collaborate to ensure the smooth operation of IoT devices, data transmission, and the provision of IoT services to end users.
-According to this generalized architecture, also known as the three layer architecture, the IoT system is divided into three layers [61] namely (i) application layer, (ii) network layer, and (iii) perception layer.
+According to this generalized architecture, also known as the three-layer architecture, the IoT system is divided into three layers [61], namely (i) the application layer, (ii) the network layer, and (iii) the perception layer.
 Every one of these layers possesses inherent security challenges [62].
-A visual representation of the Three Layer Architecture is featured in Fig.
+A visual representation of the generic three-layer architecture is featured in Fig.
 1.3.
@@
   1. 1.
-    Perception Layer
-    The perception layer, also known as the sensor layer is the foundational layer of IoT architecture [61].
+    Perception Layer: The perception layer, also known as the sensor layer, is the foundational layer of IoT architecture [61].
 This layer engages with smart devices, including but not limited to smartwatches and smart rings, employing an array of sensors and actuators.
@@
   2. 2.
-    Network Layer
-    The network layer, also known as the transmission layer is the middle layer of IoT architecture [63].
-This layer is responsible for receiving the information passed from the perception layer and determine the routes to transmit the processed data to various connected IoT devices and applications using integrated network such as wired or wireless secure connection.
-Network layer is core layer of IoT three layer architecture as this uses various devices such as routing devices, gateways, switches and hub and operate them by using various communication technologies such as WiFi, Bluetooth, 3G, LTE, Zigbee etc.
+    Network Layer: The network layer, also known as the transmission layer, is the middle layer of the IoT architecture [63].
+This layer is responsible for receiving the information passed from the perception layer and determining the routes to transmit the processed data to various connected IoT devices and applications using integrated networks such as wired or wireless secure connections.
+The network layer is the core layer of the IoT three-layer architecture, as it uses various devices such as routing devices, gateways, switches, and hubs and operates them by using various communication technologies such as WiFi, Bluetooth, 3G, LTE, Zigbee, etc.
 In summary, the network layer is responsible for transmitting data to and from several applications through interfaces and gateways using multiple communication technologies and protocols.
   3. 3.
-    Application Layer
-    Serving as the uppermost tier within the IoT architecture, denoted as the application layer or business layer, as referenced in [58], this layer is tasked with the aggregation of data from the network layer, thereby striving to attain the objective of establishing an smart environment, the ultimate aim of the IoT paradigm.
-This layer accommodates a diverse array of applications, each characterised by their own requirements.
+    Application Layer: Serving as the uppermost tier within the IoT architecture, denoted as the application layer or business layer, as referenced in [58], this layer is tasked with the aggregation of data from the network layer, thereby striving to attain the objective of establishing a smart environment, the ultimate aim of the IoT paradigm.
+This layer accommodates a diverse array of applications, each characterized by their own requirements.
 Examples of such applications are smart grids, smart cities, and smart transportation, as elaborated in [62].
@@
 Figure 1.4: Visual representation of Fog vs Cloud
-The three layer architecture is the generalized and most common architecture and several systems have integrated this architecture [64].
-Although this multi layer architecture seems to be simple at first glance, but the functionalities of the network and application layer might get complex at times.
-For example, the network layer is not only responsible for data transmission but also provides data services such as data aggregation and processing etc.
-On the other hand, the application layer is not solely responsible for providing service to customers and users but also provides data analytic and conducts data mining etc.
+The three-layer architecture is the generalized and most common architecture, and several systems have integrated this architecture [64].
+Although this multi-layer architecture seems simple at first glance, the functionalities of the network and application layers might get complex at times.
+For example, the network layer is not only responsible for data transmission but also provides data services such as data aggregation and processing, etc.
+On the other hand, the application layer is not solely responsible for providing service to customers and users but also provides data analysis, conducts data mining, etc.
 Therefore, in response to specific requirements, additional layers have been incorporated, building upon the fundamental layers.
-For instance, the four-layered or the five-layered architectures enhance the system’s flexibility.
+For instance, the four-layered or five-layered architectures enhance the system’s flexibility.
 Nonetheless, the three-layer architecture serves as the foundation for all these variations.
 Furthermore, new-generation applications require shorter response times and low energy consumption as IoT devices have limited capacity [65, 66].
-Therefore, researchers have utilized fog and cloud layers which are visualized in Fig.
+Therefore, researchers have utilized fog and cloud layers, which are visualized in Fig.
 1.4.
 1.5 Key Technologies of IoT
-To function properly, the Internet of Things requires a variety of technologies.
-Different technologies are deployed in various layers of an IoT architecture, and there are many different sorts of technologies, including hardware technology, software technology, and, most crucially, communication technology.
-This section discusses the many IoT technologies that are used to ensure the successful operation of an IoT system.
+IoT requires a variety of technologies to operate, which are deployed in various layers of an IoT architecture, and there are many different sorts of technologies, including hardware technology, software technology, and, most crucially, communication technology.
+This section discusses crucial IoT technologies that are used to ensure the successful operation of an IoT system.
 Fig.
-1.5 depicts the taxonomoy of IoT technologies.
+1.5 depicts the taxonomy of IoT technologies.
 Figure 1.5: Technology Taxonomy of IoT
 1.5.1 Hardware Platforms
-The major components of an IoT system are devices attached with sensors or wearable devices which are used to data collection.
+The major components of an IoT system are devices attached to sensors or wearable devices that are used for data collection.
 Therefore, various types of hardware platforms are used in order to build these sensor devices.
-Several key points need to be considered before selecting the hardware platforms such as what is the purpose of the IoT device, what kind of connectivity it requires and so on.
-Two most accessible and popular hardware platforms are Raspberry Pi and Arduino.
-Raspberry Pi is a Linux based single board computer whereas Arduino makes different microcontrollers, open-source development tools, kits, and software.
-Both of them have strong data acquisition, processing ans storage capabilities and and provide both wireless and wired connectivity.
-However, in terms of power management Arduino is superior than Raspberry Pi because Raspberry Pi does not contain sleep or suspend modes for power utilization but Arduino does [67].
-Intel has developed several development boards named Intel Galileo Gen and Intel Edison which uses Arduino IDE as development environment.
-Beagle Bone Black is a small sized Linux based board designed to build powerful IoT devices.
-Other than these platforms there are several other hardware platforms present in the markets such as Electric Imp 003, ARM mbed NXP LPC1768, WiPy and LoPy boards by Pycom company and many more [32].
+Several key points need to be considered before selecting the hardware platforms, such as the purpose of the IoT device or the type of connectivity it requires.
+The two most accessible and popular hardware platforms are the Raspberry Pi and Arduino.
+Both of them have strong data acquisition, processing, and storage capabilities and provide both wireless and wired connectivity.
+However, in terms of power management, Arduino is superior to Raspberry Pi because Raspberry Pi does not contain sleep or suspend modes for power utilization, while Arduino does [67].
+Intel Galileo Gen and Intel Edison are examples of using an Arduino IDE.
 1.5.2 Wireless Communication Technology
-Given the vast number of devices in the current IoT network and the expectation of diverse connected devices in future IoT applications, there’s a urge to develop various technologies to facilitate their connectivity.
-This subsection explores the existing wireless technologies designed for IoT connectivity and categorize them into three groups.
+Given the vast number of devices in the current IoT network and the expectation of diverse connected devices in future IoT applications, there’s an urge to develop various technologies to facilitate their connectivity.
+This subsection explores the existing wireless technologies designed for IoT connectivity and categorizes them into three groups.
 Table 1.2: Comparison between popular short-range technologies based on various parameters.
-(Abbreviations: CSMA/CA (Carrier Sense Multiple Access with Collision Avoidance), CSMA/CD (Carrier Sense Multiple Access with Collision Detection), FHSS (Frequency Hopping Spread Spectrum), O-QPSK (Offset Quadrature Phase Shift Keying), OOK (On-Off Keying), PPM (Pulse Position Modulation), LBT (Listen Before Talk)
 OWC
@@
 200 meters
-Energy consumtion
+Energy consumption
 Classic: High
@@
 The table summarizes the technical specifications of Bluetooth, ZigBee, Wi-Fi, LR-WPAN, VLC, and BS-ILC [68, 32].
-Bluetooth, ZigBee, LR-WPAN, and BS-ILC offer cost-effective communication solutions, whereas the overall cost of Wi-Fi and VLC may vary depending on usage.
-In terms of frequency bands, most technologies operate around 2.4 GHz, except for VLC, which utilizes a much larger frequency band.
-The frequency for BS-ILC varies by region.
-VLC and BS-ILC prioritize achieving high transmission rates of approximately 100 Gbps, while Wi-Fi transmission rates can vary across different standards.
-Notably, VLC and BS-ILC exhibit very low energy consumption, as indicated in the table.
-Long Range Technologies
-Long range technology in contrast with short range technologies refers to technologies that enable communication over extended distances.
-These technologies are designed to transmit data or signals over considerable ranges, typically much farther than short-range technologies like Bluetooth or Wi-Fi.
-Long-range technologies are commonly used in various applications such as long-distance wireless communication, remote monitoring, tracking, and more.
-They are essential for scenarios where data needs to be transmitted reliably over extensive geographical areas.
-Long-range wireless technologies can be classified into two main categories: Mobile Communication Technology and LPWAN Technologies.
-Table 1.2 presents a detailed overview of various long-range technologies utilized in the IoT environment, organized based on several parameters such as frequency band, data rate, transmission range, energy consumption, cost, and more.
-it provides a summary of the technical specifications for various versions of Mobile Communication Technologies, which include 2G, 3G, 4G, 5G, and WiMax, as well as different versions of LPWAN Technologies, which encompass UNLICENSED LPWAN and LICENSED LPWAN.
-It is evident from the table that when it comes to mobile communication technologies, 2G, 3G, and 4G operate within a frequency band ranging from approximately 850MHz to 2100MHz, while 5G utilizes a much broader frequency band of up to 80GHz.
-Additionally, all five technological variations, 5G offers the highest data rate, exceeding 20Gbps.
-In terms of energy efficiency, both 2G and 5G have low consumption rates, whereas WiMax exhibits high energy consumption.
-However, it’s worth noting that 3G, 4G, and 5G all come with higher implementation costs.
-In the UNLICENSED LPWAN category, the top contenders are LoRa and Sigfox, while the most promising options in the LICENSED LPWAN technology are LTE-M and NB-IoT (Narrowband Internet of Things).
-The key distinction highlighted in the table is that UNLICENSED LPWAN technology operates within unlicensed spectrum resources, leading to lower deployment costs, whereas LICENSED LPWAN technology relies on licensed spectrum resources, resulting in relatively higher expenses for both devices and deployment [68].
-Emerging For Massive Connectivity
-While current wireless IoT technologies have achieved some success in supporting various IoT applications, they still face challenges in meeting the future demands of IoT, where there will be a need to connect hundreds of billions of objects or things.
-For example, handling the connectivity among vast number of IoT devices with limited transmission payload and several random access protocols used by existing technologies often result in significant problems such as frequent access collisions, increased delays, and a high amount of signaling overhead for IoT devices.
-Furthermore, the constrained availability of wireless resources for connecting IoT devices creates shortages and inefficient utilization of these resources.
-Hence, there have been numerous ongoing initiatives to tackle the limitations of current technologies.
-Among these, the solutions CS, NOMA, mMIMO, and ML stand out as the most promising [68].
-These technologies all have the potential to be incorporated into future IoT connectivity standards.
-However, it is essential to acknowledge that there are existing issues and constraints that must be resolved for their effective implementation.
-These emerging technologies offer the capability not only to support massive connectivity but also to deliver high reliability and low-latency transmissions.
-In the future, it is anticipated that the development of even more advanced technologies aimed at addressing critical IoT challenges.
-Simultaneously, efforts should focus on intelligently integrating existing and emerging technologies to unlock their full potential and optimize system performance [68].
 Table 1.3: Comparison between popular long-range technologies based on various parameters.
-(Abbreviations: GSM (Global System for Mobile Communications), UMTS (Universal Mobile Telecommunications System), LTE-A (Long-Term Evolution Advanced), LTE-M (Long-Term Evolution for Machines), 3GPP (3rd Generation Partnership Project), TDMA (Time Division Multiple Access), CDMA (Code Division Multiple Access), FDMA ( Frequency Division Multiple Access), WCDMA (Wideband Code Division Multiple Access), OFDMA (Orthogonal Frequency Division Multiple Access), SC-FDMA (Single Carrier Frequency Division Multiple Access), CP-OFDM (Cyclic Prefix Orthogonal Frequency Division Multiplexing), TDD (Time Division Duplex), FDD (Frequency Division Duplex), ALHOA (Adaptive Link Hopping On Air), QPSK (Quadrature Phase Shift Keying), BPSK (Binary Phase Shift Keying), QAM (Quadrature Amplitude Modulation), GFSK (Gaussian Frequency Shift Keying), DBPSK (Differential Binary Phase Shift Keying), CSS (Chirp Spread Spectrum), OFDM (Orthogonal Frequency Division Multiplexing), GMSK (Gaussian Minimum Shift Keying))
 Mobile Communication Technology LPWAN TECHNOLOGIES
@@
 model
-Moderae to high Moderae to high
+Moderate to high Moderate to high
+Bluetooth, ZigBee, LR-WPAN, and BS-ILC offer cost-effective communication solutions, whereas the overall cost of Wi-Fi and VLC may vary depending on usage.
+In terms of frequency bands, most technologies operate around 2.4 GHz, except for VLC, which utilizes a much larger frequency band.
+The frequency of BS-ILC varies by region.
+VLC and BS-ILC prioritize achieving high transmission rates of approximately 100 Gbps, while Wi-Fi transmission rates can vary across different standards.
+Notably, VLC and BS-ILC exhibit very low energy consumption, as indicated in the table.
+Long Range Technologies
+Long-range technologies are designed to transmit data or signals over considerable distances, typically much farther than short-range technologies like Bluetooth or Wi-Fi.
+Long-range technologies are commonly used in various applications, such as long-distance wireless communication, remote monitoring, tracking, and more, and are essential for scenarios where data needs to be transmitted reliably over extensive geographical areas.
+These technologies can be classified into two main categories: mobile communication technology and LPWAN technologies.
+Table 1.3 presents a detailed overview of various long-range technologies utilized in the IoT environment.
+It provides a summary of the technical specifications for various versions of mobile communication technologies as well as different versions of LPWAN technologies.
+It is evident from the table that when it comes to mobile communication technologies, 2G, 3G, and 4G operate within a frequency band ranging from approximately 850MHz to 2100MHz, while 5G utilizes a much broader frequency band of up to 80GHz.
+Additionally, of all five technological variations, 5G offers the highest data rate, exceeding 20Gbps.
+In terms of energy efficiency, both 2G and 5G have low consumption rates, whereas WiMax exhibits high energy consumption.
+However, it is worth noting that 3G, 4G, and 5G all come with higher implementation costs.
+In the UNLICENSED LPWAN category, the top contenders are LoRa and Sigfox, while the most promising options in the LICENSED LPWAN technology are LTE-M (Long-Term Evolution for Machines) and NB-IoT (Narrowband IoT).
+The key distinction highlighted in the table is that UNLICENSED LPWAN technology operates within unlicensed spectrum resources, leading to lower deployment costs, whereas LICENSED LPWAN technology relies on licensed spectrum resources, resulting in relatively higher expenses for both devices and deployment [68].
+Emerging For Massive Connectivity
+While current wireless IoT technologies have achieved some success in supporting various IoT applications, they still face challenges in meeting the future demands of IoT.
+For example, handling the connectivity among a vast number of IoT devices with a limited transmission payload and several random access protocols used by existing technologies often results in significant problems such as frequent access collisions, increased delays, and a high amount of signaling overhead for IoT devices.
+Furthermore, the constrained availability of wireless resources for connecting IoT devices creates shortages and inefficient utilization of these resources.
+Hence, there have been numerous ongoing initiatives to tackle the limitations of current technologies.
+Among these, the solutions CS, NOMA, mMIMO, and ML stand out as the most promising [68].
+These emerging technologies offer the capability not only to support massive connectivity but also to deliver high reliability and low-latency transmissions.
+However, it is essential to acknowledge that there are existing issues and constraints that must be resolved for their effective implementation.
+It is anticipated that the development of even more advanced technologies will address critical IoT challenges.
+Simultaneously, efforts should focus on intelligently integrating existing and emerging technologies to unlock their full potential and optimize system performance [68].
 1.5.3 Cloud Solutions
@@
 These services are available on a pay-as-you-use basis, allowing users to pay only for the services they actually use.
-Cloud services are becoming increasingly popular, and cloud-based architectures are discussed in detail in Section 1.4.
 Cloud platforms can be integrated into numerous domains, including healthcare, smart cities, agriculture, education, and supply chain management.
-For simplicity purpose, Table 1.3 presents several popular platforms in the agriculture domain [32].
+For simplicity, Table 1.4 presents several popular platforms in the agriculture domain [32].
 Table 1.4: Popular IoT cloud solutions in agriculture domain
@@
 This category encompasses a range of hardware and software technologies responsible for different aspects of IoT communication channels.
-Several technologies, such as NFC and M2M, are considered as both hardware and software technologies, whereas RFID is classified as a hardware-only technology.
+Several technologies, such as NFC and M2M, are considered both hardware and software technologies, whereas RFID is classified as a hardware-only technology.
 The subsequent sections will provide brief insights into these technologies.
-Hardware Technology
 RFID
-RFID, or Radio-Frequency Identification which has also been introduced in sub section 1.2.2, is a technology that comprises of one or more readers and several RFID tags.
-These tags are small microchips, have unique codes and can be on items like products in a store or even access cards.
-When a device called an RFID reader sends out electromagnetic radio waves, the tags respond with their codes.
+RFID is a technology that consists of one or more readers and several RFID tags.
+These tags are small microchips, have unique codes, and can be on items like products in a store or even access cards.
+When an RFID reader device sends out electromagnetic radio waves, the tags respond with their unique Electronic Product Codes (EPCs) [69].
 The reader captures these codes and sends them to a computer, which can then figure out what the tags belong to.
 This technology is used in various ways, like keeping track of inventory in stores, allowing access to secure areas, or monitoring packages as they move through the delivery process.
-Hitachi has developed a tag with dimensions that comes in three configurations Passive Reader Active Tag (PRAT), Active Reader Passive Tags (ARPT) and Active Reader Active Tag (ARAT).
-An Electronic Product Code (EPC) is a typical piece of information stored in an RFID tag.
-These codes make it possible to uniquely track objects because they are encoded on RFID tags [69].
 Near Field Communication (NFC)
-Near Field Communication (NFC) is a short range wireless connectivity technology that uses magnetic field induction to establish communicating between devices when they are brought within near proximity of each other without the need of prior connection establishment.
-It is an extension of RFID technology with additional and more dynamic features by integrating smartphones.
-Most modern phones are equipped with NFC chips and support applications like Apple Pay and Google Pay.
+NFC, an extension of RFID technology, is a short-range wireless connectivity technology that uses magnetic field induction to establish communication between devices when they are brought within close proximity of each other without the need for prior connection establishment.
+NFC chips are equipped in most modern phones, supporting applications like Apple Pay and Google Pay.
 NFC operates in the unlicensed radio frequency band at 13.56MHz.
 Its typical range is about 20 meters, with the actual distance often determined by the size of the antenna in the device.
-NFC technology is expected to play a vital role in the future of IoT by providing a wireless connection tool to link with other smart objects.
-For instance, using NFC on a mobile device, a user can turn their phone into various objects, like using it as a credit card for transactions [69].
-Machine to Machine (M2M)
-Machine-to-Machine (M2M) communication is gaining popularity and involves direct communication between computers, embedded processors, smart sensors, actuators, and mobile devices.
-It comprises four main components: sensing, heterogeneous access, information processing, and applications & processing.
-In practical terms, M2M functions within a five-part framework:
-  1. 1.
-    M2M Device: This is a device capable of responding to requests for data stored within it.
-  2. 2.
-    M2M Area Network (Device Domain): It provides the connectivity that links M2M Devices with M2M Gateways.
-  3. 3.
-    M2M Gateway: These gateways utilize M2M capabilities to ensure that M2M Devices can interact and connect to the communication network.
-  4. 4.
-    M2M Communication Networks (Network Domain): These networks facilitate communication between the M2M Gateway(s) and M2M applications.
-  5. 5.
-    M2M Applications: This layer serves as middleware, where data passes through various application services and is utilized by specific business processing engines.
-Machine-to-Machine (M2M) technology finds applications across various sectors, including healthcare, smart robotics, cyber transportation systems (CTS), manufacturing systems, smart home technologies, and smart grids [69].
+It is expected that NFC technology will play a vital role in the future of IoT by providing a wireless connection tool to link with other smart objects.
+For instance, using NFC on a mobile device, a user can transform their phone into various objects, such as using it as a credit card for transactions [69].
+M2M
+M2M communication is gaining popularity and involves direct communication between computers, embedded processors, smart sensors, actuators, and mobile devices.
+It comprises four main components: sensing, heterogeneous access, information processing, and applications and processing.
+In practical terms, M2M functions within a five-part framework: a device for responding to requests, gateways to interact and connect, an area network for providing connectivity between devices and gateways, applications serving as middleware, and a communication network to facilitate communication between gateways and applications.
+M2M technology is applied across various sectors, including healthcare, smart robotics, cyber transportation systems (CTS), manufacturing systems, smart home technologies, and smart grids [69].
 Vehicle-to-Vehicle Communications (V2V)
-Vehicle-to-Vehicle Communications (V2V) treats each vehicle as a node and enables wireless data exchange between vehicles regarding their speed, location, and more.
-This network allows vehicles to broadcast and receive messages in all directions.
+V2V Communications treats each vehicle as a node and enables wireless data exchange (omnidirectional) between vehicles regarding their speed, location, and more.
 Vehicles equipped with appropriate software, often referred to as safety applications, can use messages from nearby vehicles to detect potential collision risks in real-time.
@@
 1.6.1 Smart Cities
-The concept of smart cities can be seen as a complex IoT paradigm, where the management of public affairs incorporates the introduction of Information and Communication Technology (ICT) solutions [70].
+The concept of smart cities can be seen as a complex IoT paradigm where the management of public affairs incorporates the introduction of information and communication technology (ICT) solutions [70].
 A smart city utilizes public resources to digitize the city and enhance the quality of life.
@@
 Barcelona has embarked on the CityOS project, with the primary goal of creating a centralized operating system to manage all the smart devices and services available within the city.
-Their focus has predominantly centered on enhancing smart transportation and water systems.
+Their focus has predominantly been on enhancing smart transportation and water systems.
 Similarly, Stockholm has also placed significant emphasis on these two domains and holds the distinction of being one of the pioneering cities to implement the concept of congestion charges, wherein users are imposed fees for entering congested areas [4].
-Smart cities are characterized by technology-based infrastructure and highly efficient smart public services.
-These smart cities are divided into multiple subsystems, and some of them are described below.
-In addition, Table 1.4 summarises related research works in this domain.
+In addition, Table 1.5 summarizes related research works in this domain.
 Table 1.5: Relevant applications in different Smart Cities domain
-Applications Related Works
-Structural health of buildings [72, 70]
-Waste management [70]
-Air quality and noise monitoring [73, 74]
-Smart Transport & Traffic congestion
+Applications Related Works Applications Related Works
+Structural health of buildings [72, 70] Waste management [70]
+Air quality and noise
+monitoring
+[73, 74]
+Smart Transport and
+Traffic congestion
 [4, 75, 76, 77, 78]
-Smart grid [79, 80]
-Smart water systems [81, 4]
+Smart grid [79, 80] Smart water systems [81, 4]
 Structural Health of Buildings
-This service involves continuous monitoring of areas prone to various external factors, and measuring the health of any building.
+This service involves continuous monitoring of areas prone to various external factors and measuring the health of any building.
 IoT sensors connected to the buildings can store information about the building’s strength, which will help analyze how sturdy the building is or if it requires any refinement [72].
-Additionally, various kinds of sensors can be employed depending on the usage, such as sensors for tracking vibrations to measure the stress on the building, temperature and humidity sensors, or other types of atmospheric sensors to assess the level of pollution in an area [70].
+Depending on the usage, one can employ various kinds of sensors, including those for tracking vibrations to measure the stress on the building, temperature and humidity sensors, or other types of atmospheric sensors to assess the level of pollution in an area [70].
 Employing IoT in this area can reduce manual labor, where humans have to manually assess the building’s health and environmental conditions, and it can also reduce the overall cost to a great extent.
@@
 Thus, an efficient waste management system is essential for any society.
-Integrating IoT into waste management offers multiple benefits.
+Integrating the IoT into waste management offers multiple benefits.
 It enables the detection of waste levels and real-time tracking of garbage truck routes, leading to more efficient route planning.
-Moreover, it can streamline manual labor involved in waste separation and monitor the disposal process.
-These tasks are achieved using sensors on garbage vehicles connected to a central software system that analyzes and controls the system based on the data collected by the sensors [70].
+Moreover, it can streamline the manual labor involved in waste separation and monitor the disposal process.
+Sensors on garbage vehicles connected to a central software system achieve these tasks by analyzing and controlling the system based on the data collected [70].
 This approach reduces the cost of manual processes and enhances recycling management.
@@
 Likewise, it is essential to maintain a balanced noise level for all living beings in society, including humans and animals.
-Noise sensors can be used to measure the decibel levels, and the data collected can help the central authority identify noisy areas and regulate noise levels to keep them within acceptable limits.
-Smart Transport & Traffic Congestion
+Using noise sensors to measure decibel levels, the central authority can collect data to identify noisy areas and regulate noise levels to keep them within acceptable limits.
+Smart Transport and Traffic Congestion
 In today’s world, traffic congestion is a widespread issue affecting nearly every country and city, particularly as more than half of the world’s population now resides in urban areas [4].
-To tackle this issue, many cities have adopted IoT solutions to establish smart transportation systems aimed at managing traffic congestion.These initiatives include smart traffic lights and smart parking systems, which collectively enhance transportation capacity and improve safety and speed for travelers.
+To tackle this issue, many cities have adopted IoT solutions to establish smart transportation systems aimed at managing traffic congestion.
+These initiatives include smart traffic lights [4] and smart parking system [77, 78], which collectively enhance transportation capacity and improve safety and speed for travelers.
 The primary advantages of smart transportation systems are reducing traffic congestion, ensuring hassle-free travel, and facilitating easy parking.
 Moreover, these systems enable quicker responses in case of accidents and contribute to accident reduction by effectively managing traffic flow [4].
-A variety of sensor technologies are employed to achieve these objectives, such as accelerometers to measure speed, RFIDs for vehicle identification, GPS sensors for location tracking, gyroscopes for direction detection, and cameras for recording traffic patterns and vehicle movements.
-Practical implementations of these sensors are elaborated upon below.
-  1. 1.
-    Traffic Management and Monitoring Applications
-    In a smart transportation system, vehicles are seamlessly interconnected via wireless networks.
-Each vehicle is equipped with essential IoT sensors, such as RFID and GPS.
-Traffic congestion is identified through the use of GPS and accelerometers, which are capable of tracking vehicle movements across different areas.
-The collected data is then analyzed by mapping systems to pinpoint congested areas.
-This information is subsequently transferred to the driver, providing guidance on selecting the optimal route and avoiding potential traffic jams [4, 75].
-  2. 2.
-    Application to Ensure Safety
-    Numerous IoT applications have been developed with a focus on monitoring driver behavior continuously to enhance road safety.
-This is achieved by detecting signs of drowsiness or fatigue in drivers and providing appropriate recommendations, such as suggesting rest breaks.
-Various IoT applications utilize techniques like eye movement detection, facial expression analysis, and pressure sensing.
-These applications are typically integrated into smartphones, equipped with sensors like gyroscopes, accelerometers, and GPS, which analyze the collected data to make informed decisions.
-Additionally, through remote vehicle monitoring, it becomes feasible to track factors such as speed and tire pressure.
-This information allows centralized systems to make informed decisions aimed at preventing potential accidents [76].
-  3. 3.
-    Parking Guidance and Information System
-    To address vehicle parking challenges, a range of IoT-based applications have been introduced.
-The primary goal of these applications is to assist drivers in locating available parking spaces remotely, eliminating the need to spend valuable time searching for vacant spots.
-This, in turn, contributes to reducing traffic congestion [77].
-Parking guidance and information (PGI) applications employ barcodes to track the number of vehicles entering or leaving a parking area.
-Furthermore, parking zones are equipped with cameras and sensors that detect and monitor vehicles, facilitating the identification of vacant and suitable parking spots for incoming vehicles [78].
-  4. 4.
-    Smart Traffic Lights
-    Smart traffic lights utilize data collected from sensors to make decisions based on the degree of traffic congestion.
-These intelligent systems can serve various purposes, such as prioritizing public transportation vehicles or implementing special controls for emergency service vehicles when needed [4].
-  5. 5.
-    Applications for Detecting Accidents
-    These applications have the capability to identify unusual situations, such as road accidents, through the use of accelerometer sensors or the assistance of acoustic information.
-Additionally, they can transmit images of the surrounding area to the central authority.
-Based on this information, the authority can make informed decisions regarding the extent of medical assistance required.
+These objectives are achieved through the use of a variety of sensor technologies, including accelerometers for measuring speed, RFIDs for vehicle identification, GPS sensors for location tracking, gyroscopes for direction detection, and cameras for recording traffic patterns and vehicle movements.
+Some real-world uses of these sensors can be seen in applications for managing and monitoring traffic [4, 75], applications to ensure safety [76], and application for detecting accidents.
 Smart grid
-A traditional grid system is an electrical grid that includes transmission lines, transformers, and various communication utilities, responsible for delivering electricity from power plants to homes or businesses.
+A traditional grid system is an electrical grid that includes transmission lines, transformers, and various communication utilities responsible for delivering electricity from power plants to homes or businesses.
 One significant limitation of the traditional grid is its one-way communication, which prevents power plants from efficiently responding to increasing power demands from consumers.
-To address this challenge, the smart grid establishes a two-way communication system between utilities and consumers which enables more effective management of economic, sustainable, and secure power resources [79].
-By integrating IoT into grid systems, houses and businesses can be equipped with smart meters that monitor energy generation, storage, and consumption, transmitting this data to the smart grid.
-The smart grid, which allows for two-way communication between utilities and consumers, helps consumers in better understanding the power consumption, reports real-time increases in energy demand and calculates electricity pricing for customers, departing from the traditional monthly billing cycle [80].
-Additionally, with the aid of fog/edge computing, utility providers can efficiently dispatch energy from the grid, enabling consumers to reduce their energy consumption, utilize resources efficiently, and ultimately reduce their electricity costs.
+To address this challenge, the smart grid establishes a two-way communication system between utilities and consumers, which enables more effective management of economic, sustainable, and secure power resources [79].
+Integrating IoT into grid systems enables equipping houses and businesses with smart meters that monitor energy generation, storage, and consumption, and transmit this data to the smart grid.
 Smart Water Systems
@@
 The primary role of smart water systems is to monitor, measure, and efficiently distribute water usage.
-One notable model in this field is the smart water meter designed by Hauber-Davidson and Idris [81].
-These meters can detect water inflow and outflow and identify any potential leaks.
-Smart water meters can also be used in conjunction with data from smart river sensors and weather information to aid in flood prediction [4].
+Hauber-Davidson and Idris have designed a notable model in this field, the smart water metre [81].
+These metres can detect water inflow and outflow and identify any potential leaks.
+Additionally, smart water metres can utilize data from smart river sensors and weather information to assist in flood prediction [4].
+Table 1.6: Relevant works in medical and healthcare applications
+Applications Related Works Applications Related Works
+ECG Monitoring [82, 83] Glucose level monitoring [84, 85]
+Temperature monitoring [86, 87] Blood pressure monitoring [88, 89]
+Oxygen saturation monitoring [90, 91] Mood monitoring [92, 93]
+Medication management [94, 95] Wheelchair management [96, 97]
+Rehabilitation [98, 99] Fitness [100, 101]
+Other notable application
+[102, 103, 104, 105],
+[106, 107, 108, 109]
 1.6.2 Medical and Healthcare
@@
 Furthermore, sharing data collected from these wearable devices with healthcare researchers contributes to the development of safer and more timely healthcare solutions and aids in the discovery of cures and vaccines for emerging diseases.
-This section explores several IoT applications in healthcare and a summary of the associated research papers is presented in Table 1.5.
-Table 1.5: Relevant works in medical and healthcare applications
-Applications Related Works
-ECG Monitoring [82, 83]
-Glucose level monitoring [84, 85]
-Temperature monitoring [86, 87]
-Blood pressure monitoring [88, 89]
-Oxygen saturation monitoring [90, 91]
-Mood monitoring [92, 93]
-Medication management [94, 95]
-Wheelchair management [96, 97]
-Rehabilitation system [98, 99]
-Fitness [100, 101]
-Other notable applications
-[102, 103, 104, 105],
-[106, 107, 108, 109]
-ECG Monitoring
-An Electrocardiogram (ECG) measures the heart’s electrical signals and serves as an indicator of heart health, aiding in the detection of conditions like arrhythmia, prolonged QT interval, and myocardial ischemia.
-IoT integration in ECG monitoring has enabled early detection of heart abnormalities.
-Wu et al.
-[82] have introduced an ECG data monitoring system utilizing a bipotential chip attached to the user’s t-shirt, transmitting data to a smartphone via Bluetooth.
-Real-time ECG data monitoring becomes feasible when IoT systems are combined with big data analytics.
-Bansals and Gandhi [83] have presented a scheme that amalgamated nanoelectronics, IoT, and big data analytics for long-term and continuous ECG data monitoring.
+This section explores several IoT applications in healthcare, and a summary of the associated research papers is presented in Table 1.6.
+Electrocardiogram (ECG) Monitoring
+An ECG measures the heart’s electrical signals and serves as an indicator of heart health, aiding in the detection of conditions like arrhythmia, prolonged QT interval, and myocardial ischemia.
+An interesting example would be Wu et al.’s ECG data monitoring system [82], where a bipotential chip is utilized by attaching to the user’s t-shirt and transmitting data to a smartphone via Bluetooth.
+Combining IoT systems with big data analytics enables real-time ECG data monitoring.
 Glucose Level Monitoring
 Diabetes is a medical condition characterized by higher blood glucose levels than those found in individuals without diabetes.
-Various types of diabetes exist, and doctors use three primary methods to identify them: random plasma glucose tests, fasting plasma glucose tests, and oral glucose tolerance tests.
-Nevertheless, the fingerstick method, involving a small pinprick to the fingertip followed by blood glucose level measurement, remains the most commonly used diagnostic approach.
+Among various approaches to identifying diabetes, the fingerstick method, involving a small pinprick to the fingertip followed by blood glucose level measurement, remains the most commonly used diagnostic approach.
 The advent of IoT technology has brought about improvements in this process, making it quicker and more convenient for patients.
@@
 [84] have introduced an IoT-integrated noninvasive blood glucose monitoring device to continuously monitor glucose levels, eliminating the need for fingerstick testing.
-It is worth noting that optical sensors, such as infrared LED and near-infrared photodiode setups, have also been employed for glucose level measurements.
-In these approaches, the reflected light signal from the human body is used to calculate the glucose level [85].
+It is worth noting that optical sensors, such as infrared LED and near-infrared photodiode setups, have also been used for glucose level measurements.
 Temperature Monitoring
-Monitoring body temperature is crucial for maintaining homeostasis and plays a pivotal role in various diagnostic processes.
 Traditionally, temperature measurement methods involve using thermometers placed in the mouth, ear, or rectum, but these methods often cause discomfort for the patient and pose an elevated risk of infection.
 However, recent advancements in IoT-based temperature monitoring applications have effectively addressed these issues.
-For instance, one innovative approach by [86] where authors have introduced a 3D printed wearable device designed to be inserted into the ear.
-This device utilizes an infrared sensor to measure temperature from the tympanic membrane, ensuring accuracy while remaining unaffected by external factors like environmental conditions or physical activities.
-In another significant development, Gunawan [87] has created an IoT-based temperature monitoring system employing both Arduino and Raspberry Pi.
-This system records temperature data into a database, which users can conveniently access through a web page via their mobile phones or computers.
+For instance, authors in [86] have introduced a 3D-printed wearable device designed to be inserted into the ear.
+This device utilizes an infrared sensor to measure temperature from the tympanic membrane, ensuring accuracy while remaining environment-independent.
 Blood Pressure Monitoring
@@
 Therefore, the integration of IoT into blood pressure monitoring has been a blessing for both doctors and patients.
-In [88], a wearable cuffless gadget has been proposed, capable of measuring both systolic and diastolic pressure, with the results stored in the cloud.
-A cloud and fog-based architecture have been proposed in [89] for long-term, real-time blood pressure monitoring, with the data also being stored for future references.
+A wearable cuffless gadget [88], for example, is capable of measuring both systolic and diastolic pressure, with the results stored in the cloud.
 Oxygen Saturation Monitoring
@@
 The integration of IoT-based technology has led to significant advancements in pulse oximetry, particularly in the healthcare industry.
-In the he study [90], an advanced noninvasive pulse oximetry system has been proposed, capable of measuring oxygen levels, heart rate, and pulse parameters, while transmitting this data to a central server.
-Another study [91], introduced an integrated alert system that sends notifications when oxygen levels reach a critical threshold.
-This system incorporates a pulse oximeter and WLAN router connected through the Blynk server.
+In a study [90], an advanced noninvasive pulse oximetry system has been proposed, capable of measuring oxygen levels, heart rate, and pulse parameters while transmitting this data to a central server.
 Mood Monitoring
-Monitoring a person’s mood is crucial for understanding their mental state and promoting a mentally healthy life.
-It also aids healthcare specialists and psychologists in treating various mental health conditions such as depression, stress, and bipolar disorder.
-Self-monitoring emotional stability empowers individuals to proactively address their mental well-being.
-The integration of IoT into this domain offers numerous advantages.
-Continuous remote monitoring becomes possible, greatly benefiting individuals with severe mental illnesses by enabling real-time responses to emergencies.
-Furthermore, IoT can detect a person’s mental state by analyzing heartbeats through wearable devices.
+The integration of IoT into the mood monitoring domain offers numerous advantages.
+IoT can detect a person’s mental state by analyzing heartbeats through wearable devices.
 Kaur et al.
@@
 The intelligent system, by analyzing emotion variations, determines whether the driver has entered a subconscious state and stops the vehicle’s DC motor accordingly.
-In another study, machine learning and IoT were combined to monitor a user’s mood by analyzing their social networking activities and heart rate through a wearable device [93].
 Medication Management
-Adhering to a prescribed medication schedule is crucial for maintaining one’s health.
-However, this can be particularly challenging for elderly individuals who may experience memory issues, dementia, or cognitive decline.
-Managing daily medication routines becomes increasingly difficult for them.
+Adherence to medication schedules is vital but challenging for elderly individuals with memory issues.
 Fortunately, the integration of IoT offers a solution to this problem, and numerous research efforts have explored using IoT to track patients’ medication compliance.
-In one study [94], a medical box was developed with the primary function of reminding individuals to take their medications.
-Notably, this system features three separate trays for different times of the day (morning, afternoon, and night).
-Additionally, it can measure vital health parameters such as blood pressure, temperature, blood oxygen levels, and ECG readings.
-A mobile application was designed to facilitate two-way communication between patients and doctors, and the recorded information is accessible to both parties.
-Furthermore, authors in [95] have proposed storage conditions, including temperature and humidity solutions, to help users maintain an optimal environment for their medications.
+In [94], a medical box was created to remind individuals to take medications, featuring three trays for different times of the day.
+The system also measures vital health parameters (i.e., blood pressure, temperature, blood oxygen levels, etc.) and facilitates two-way communication between patients and doctors through a mobile application.
 Wheelchair Management
@@
 However, for individuals with brain damage who lack the capability to operate a wheelchair, researchers have been exploring the addition of navigation and tracking systems, with IoT playing a crucial role in these wheelchair advancements.
-In one study [96], an IoT-based steering and real-time obstacle detection system for wheelchairs has been proposed.
-This system can identify obstacles through image processing methods, enhancing the wheelchair’s safety and usability.
-An advanced and automated smart wheelchair was reported in another study [97].
-This wheelchair not only monitored movement but also offered features such as an umbrella, foot mat, head mat, and obstacle detection.
-These innovations significantly improved interaction with the living environment and enhanced the user’s overall experience.
+An advanced and automated smart wheelchair was reported in a study [97], which not only monitors movement but also offers features such as an umbrella, foot mat, head mat, and obstacle detection.
+These innovations have significantly improved interaction with the living environment and enhanced the user’s overall experience.
 Rehabilitation System
-Rehabilitation is a critical aspect of restoring an individual’s functional abilities, aiming to help those with diseases or disabilities regain their normal lifestyle.
 The application of IoT in this field is versatile and has proven effective in various areas, including cancer treatment, sports injury recovery, stroke rehabilitation, and addressing physical disabilities.
 For instance, an innovative smart walker was introduced in a particular study [98].
-This walker utilizes a multi-modal sensor to monitor the patient’s walking pattern, and the collected data can be accessed by doctors and caregivers through a mobile application, facilitating better monitoring and support.
-A sports rehabilitation system was reported which monitors various parameters such as temperature, motion posture, electromyography (EMG), electrocardiography (ECG), among others, and provides real-time feedback to athletes [99].
-These examples highlight the potential of IoT integration in modernizing rehabilitation systems, making them more efficient and effective.
+Doctors and caregivers can access the collected data through a mobile application, facilitating better monitoring and support, as this walker utilizes a multi-modal sensor to monitor the patient’s walking pattern.
 Fitness
-The quality of an individual’s life is significantly influenced by their level of fitness and daily physical activity.
-To facilitate fitness monitoring and promote healthier lifestyles, various applications leveraging the IoT have been developed.
-One approach involves utilizing smartphone accelerometer data to assess users’ activity levels.
-This data can reveal metrics such as the number of steps taken, the duration of physical activity, and periods of inactivity.
-Wearable fitness trackers are readily available in the market and have gained popularity as convenient devices for monitoring fitness levels.
-Additionally, IoT sensors can be integrated with gym equipment to provide real-time monitoring of individuals’ fitness progress.
-For instance, in a notable study [100], a “smart mat” have been introduced.
-This mat is designed to measure the number of exercise steps performed on its surface, offering valuable insights into users’ workout routines.
-In the field of sports science, fitness assessment and training load monitoring have proven to be invaluable tools for both coaches and athletes.
-IoT technology enables the continuous and non-invasive monitoring of physiological biomarkers, such as saliva or sweat.
-This data can be leveraged to optimize hydration strategies, ultimately enhancing athletes’ overall performance [101].
-These IoT-driven fitness monitoring applications and devices contribute to a more proactive approach to health and fitness, promoting active lifestyles and well-being.
+Regular physical activity and maintaining a high level of fitness significantly influence the quality of an individual’s life.
+Developers have created various applications leveraging the IoT to facilitate fitness monitoring and promote healthier lifestyles.
+These approaches include assessing users’ activity levels and revealing metrics, including the duration of physical activity and periods of inactivity, by utilizing smartphone accelerometer data.
+Nowadays, wearable fitness trackers are readily available in the market and have gained popularity as convenient devices for monitoring fitness levels, for instance, smart mats to provide insights into users’ workout routines [100] and fitness assessment and training load monitoring to optimize athletes’ hydration strategies [101].
 Other Notable Applications
@@
 There are numerous domains where IoT has already been implemented and where its potential benefits are being realized, leading to a significant increase in the adoption of healthcare IoT (HIoT) technology.
-For instance, in cancer treatment, IoT-based methods have emerged as powerful tools.
-An innovative IoT-based cancer treatment approach is introduced in a recent study, that encompasses various stages, including chemotherapy and radiotherapy [102].
+For cancer treatment, IoT-based methods, for example, have emerged as powerful tools.
+An innovative IoT-based cancer treatment approach is introduced in a recent study that encompasses various stages, including chemotherapy and radiotherapy [102].
 Additionally, this system securely stores lab test results on a cloud server, allowing physicians to monitor medication dosages and enabling remote consultations through a dedicated mobile application.
-Furthermore, HIoT has found applications in detecting skin lesions [103], with notable advancements in lung cancer detection through state-of-the-art machine learning algorithms and IoT-based systems [104, 105, 106].
-In the realm of surgical training and medical procedures, IoT has been harnessed to create next-generation solutions.
+Furthermore, HIoT has found applications in detecting skin lesions [103], with notable advancements in lung cancer detection through state-of-the-art ML algorithms and IoT-based systems [104, 105, 106].
+IoT has revolutionized the realm of surgical training and medical procedures by creating next-generation solutions.
 One such development involves a surgical training framework that employs virtual reality to simulate realistic training environments.
 This framework also enables interaction with surgeons from around the world, fostering collaborative learning and expertise sharing [107].
-Additionally, a human-robot collaborative system has been devised for performing minimally invasive surgeries [108].
-Monitoring hemoglobin levels in the blood has become more accessible through portable devices equipped with photoplethysmography (PPG) sensors, light-emitting diodes (LEDs), and photodiodes.
-These devices enable the non-invasive measurement of hemoglobin levels, enhancing healthcare monitoring and diagnosis [109].
-Asthma Monitoring is another notable sector where sensor-based asthma monitoring system is used to track the user’s respiratory rate [110].
+Monitoring haemoglobin levels in the blood has become more accessible through portable devices equipped with photoplethysmography sensors, light-emitting diodes (LEDs), and photodiodes.
+These devices enable the non-invasive measurement of haemoglobin levels, enhancing healthcare monitoring and diagnosis [109].
 Numerous other HIoT applications are currently in use or under research, underscoring the ongoing revolutionary impact of IoT in the field of healthcare.
-It is expected that IoT will continue to drive advancements and improvements in healthcare delivery and patient outcomes.
+It is expected that the IoT will continue to drive advancements and improvements in healthcare delivery and patient outcomes.
 1.6.3 Smart Agriculture and Environment
 Agriculture holds a crucial position in a country’s economic progress.
-However, crop yields can be significantly impacted by various factors, including soil moisture and environmental variables like carbon dioxide levels, temperature, and humidity.
+Various factors, including soil moisture and environmental variables like carbon dioxide levels, temperature, and humidity, can significantly impact crop yields.
 To enhance agricultural outcomes, it becomes essential to implement robust surveillance systems in the fields.
-Achieving this efficiently is made possible through the integration of the IoT.
-The following section has explored several applications in this agricultural domain as well as a summary of the related research work in Table 1.6.
+The integration of the IoT enables efficient achievement of this goal.
+The following section has explored several applications in this smart agricultural domain as well as a summary of the related research work in Table 1.7.
 Table 1.7: Relevant studies in smart agriculture and environment domain
-Applications Related Works
-Water-saving irrigation [111, 112]
-Crop growth environment monitoring [113, 114]
-Animal and plant life information monitoring [115, 116]
-Intelligent agricultural machinery [117, 118]
-Agricultural product quality safety and traceability [119, 120]
-Diseases Monitoring [121, 122]
+Applications Related Works Applications Related Works
+Water-saving irrigation [111, 112] Diseases Monitoring [113, 114]
+Animal and plant life
+information monitoring
+[115, 116]
+Intelligent agricultural
+machinery
+[117, 118]
+Agricultural product quality
+safety and traceability
+[119, 120]
+Crop growth environment
+monitoring
+[121, 122]
 Water Saving Irrigation
-Water is a crucial factor in crop growth, yet its scarcity is an escalating concern.
-Different crops have varying water requirements, necessitating a dynamic irrigation approach.
-The integration of IoT technology in agriculture addresses this challenge by revolutionizing traditional flood irrigation methods, ultimately resolving water shortage problems.
-To address these issues, Yang et al.
-[111] have proposed an intelligent system leveraging neural networks for water-efficient irrigation.
+Water scarcity in agriculture is a growing concern, necessitating a dynamic irrigation approach due to varying crop water requirements.
+IoT integration revolutionizes traditional flood irrigation, offering a solution to water shortage problems in crop growth.
+Yang et al.
+[111] have proposed an wireless sensor network based system leveraging neural networks for water-efficient irrigation.
 This method enhances irrigation efficiency by minimizing human intervention and reducing wastage due to excessive drainage.
-Li et al.
-[112] have developed a series of GPRS DTU-based intelligent agricultural irrigation control systems capable of tailoring irrigation to crop growth needs.
 Crop Growth Environment Monitoring
@@
 IoT devices integrated into agricultural systems can sense and analyze these environmental factors, enabling remote field monitoring and the creation of an optimal farming environment tailored to these variables.
-Chen et al.
-[113] have proposed an environmental monitoring solution that utilizes WSNs for millet farmland.
-In contrast, Lin et al.
+Lin et al.
 [114] have designed a wireless environmental monitoring system that harnesses soil energy to enable cost-effective remote monitoring of farmland environments.
 Animal and Plant Life Information Monitoring
-Efficient agricultural production necessitates the monitoring of both animal and plant information, a pivotal phase in the process.
-This monitoring plays a substantial role in advancing improved production, increasing profitability, and ensuring high-quality product development.
+Effective agricultural production requires comprehensive monitoring of both plant and animal information, which is crucial for enhancing production, increasing profitability, and ensuring high-quality product development.
   1. 1.
-    Animal Life Information Monitoring
-    Monitoring diverse aspects of animal behavior, including their food consumption, body temperature, activity levels, and health status, enables the tracking of their physiological and nutritional well-being, ensuring their healthy development.
-Similar to other monitoring systems, animal information monitoring includes various types of sensors, actuators, wireless connections, and terminal equipment.
+    Animal Life Information Monitoring: Monitoring diverse aspects of animal behavior, including their food consumption, body temperature, activity levels, and health status, enables the tracking of their physiological and nutritional well-being, ensuring their healthy development.
 In [115], the authors have proposed an infrared-based body temperature measurement system for pigs, allowing the early detection of diseases.
   2. 2.
-    Plant Life Information Monitoring
-    Plant life information can be assessed through two distinct methods: external information and internal information.
-External data involves analyzing visual factors such as diseases, pests, and leaf color, while internal information includes parameters like chlorophyll content and photosynthetic rate.
-Various wireless sensor devices can be connected to plants to facilitate remote and continuous monitoring of their information, ultimately allowing for the early detection of diseases and promoting healthy plant growth.
+    Plant Life Information Monitoring: Wireless sensor devices, when connected to plants, enable remote and continuous monitoring of both external factors (such as diseases, pests, and leaf color) and internal factors (including chlorophyll content and photosynthetic rate).
+This technology allows for early disease detection and promotes overall healthy plant growth.
 Porto et al.
-[116] have introduced a system for citrus traceability that examines the environmental conditions suitable for the optimal growth of citrus.
-This system is capable of identifying and preventing plant diseases, ensuring the robust health of citrus crops.
+[116] have introduced a citrus traceability system that assesses environmental conditions for optimal growth, identifying and preventing plant diseases to ensure robust citrus crop health.
 Intelligent Agricultural Machinery
-Intelligent machinery demonstrates the capacity to autonomously manage a wide range of agricultural operations, encompassing tasks such as cultivation, sowing, transplanting, fertilization, pesticide application, feeding, irrigation, picking, and harvesting, all executed with precision and efficiency.
-Moreover, it has the capability to gather a variety of data about the farm, including soil moisture and water quality, as well as ambient information like temperature and humidity.
-These data resources can be effectively harnessed to implement precision agriculture and enhance breeding practices [123].
-IoT technology plays a pivotal role in reducing manual labor in agricultural machinery operations.
-It enables remote monitoring and standardizes machinery functions through various technologies such as sensors and wireless communication methods.
+Intelligent machinery autonomously manages a wide range of agricultural operations, such as cultivation, sowing, transplanting, fertilization, pesticide application, feeding, irrigation, picking, and harvesting, all executed with precision and efficiency.
+Moreover, it has the capability to gather a variety of data about the farm, including soil moisture and water quality, as well as ambient information like temperature and humidity, which can be effectively harnessed to implement precision agriculture and enhance breeding practices [123].
+IoT technology plays a key role in minimizing manual labor in agriculture by enabling remote monitoring and standardizing machinery functions through sensors and wireless communication [123].
 Sowjanya et al.
 have introduced a versatile autonomous robot vehicle in [117] where the vehicle is equipped with Bluetooth technology for remote control and is capable of independently executing a range of tasks including farming, seeding, and irrigation.
-In a similar vein, researchers in [118] have presented a robot system designed for fruit detection.
-This system employs a single-lens multi-box detection method and a stereo camera to identify the position of fruits in three dimensions, enabling efficient fruit picking.
 Agricultural Product Quality Safety and Traceability
-IoT plays a pivotal role in enhancing agricultural product quality, safety, and traceability, particularly in the domains of product warehousing, logistics, and distribution.
-It enables automatic product identification, tracking, and the accurate tallying of incoming and outgoing items within warehouses.
-By leveraging WSNs, real-time remote monitoring of vehicles or containers used for product delivery can be achieved, ensuring the traceability of agricultural product origins and destinations.
-Numerous countries have recognized the imperative for real-time traceability and have developed intelligent systems to address this need.
-These systems include the American agricultural product traceability system, the European beef traceability system, the Swedish agricultural traceability system, the Japanese food traceability system, and the Australian livestock traceability system.
-In the pursuit of agricultural product safety and traceability, Jiang et al.
-[119] have established a comprehensive agricultural product safety traceability platform.
-This platform enables the real-time automatic collection, processing, and display of agricultural product data, thereby enhancing product traceability and reducing the costs associated with tracking and monitoring agricultural items.
-Gu et al.
-have introduced a system in [120] designed to trace the complete lifecycle of pigs, covering stages such as slaughter, processing, and sales, thereby ensuring transparency and traceability throughout the pork production process.
+IoT significantly improves agricultural product quality, safety, and traceability, particularly in warehousing, logistics, and distribution, by enabling automatic identification, tracking, and accurate tallying of products.
+Various countries have implemented real-time traceability systems, such as the American, European, Swedish, Japanese, and Australian systems, recognizing the crucial need for effective tracking in agriculture.
+Jiang et al.
+[119] developed a comprehensive agricultural product safety traceability platform, facilitating real-time automatic data collection, processing, and display to enhance traceability and reduce associated tracking costs.
 Diseases Monitoring
 The integration of IoT for real-time and continuous monitoring offers farmers and relevant authorities the capability to identify diseases at an early stage and implement preventive measures before they escalate.
-IoT, when combined with machine learning, artificial intelligence, or image processing, has yielded various systems for early disease detection and prevention in agriculture.
-The farm’s environment plays a pivotal role in disease occurrence, and in [121], authors have presented a framework that combines various sensor devices through wireless sensor networks for monitoring various environmental factors.
-This system collects real-time data and employs multiple machine learning models like KNN, Random Forest, Logistic Regression, and Linear Regression to analyze insights from this data.
-Additionally, a framework for environmental observation-based crop disease monitoring, which continuously monitors plant growth and diseases and notifies farmers when necessary, has been proposed in [122].
+The farm’s environment plays a pivotal role in disease occurrence.
+For example, the framework introduced in [121] integrates various sensor devices through wireless sensor networks for monitoring various environmental factors.
 1.6.4 Smart Home (SH)
-The adoption of smart home technology has seen a significant and continuous increase in recent years.
-As technologies such as sensors, actuators, and wireless networks have matured, people’s trust in these advancements has grown, leading to a greater reliance on technology to enhance their quality of life.
-In a smart home, various types of sensors are strategically deployed, each with its specific function.
+In a smart home, various types of sensors are strategically deployed, each with its own specific function.
 Smart homes simplify daily tasks for users, proving especially beneficial for those prone to forgetting routine actions like locking doors or turning off appliances.
@@
 Smart homes contribute to improved energy efficiency by automatically turning off devices not in use and notifying users of any unusual incidents.
-MavHome, short for “Managing An Intelligent Versatile Home” [5], represents a smart home system that functions as an intelligent agent.
-It employs prediction algorithms to perform various tasks in response to user-initiated events.
-As for energy conservation, an intelligent home achieves energy conservation through the utilization of sensors and the context-aware capabilities of IoT.
-Data gathered by these sensors are transmitted to a context aggregator, which then forwards the data to a context-aware service engine.
+MavHome [5], for example, employs prediction algorithms to perform various tasks in response to user-initiated events.
+As for energy conservation, an intelligent home achieves it through the utilization of sensors and the context-aware capabilities of IoT.
+Data gathered by these sensors is transmitted to a context aggregator, which then forwards the data to a context-aware service engine.
 This engine analyzes the data and determines appropriate actions.
@@
 1.6.5 Smart Manufacturing System (SMS)
-With the development and evolution of IoT (Internet of Things), Industrial IoT, Artificial Intelligence, and Cyber-Physical Systems, many countries have opted to transform their manufacturing systems into smart manufacturing systems.
+With the development and evolution of IoT, industrial IoT, artificial intelligence, and cyber-physical systems, many countries have opted to transform their manufacturing systems into smart manufacturing systems.
 Through the integration of smart technologies, these systems facilitate a rapid and extensive flow of data within and among manufacturing processes.
-Equipped with these data and employing advanced information and communication technology, smart manufacturing systems possess the capability to swiftly respond to global demands, efficiently utilise materials, energy, and labour resources, and deliver customised products on time [125].
+Equipped with this data and employing advanced information and communication technology, smart manufacturing systems possess the capability to swiftly respond to global demands, efficiently utilise materials, energy, and labour resources, and deliver customised products on time [125].
 What sets the smart manufacturing model apart from other manufacturing paradigms is its vision of the next generation of manufacturing with enhanced capabilities [126].
-Smart manufacturing systems adapt to new circumstances by leveraging real-time information for intelligent decision-making and by proactively predicting and preventing potential failures.
+These systems adapt to new circumstances by leveraging real-time information for intelligent decision-making and by proactively predicting and preventing potential failures.
 1.6.6 Internet of Robotics Things (IoRT)
-The Internet of Robotic Things (IoRT) is a concept that combines the principles of IoT (Internet of Things) and Robotics.
+The Internet of Robotic Things is a concept that combines the principles of IoT and robotics.
 IoRT represents an emerging technology that incorporates robots within an IoT ecosystem as objects, enabling communication, collaboration, and automation.
-These robots are seamlessly integrated into smart environments, where they perform a wide range of tasks.
+These robots seamlessly integrate into smart environments, performing a wide range of tasks.
 These tasks span from personal activities within smart homes to applications in the healthcare industry.
@@
 1.6.7 Oil and Gas Industry
-The integration of IoT into the oil and gas industry can enable automated and optimized execution of industry operations, leading to more efficient decision-making.
-IoT-based applications in the oil and gas sector offer numerous advantages, including enhanced operational efficiency through automation and optimization, resulting in overall cost reduction.
-Environmental monitoring sensors deployed in this context can detect and alert to the presence of poisonous or harmful substances in the environment.
-Moreover, IoT technology provides real-time data, significantly improving decision-making processes.
-Importantly, IoT integration enables the early detection of potential hazards, empowering authorities to take prompt and effective preventive measures.
-The oil and gas industry entails substantial expenditures and significant health and safety risks, highlighting the importance of continuous and real-time monitoring within this sector.
-To achieve this objective, the oil and gas industry is increasingly adopting IoT-based remote monitoring solutions.
+The oil and gas industry, facing substantial costs and safety risks, embraces IoT-based remote monitoring for real-time field equipment oversight and data-driven decision-making.
 These solutions allow for the remote monitoring of field equipment, the analysis of field data, collaborative data-driven decision-making, and the implementation of control commands to optimize asset performance while mitigating health, safety, and environmental (HSE) hazards [127].
-Furthermore, IoT integration in this industry aims to reduce human labor, minimize time wastage, and enhance accuracy through automation.
-An illustrative example is Equinor’s deployment of a well optimization system in the Bakken oil field.
-In this case, Equinor have utilized IoT devices and machine learning algorithms in approximately 50 wells.
-This implementation led to optimized well operation and maintenance, resulting in a remarkable 33% increase in oil production from these.
+Furthermore, IoT integration in the oil industry focuses on reducing human labor, minimizing time wastage, and improving accuracy through automation, as exemplified by Equinor’s well optimization system in the Bakken oil field [freeman2018enabling].
+By deploying IoT devices and ML algorithms in around 50 wells, Equinor achieved a 33% increase in oil production through optimized well operation and maintenance.
 1.6.8 Smart Retail
@@
 This shift allows the entire retail sector to migrate from offline to online, enabling customers to independently conduct their shopping through self-service while facilitating smooth interactions between retailers and their customers.
-Furthermore, IoT technologies like RFID can be employed to monitor products, and sensors can be deployed to collect customer data, which can then be utilized for analyzing customer buying behavior and enhancing business profitability.
+Furthermore, retailers can employ IoT technologies like RFID to monitor products and deploy sensors to collect customer data, which they can then utilize for analyzing customer buying behavior and enhancing business profitability [129, 130].
 Additionally, customers have the option to make payments through online transactions and monitor their orders using online services [128].
-In the study presented in [129], the authors have introduced a smart unstaffed retail store that relies on AI and IoT for commodity detection.
-Furthermore, in [130], two models, namely the Retail 4.0 IoT Consumer Model and the Retail 4.0 IoT Retailer Model, have been put forward.
-These models are capable of analyzing consumer behavior and discerning consumer buying patterns with ease.
 1.6.9 Industrial Internet of Things (IIoT)
-The IIoT, often recognized as The Industrial Internet of Things (IIoT), holds significant potential according to numerous market researchers.
-It serves as an extension of IoT specifically customized for the industrial sector and its applications.
-IIoT empowers industries and enterprises to enhance and optimize their operations by leveraging machine-to-machine communication, big data analytics, and machine learning.
+The IIoT holds significant potential, according to numerous market researchers, by serving as an extension of the IoT specifically customized for the industrial sector and its applications.
+It empowers industries and enterprises to enhance and optimize their operations by leveraging M2M communication, big data analytics, and ML.
 The scope of IIoT is extensive, encompassing a wide array of connected industrial devices and systems.
-These include, but are not limited to, connected electric meters, wastewater systems, flow gauges, pipeline monitors, manufacturing robots, and various other types of industrial equipment and devices [131].
+Connected electric meters, wastewater systems, flow gauges, pipeline monitors, manufacturing robots, and various other types of industrial equipment and devices are included in this list [131].
 One notable application of IIoT is in the mining industry, where companies like CISCO have implemented IIoT solutions to improve safety and efficiency in underground mines.
 These solutions involve connecting people, tracking the locations of miners and vehicles, monitoring vehicle statuses, and automating building controls.
-Additionally, the mine’s blasting system is integrated with the personnel location tracking system to enhance miner safety [132].
-Moreover, IoT has the potential to revolutionize the manufacturing field by enhancing productivity, maintenance practices, machine monitoring, and ensuring the safety of workers [133].
 1.6.10 Social Life and Entertainment
-In addition to work or professional activities, social life and entertainment are integral parts of a person’s life, and several applications have been developed to monitor and enhance human social activities.
+Several applications have been developed to monitor and enhance human social activities, in addition to work or professional activities, as social life and entertainment are integral parts of a person’s life.
 Portable devices such as mobile phones and tablets possess sensing capabilities and communication technologies that facilitate interactions between individuals.
 Integrating IoT into an individual’s social life can contribute to emotion detection, community building, and emotional support.
-For instance, CircleSense [134] is an application designed to identify a person’s social circle by analyzing their social activities using various sensors.
+CircleSense [134] is an application that analyzes a person’s social activities using various sensors to identify their social circle.
 It also tracks the person’s location via location sensors and employs Bluetooth technology to identify people in proximity.
-Affective computing is a technology that analyzes various aspects of a person’s behavior, including facial expressions, speech, body gestures, hand movements, and sleep patterns, to identify and respond to their emotions.
-An application of affective computing is Camy [135].
-Camy is an artificial pet dog designed to interact with human beings, displaying affection and empathy.
-It assesses the owner’s mood and provides suggestions, emotional support, and encourages playful behavior accordingly.
+Camy, an artificial pet dog, expresses affection and empathy through the use of effective computing technology.
+This technology analyzes multiple aspects of a person’s behavior, such as facial expressions, speech, body gestures, hand movements, and sleep patterns, to identify and appropriately respond to their emotions [135].
 A Table 1.8 has been provided for presenting information about relevant works in the aforementioned domains.
 Table 1.8: Relevant works in other applications
-Applications Related Works
-Smart home (SH) [5, 136]
-Smart manufacturing system (SMS) [125, 126]
-Internet of robotics things (IoRT) [137, 138]
-Oil and gas [127]
-Smart Retail [129, 130]
-Industrial IoT [132, 133]
-Social Life and Entertainment [134, 135]
+Applications Related Works Applications Related Works
+Social Life and
+Entertainment
+[134, 135]
+Smart manufacturing
+system
+[125, 126]
+Internet of robotics things [137, 138] Oil and gas [127]
+Smart Retail [129, 130] Industrial IoT [132, 133]
+Smart home [5, 136]
 1.7 Challenges and Future Directions
@@
 Nonetheless, IoT also grapples with numerous issues and challenges that pose hurdles to its seamless implementation and expansion.
-This section offers an in-depth exploration of the key challenges confronting IoT.
+This section offers an in-depth exploration of the key challenges confronting the IoT.
 1.7.1 Broad and Open Research Challenges
-Broad and open research challenges refers to complex and critical challenges in the IoT system that do no have any straightforward or predefined solutions.
-These challenges require extensive research, exploration in order to propose efficient solutions.
+Broad and open research challenges refer to complex and critical challenges in the IoT system that do not have any straightforward or predefined solutions.
+These challenges require extensive research and exploration in order to propose efficient solutions.
 They are “broad” in the sense that they encompass a wide range of related issues and considerations, and they are “open” because they may not have clear-cut or definitive answers, leaving room for ongoing research, experimentation, and discovery.
@@
 Building Intelligent Environments based on IoT Paradigm
-Creating a smart environment requires a vast number of devices, sensors, and complementary technologies to facilitate their inter-connectivity.
+Creating a smart environment requires a vast number of devices, sensors, and complementary technologies to facilitate their interconnectivity.
 Managing this large volume of objects constitutes the initial challenge in the realm of IoT and intelligent environments.
@@
 One approach is to employ decentralized systems instead of centralized ones, reducing the volume of data sent to the cloud for processing.
-Techniques like data filtering, compression, and load balancing can further minimize data size.
+Techniques like data filtering, compression, and load balancing can further minimize the size of the data.
 Utilizing IoT technologies with robust device management and maintenance capabilities is also beneficial.
@@
 The heterogeneity of IoT devices and the diversity of various IoT applications result in various security and privacy issues.
-People are primarily concerned about potential privacy invasion and security threats when using these technological devices.
-Further details are provided below.
+People are primarily concerned about potential privacy invasions and security threats when using these technological devices.
+More information is provided below.
   1. 1.
-    Security
-    Different layers of IoT are vulnerable to various kinds of attacks based on the technologies and protocols used in these respective layers.
+    Security: Different layers of the IoT are vulnerable to various kinds of attacks based on the technologies and protocols used in these layers.
 According to [140], IoT layers including perception layer, network layer, and application layer face various security attacks.
-As the major role of perception layer is to collect data, so the security challenges in this layer focuses in Falsifying the data and destroying perception devices.
-Attacks including Node Capture Attacks, Malicious code Injection Attacks, False Data Injection Attacks, Replay Attacks, Cryptanalysis Attacks & Side Channel Attacks, Eavesdropping and Interference and Sleep Deprivation Attacks are faced by this layer [140].
+As the major role of the perception layer is to collect data, the security challenges in this layer focus on falsifying the data and destroying perception devices.
+Attacks including Node Capture Attacks, Malicious code Injection Attacks, False Data Injection Attacks, Replay Attacks, Cryptanalysis Attacks and Side Channel Attacks, Eavesdropping and Interference, and Sleep Deprivation Attacks are faced by this layer [140].
     Since the basic function of network layer is to transmit collected data, particularly using wireless technologies, security challenges in this layer revolve around the availability of network resources and the wireless network.
 Challenges at the Network Layer encompass Denial-of-Service (DoS) Attacks, Spoofing Attacks, Sinkhole Attacks, Wormhole Attacks, Man-in-the-Middle Attacks, Routing Information Attacks, Sybil Attacks, and Unauthorized Access [140].
-Application layer, which focuses on providing user-requested services, challenges primarily revolve around software attacks, including: Phishing Attack, Malicious Virus/Worm, Malicious Scripts [140].
+Application layer, which focuses on providing user-requested services, challenges primarily revolve around software attacks, including phishing attacks, malicious viruses/worms, malicious scripts [140].
     It is imperative to propose secure, robust, and reliable authentication schemes to detect and defend unauthorized access [141].
-To defend against worms, viruses, and malicious scripts bridging firewalls, virus detection techniques and script detection techniques such as honeypot techniques, static code analysis, and dynamic action detection must be implemented.
-Along with that, secured routing protocols must be introduced to ensure secure routing.
+Virus detection techniques and script detection techniques such as honeypot techniques, static code analysis, and dynamic action detection must be implemented to defend against worms, viruses, and malicious scripts bridging firewalls.
+Furthermore, introducing secured routing protocols is essential to ensure secure routing.
   2. 2.
-    Privacy
-    As previously mentioned, IoT devices continuously generate vast amounts of real-time data, which undergo three main stages: (1) Data Collection, (2) Data Aggregation , and (3) Data Mining and Analytics.
-While these processes enhance our lives by providing various services, they also raise concerns about data privacy in IoT.
+    Privacy: IoT devices continuously generate vast amounts of real-time data, which undergoes three main stages: (1) Data Collection, (2) Data Aggregation , and (3) Data Mining and Analytics.
+While these processes enhance our lives by providing various services, they also raise concerns about data privacy in the IoT.
 Privacy breaches in IoT can have serious repercussions for both the IoT network and its users, including financial losses, property damage, and even risks to human safety and security [140, 142].
-    For instance, consider the smart grid, Adversaries can readily seize control of the smart meters, allowing them to access or manipulate the collected data.
+    For instance, consider the smart grid, where adversaries can readily seize control of the smart metres, allowing them to access or manipulate the collected data.
 This could potentially compromise the confidentiality and privacy of energy consumption data.
-By using these altered data, utility providers may make inaccurate assessments of energy supply and demand within the grid, resulting in erroneous energy dispatch decisions.
+By using this altered data, utility providers may make inaccurate assessments of energy supply and demand within the grid, resulting in erroneous energy dispatch decisions.
 This, in turn, could lead to imbalances in energy supply and demand, potentially causing widespread power outages.
@@
 Therefore, it is imperative to deploy privacy preservation schemes to prevent data leakage and ensure that private data remains inaccessible to adversaries [140, 143].
-    Privacy-preserving mechanisms in the context of IoT data processing can be categorized into three main groups: (i) privacy preservation during data collection, (ii) privacy preservation during data aggregation, and (iii) privacy preservation during data mining and analytics.
-While various techniques such as encryption and key management can be applied to protect privacy in data collection, mining, and analytics, the majority of efforts in IoT privacy preservation have focused on data aggregation.
+    There are three main groups for categorizing privacy-preserving mechanisms in the context of IoT data processing: (i) privacy preservation during data collection, (ii) privacy preservation during data aggregation, and (iii) privacy preservation during data mining and analytics.
+While various techniques, such as encryption and key management, can be applied to protect privacy in data collection, mining, and analytics, the majority of efforts in IoT privacy preservation have focused on data aggregation.
 Data aggregation involves processing relevant data in multiple locations, making it challenging to ensure privacy using traditional encryption methods.
-Therefore, several privacy-preserving mechanisms have been developed specifically for data aggregation, falling into the following categories:
+As a result, researchers have developed several privacy-preserving mechanisms specifically for data aggregation, which can be categorized as follows:
     1. (a)
@@
     2. (b)
-      Encryption-based privacy preservation, which utilizes encryption techniques like homomorphic encryption, commitment mechanisms, secret sharing, and zero-knowledge proofs to prevent data from being eavesdropped by adversaries during aggregation [144, 145].
+      Encryption-based privacy preservation, which prevents adversaries from eavesdropping on data during aggregation by utilizing encryption techniques such as homomorphic encryption, commitment mechanisms, secret sharing, and zero-knowledge proofs [144, 145].
     3. (c)
-      Perturbation-based privacy preservation, where techniques such as data customization, data sharing, and random noise injection are used to perturb raw data, ensuring privacy during aggregation.
+      Perturbation-based privacy preservation, where techniques such as data customization, data sharing, and random noise injection perturb raw data to ensure privacy during aggregation.
     Among these, perturbation-based privacy-preserving schemes are popular in IoT due to their direct operation on raw data.
-However, many of these perturbation-based privacy preserving schemes sacrifice data utility to achieve privacy.
+However, many of these perturbation-based privacy-preserving schemes sacrifice data utility to achieve privacy.
 This reduction in data utility can hinder the support of services requested by IoT applications.
-Therefore, a significant challenge, in the field of data privacy preservation in IoT, is designing schemes that strike a balance between privacy and data utility, making it a crucial area for future research [21, 140].
-In summary, safeguarding data privacy in IoT is crucial to prevent these adverse outcomes and maintain the security and integrity of both individuals and the IoT ecosystem.
+Therefore, a significant challenge in the field of data privacy preservation in the IoT is designing schemes that strike a balance between privacy and data utility, making it a crucial area for future research [21, 140].
+In summary, safeguarding data privacy in the IoT is crucial to preventing these adverse outcomes and maintaining the security and integrity of both individuals and the IoT ecosystem.
 Compatibility
-In an IoT network, devices from various vendors are interconnected, which can lead to monitoring and management challenges.
+Interconnecting devices from various vendors in an IoT network can pose monitoring and management challenges.
 Different industries currently rely on a multitude of standards to support their applications.
@@
 Scalability
-IoT is an ever-expanding network, and heterogeneous devices are expected to continuously join the IoT network in the future.
-As a result, as the number of devices increases, ensuring smooth connectivity, effective data management, and overall system performance at a small scale becomes increasingly challenging.
+In the future, heterogeneous devices are expected to continuously join the ever-expanding IoT network.
+As a result, as the number of devices increases, ensuring smooth connectivity, effective data management, and overall system performance on a small scale becomes increasingly challenging.
 Therefore, the scalability of IoT poses an ongoing challenge for the future of this technology.
@@
 Energy Efficiency
-Small smart devices that make up IoT systems often have limited battery power, which cannot be easily replaced.
+Small smart devices that comprise IoT systems often have limited battery power, which is not easily replaceable.
 This limitation can lead to a global energy crisis and high power consumption, as well as constraints on memory and processing capabilities.
@@
 While some routing protocols do support low-power communication, they are still in the early stages of development, and the constrained energy of smart devices may not be sufficient to fully utilize these WSN routing protocols.
-To tackle these challenges effectively, it’s essential to emphasize the creation of low-power hardware and the adoption of energy-efficient protocols like MQTT-SN or CoAP, rather than relying on more power-intensive alternatives such as HTTP.
+To tackle these challenges effectively, it’s essential to emphasize the creation of low-power hardware and the adoption of energy-efficient protocols like MQTT-SN or CoAP rather than relying on more power-intensive alternatives such as HTTP.
 Additionally, harnessing over-the-air (OTA) firmware updates can ensure devices remain optimized and bug-free, thus diminishing the necessity for physical maintenance visits.
@@
 Mobility Management
-Mobility management in IoT refers to the ability to handle devices that move within the network seamlessly.
+Mobility management in the IoT refers to the ability to handle devices that move within the network seamlessly.
 It is a crucial aspect because many IoT devices are not stationary and need to communicate as they change locations.
 The presence of mobile devices in IoT setups can lead to challenges in how routing protocols and IoT networks work efficiently.
-The current methods used for devices that move, like in sensor networks, mobile adhoc networks, and vehicular networks, can’t effectively handle the various problems related to routing because of these sensors having limited processing power and energy resources.
+The current methods used for devices that move, like in sensor networks, mobile adhoc networks, and vehicular networks, can’t effectively handle the various problems related to routing because these sensors have limited processing power and energy resources.
 To address these challenges, IoT systems employ various mobility management techniques and protocols, aiming to provide reliable and seamless communication for mobile devices in the IoT ecosystem.
@@
 Processing, Analysis and Management of Data
-The procedure for Processing, Analysis and Management of Data is tremendously challenging because of the heterogeneous nature of IoT devices and large sale of data generation.
-Currently, most systems utilize the centralized cloud based system for performing computationally intensive tasks and delivery of data.
+The procedure for processing, analysing, and managing data is tremendously challenging because of the heterogeneous nature of IoT devices and the large scale of data generation.
+Currently, most systems utilize the centralized cloud-based systems for performing computationally intensive tasks and delivering data.
 However, an ongoing concern revolves around the limitations of traditional cloud architectures when it comes to efficiently handling the vast amounts of data generated and utilized by IoT-enabled devices.
@@
 To address this challenge, most systems are currently relying on existing solutions like mobile cloud computing and fog computing, both of which utilize edge processing [146].
-Figure 1.6: Ethical Issues in IoT
 Other Challenges
@@
 The IoT network introduces ambiguity, making it challenging to distinguish between physical and virtual devices and even humans due to the ease of transformation between these categories.
-Quality and traffic control have become more complex due to the miniaturization and the huge number of IoT devices.
+Quality and traffic control have become more complex due to the miniaturization and huge number of IoT devices.
 Managing unique identifications for each IoT device is also a growing concern.
-Furthermore, IoT goes beyond geographical boundaries, with applications like healthcare offering services internationally.
-This global reach poses challenges for nations, as data generated within their borders can be collected and transmitted to service providers located anywhere in the world, raising issues related to data privacy and jurisdiction.
+Furthermore, the IoT goes beyond geographical boundaries, with applications like healthcare offering services internationally.
+Nations face challenges due to the global reach of IoT, as data generated within their borders can be collected and transmitted to service providers located anywhere in the world, giving rise to concerns regarding data privacy and jurisdiction.
 Addressing these multifaceted challenges will require careful consideration and international collaboration to ensure the effective and secure implementation of IoT technology.
@@
 These decisions frequently revolve around determining what is ethically correct or incorrect.
-Figure 1.6 illustrates prevalent such ethical dilemmas within the IoT domain.
-The ethical concerns related to IoT can be divided into five main categories, as adapted from [148], and are outlined below.
-These issues are put forth with the aim of safeguarding privacy rights through the regulation of how organizations manage information generated by IoT devices.
-While these ethical standards are designed to establish guidelines for organizations, they should also prompt individual concerns regarding their privacy, as these principles serve as legal safeguards to protect individuals.
-  • Information privacy considerations
-    Organizations must handle the produced data with both openness and transparency.
-Except in specific cases, they should offer choices to individuals allowing them not to disclose their identity or to use a pseudonym.
-  • Information collection
-    Organizations can gather requested data by implementing more stringent criteria for acquiring ‘sensitive’ information.
+Figure 1.6 illustrates prevalent ethical dilemmas within the IoT domain.
+Figure 1.6: Ethical Issues in IoT
+Adapted from [148], the following outlines the five main categories that divide the ethical concerns related to IoT.
+The aim of putting forth these issues is to safeguard privacy rights by regulating how organizations manage information generated by IoT devices.
+These ethical standards, designed to establish guidelines for organizations, should also prompt individual concerns regarding privacy, as they serve as legal safeguards to protect individuals.
+  1. 1.
+    Information privacy considerations: Organizations must handle the produced data with both openness and transparency.
+Except in specific cases, they should offer choices to individuals, allowing them not to disclose their identity or to use a pseudonym.
+  2. 2.
+    Information collection: Organizations can gather requested data by implementing more stringent criteria for acquiring ‘sensitive’ information.
 Conversely, they should specify their approach to handling unsolicited information.
 In both situations, organizations are required to delineate the circumstances surrounding the collection of this information and provide prior notification to relevant parties.
-  • Managing data
-    Organizations should specify the situations in which they might utilize or share the information they have gathered.
+  3. 3.
+    Managing data: Organizations should specify the situations in which they might utilize or share the information they have gathered.
 Under specific conditions, an organization can employ personal data for direct marketing purposes.
-Nonetheless, they have the option of disclosing it internationally; but, before doing so, they must set the safeguards that will be used to protect this information.
-  • Integrity of information
-    The information gathered and shared by entities should be precise, current, and comprehensive.
+Nonetheless, they have the option of disclosing it internationally, but before doing so, they must set the safeguards that will be used to protect this information.
+  4. 4.
+    Integrity of information: Entities should gather and share precise, current, and comprehensive information.
 Reasonable precautions must be in place to prevent misuse, interference, loss, and unauthorized access, changes, or disclosure.
-  • Information rectification and availability
-    When requesting access to their information, entities should clearly state their responsibilities for allowing access and making corrections to the information they possess.
+  5. 5.
+    Information rectification and availability: When requesting access to their information, entities should clearly state their responsibilities for allowing access and making corrections to the information they possess.
 This involves the obligation to grant access and make necessary modifications, except in cases where a particular exception is applicable.
@@
 After identifying the challenges and ethical concerns, there is a need to consider the legal aspects concerning the effectiveness of current laws in safeguarding users within this context.
-This concern is significant because the boundary between the physical and virtual realms in IoT is increasingly blurred.
-The ensuing questions serve as examples of the issues that require discussion, as referenced in [147].
-  1. 1.
-    What are the consequences in the event of an Internet connection failure?
-  2. 2.
-    Who bears the responsibility or liability for maintaining the security of IoT devices, routers, and connections to the cloud?
-  3. 3.
-    Can one be certain that unauthorized access to a home’s internal network will not occur due to hacking within IoT services on the cloud side?
-  4. 4.
-    What are the consequences when an IoT service provider encounters a period of unavailability for essential life-support devices?
-  5. 5.
-    What are the consequences of an IoT device operating without the explicit consent of its owner or functioning in unintended manners, such as erroneously placing orders or engaging in excessive vacuuming during unconventional hours?
-  6. 6.
-    What occurs if the vendor of an IoT product ceases operations and discontinues product support?
-  7. 7.
-    To whom does the data gathered by IoT devices belong?
-  8. 8.
-    Are there situations in which it is inappropriate for IoT devices to gather data?
-  9. 9.
-    What occurs when the user decides to withdraw their participation?
-  10. 10.
-    How can one address the issue of individuals who lack access to smart devices or the skills to operate them? (Digital divide)
-Considering the challenges mentioned earlier, it is evident that IoT faces several obstacles, including financial constraints, security vulnerabilities, and data privacy concerns, which can have life-threatening implications, especially in healthcare data breaches.
+The significance of this concern arises from the increasing blurring of the boundary between the physical and virtual realms in IoT.
+The ensuing questions, depicted in Fig.
+1.7, serve as examples of the issues that require discussion, as referenced in [147].
+Figure 1.7: Legal and Regulatory Issues of IoT
+Considering the challenges mentioned in Section 1.7, it is evident that IoT faces several obstacles, including financial constraints, security vulnerabilities, and data privacy concerns, which can have life-threatening implications, especially in healthcare data breaches.
 To address these challenges effectively, a proactive approach involving extensive research is crucial.
@@
 The introduction of new laws and standards is also necessary, integrating existing regulations like HIPPA, FIPPS, the Electronic Communication Privacy Act, and others to comprehensively address security, privacy, and legal issues.
-Additionally, addressing technical challenges involves the introduction of adaptable and new standards, as well as the implementation of standard address identification.
+Additionally, addressing technical challenges involves the introduction of adaptable and new standards as well as the implementation of standard address identification.
 Some examples of these technical solutions include advanced encryption techniques, electronic signatures, the integration of standard protocols, and regulations to limit third-party data usage.
-This holistic approach aims to overcome the multifaceted challenges that IoT faces.
+This holistic approach aims to overcome the multifaceted challenges that the IoT faces.
 1.8 Conclusions
 The IoT has rapidly become an integral part of the 21st century, enhancing daily decision-making and ushering in innovative consumer services like pay-as-you-use.
-The seamless integration of smart devices and automation technologies has revolutionized our lives.
+The seamless integration of smart devices and automation technologies has revolutionized every aspect of our lives.
 However, amidst this technological marvel, we must acknowledge significant concerns related to security, privacy, intellectual property rights, safety, and trust.
 These concerns continue to demand further investigation.
-This chapter has provided a comprehensive overview of IoT for newcomers seeking to explore this domain and gain a thorough understanding to make future contributions.
-Specifically, the chapter begins by discussing the fundamental concept of IoT, its historical development, network components, and key characteristics.
-Subsequently, it explores the advantages of employing IoT in various domains and introduces readers to the fundamental layered architectures that underpin IoT.
-Furthermore, the chapter delves into the taxonomy of IoT technologies, which not only encompasses hardware, software, and wireless communication technologies but also various hardware platforms and cloud solutions.
-It also reviews existing applications across major domains such as smart cities, healthcare, and agriculture, offering readers insights into the breadth of IoT applications.
-Finally, the chapter addresses the prevalent issues and challenges associated with the deployment and design of IoT applications.
-It explores security threats across architectural layers, ethical dilemmas, privacy concerns faced by users, and issues related to trust.
+This chapter has provided a comprehensive overview of IoT for newcomers seeking to explore this domain and gain a thorough understanding to make future contributions.The chapter covers fundamental IoT concepts, historical development, architectures, advantages, and technology taxonomy.
+It explores diverse applications in domains such as smart cities and healthcare while addressing challenges and providing possible future directions.
 This discussion serves as a solid foundation for researchers interested in developing practical IoT projects or pioneering new theoretical approaches within the IoT field, equipping them with a deep understanding of various aspects of IoT.
-This, in turn, provides a good ground for the researchers who are whether interested in designing realistic IoT projects or developing novel theoretical approaches in the IoT field by acquiring deep knowledge in different IoT aspects.
+This, in turn, provides a good ground for researchers who are interested in designing realistic IoT projects or developing novel theoretical approaches in the IoT field by acquiring deep knowledge in different IoT aspects.